### v2

# A Survey on Vision-Language-Action Models for Embodied AI

Abstract
Deep learning has demonstrated remarkable success across many domains, including computer vision, natural language processing, and reinforcement learning. Representative artificial neural networks in these fields span convolutional neural networks, Transformers, and deep Q-networks. Built upon unimodal neural networks, numerous multi-modal models have been introduced to address a range of tasks such as visual question answering, image captioning, and speech recognition. The rise of instruction-following robotic policies in embodied AI has spurred the development of a novel category of multi-modal models known as vision-language-action models (VLAs). Their multi-modality capability has become a foundational element in robot learning. Various methods have been proposed to enhance traits such as versatility, dexterity, and generalizability. Some models focus on refining specific components. Others aim to develop control policies adept at predicting low-level actions. Certain VLAs serve as high-level task planners capable of decomposing long-horizon tasks into executable subtasks. Over the past few years, a myriad of VLAs have emerged, reflecting the rapid advancement of embodied AI. Therefore, it is imperative to capture the evolving landscape through a comprehensive survey.
Index Terms:
Embodied AI, Robotics, Multi-modality, Computer Vision, Natural Language Processing, Large Language Model
I Introduction
[…]
rect/.append style=rectangle, rounded corners=2pt, dir tree switch/.style args=at #1for tree= fit=rectangle, , where level=#1 for tree= folder, grow’=0, , delay=child anchor=north, , before typesetting nodes= for tree= content/.wrap value=##1, , if=isodd(n_children(”!r”)) for nodewalk/.wrap pgfmath arg=fake=r,n=##1calign with current edgeint((n_children(”!r”)+1)/2), , , ,
{forest}
[…]
Figure 1: The taxonomy of Vision-Language-Action models.
Vision-language-action models (VLAs) are a class of multi-modal models within the field of embodied AI, designed to process information from vision, language, and action modalities. Unlike conversational AI such as ChatGPT [1], embodied AI requires controlling physical embodiments that interact with the environment. Consequently, robotics is the most prominent domain of embodied AI. In language-conditioned robotic tasks, the policy must possess the capability to understand language instructions, visually perceive the environment, and generate appropriate actions, necessitating the multi-modal capabilities of VLAs. The term was recently coined by RT-2 [2]. Compared to earlier deep reinforcement learning approaches, VLAs offer superior versatility, dexterity, and generalizability in complex environments. As a result, they are well-suited not only for controlled settings like factories but also for everyday tasks in household environments [3].
Early developments in deep learning primarily consist of unimodal models. In computer vision (CV), models like AlexNet [4] showcased the potential of artificial neural networks (ANNs) [5]. Recurrent neural networks (RNNs) [6] laid the groundwork for numerous natural language processing (NLP) models but have seen a transition in recent years, with Transformers [7] taking precedence. The Deep Q-network demonstrated that ANNs can successfully tackle reinforcement learning problems. Leveraging advancements of unimodal models across diverse machine learning fields, multi-modal models have evolved to be powerful enough to tackle a variety of tasks [8], such as visual question answering, image captioning, speech recognition, etc.
Conventional robot policies based on reinforcement learning mostly concentrate on a restricted set of tasks, typically within controlled environments like factories and laboratories. For instance, [9] trains a policy specifically for grasping items. However, there is a growing need for more versatile multi-task policies, akin to recent developments in large language models (LLMs) [10, 1] and vision-language models (VLMs) [11]. Developing a multi-task policy is more challenging as it requires learning a broader set of skills and adapting to dynamic and uncertain environments. Additionally, task specification adds another layer of complexity. Some approaches use a one-hot vector to select tasks [12], but they are limited by the number of tasks in the training set.
Built upon the success of pretrained vision foundation models, LLMs, and VLMs, Vision-Language-Action models have demonstrated their competency in addressing these challenges. Pretrained visual representations from state-of-the-art vision encoders assist VLAs in perceiving complex environments [13, 14, 15], providing more precise estimations such as object class, object pose, and object geometry. With the increasing power of language models [10, 2], task specification based on language instructions becomes a viable option. Numerous ways of integrating vision models and language models have been explored by foundation VLMs, including BLIP-2 [16], Flamingo [11], etc. These innovations from different fields empower VLAs to address the challenges of embodied AI.
Different VLAs exhibit distinct emphases, as depicted by the taxonomy in Figure 1. Some works strive to enhance specific constituents of VLAs, such as pretrained visual representations, dynamics learning, and world models. Meanwhile, a substantial body of work is dedicated to robotic control policy. In this category, language instructions are input into the control policy, which generates primitive actions based on the environment. In contrast, another category of VLAs functions as high-level task planners that abstract away low-level controls. Instead, these models focus on decomposing long-horizon robotic tasks into subtasks. These subtasks can then be completed by control policies one by one, eventually fulfilling the overall task.
Related Work. Although surveys for VLAs are currently lacking, existing surveys in related domains provide valuable insights for VLA research. In the field of computer vision, surveys cover a wide range of visual models, ranging from convolutional neural networks [17] to Transformers [18]. Natural language processing models are comprehensively summarized in surveys [19, 20]. In-depth reviews of reinforcement learning can be found in surveys [21, 22, 23]. Surveys on graph neural networks are also available [24]. Furthermore, existing surveys comparing vision-language models offer inspiration for VLAs [8, 25, 26, 27]. Additionally, a survey on earlier works of embodied AI exists [28].
Contributions. This survey stands as the first thorough review of the emerging vision-language-action model in the field of embodied AI.
  • Comprehensive Review. We present a thorough review of emerging VLA models in embodied AI, covering various aspects including architectures, training objectives, and robotic tasks.
  • Taxonomy. We introduce a taxonomy of the hierarchical structure in current robot systems, comprising three main components: pretraining, control policy, and task planner. Pretraining techniques aim to enhance specific aspects of VLAs, such as the vision encoder or dynamics model. Low-level control policies execute low-level actions based on specified language commands and the perceived environment. High-level task planner decomposes long-horizon tasks into subtasks executable by control policies.
  • Abundant Resources. We provide an overview of necessary resources for training and evaluating VLA models. We investigate recently introduced datasets and simulators by comparing their key characteristics. Additionally, we include widely-adopted benchmarks for tasks such as robot control and embodied reasoning.
  • Future Directions. We outline current challenges and promising future opportunities in the field, such as addressing data scarcity, enhancing robot dexterity, achieving generalization across different tasks, environments, and embodiments, as well as improving robot safety.
Paper Organization. §II-A provides an overview of representative developments and milestones in unimodal models. Because vision-language models are closely related to vision-language-action models, recent advancements in vision-language models are compared in §II-B. §III explores various types of vision-language-action models. §IV summarizes recent datasets and benchmarks for embodied AI. Challenges and future directions are included in §V.
II Background
Figure 2: A brief timeline traces the evolution from unimodal models to multimodal models, laying the groundwork for the introduction of VLA models.
II-A Unimodal Models
Vision-language-action models integrate three modalities, often relying on existing unimodal models. The transition from convolutional neural networks (e.g., ResNet [29]) to visual Transformers (e.g., ViT [30], SAM [31]) in computer vision has facilitated the development of more general vision models. In natural language processing, the evolution from recurrent neural networks (e.g., LSTM [6], GRU [32]) to Transformers [7] initially led to the pretrain-finetune paradigm (e.g., BERT [33], ChatGPT [1]), followed by the recent success of prompt tuning driven by large language models. Reinforcement learning (e.g., DQN [34], AlphaGo [35], PPO [36], Dactyl [37]) has also witnessed a shift towards employing Transformers (§III-A1) to model the Markov Decision Process as autoregressive sequential data.
II-B Vision-Language Models
Vision-language tasks, encompassing image captioning [38], visual question answer [39], visual grounding [40], require the fusion of computer vision and natural language processing models. Early attempts like Show and Tell [41] capitalized on the success of early CNNs and RNNs. The introduction of high-capacity language models such as BERT [33] and GPT [42] ushered in a new era of Transformer-based VLMs, which have evolved in three major directions: self-supervised pretraining, contrastive pretraining, and multi-modal LLMs. One of the pioneering self-supervised pretraining methods is ViLBERT [43], while CLIP [44] popularized contrastive pretraining approaches for multi-modal alignment. The recent emergence of large language models has driven the development of multi-modal LLMs (MLLMs) or large multi-modal models (LMMs), which achieve state-of-the-art performance on multi-modal instruction-following tasks. Representative MLLMs include Flamingo [11], BLIP-2 [16], LLaVA [45]. VLMs share a close relationship with VLAs, as the multi-modal architectures of VLMs can be readily adopted for VLAs. For example, attaching an action decoder to VLMs can transform them into VLAs for low-level control [46]. VLMs can also serve as high-level task planners if they possess sufficient reasoning abilities.
II-C Embodied AI & Robot Learning
Embodied AI is a unique form of artificial intelligence that actively interacts with the physical environment. This sets it apart from other AI models, such as conversational AI, which primarily handles textual conversations (ChatGPT [1]), or generative AI models that focus on tasks like text-to-video generation (Sora [47]). Embodied AI encompasses a broad spectrum of embodiments, including smart appliances, smart glasses, autonomous vehicles, and more. Among them all, robots stand out as one of the most prominent embodiments.
Robot learning is also usually framed as reinforcement learning problems, represented as a Markov Decision Process (MDP) consisting of states ($s$), actions ($a$), and rewards ($r$). A MDP trajectory is denoted as $\tau=(s_{t=1},a_{t=1},r_{t=1},\dots,s_{t=T},a_{t=T},r_{t=T})$. In certain scenarios, robotic tasks may also be viewed as partially-observable Markov Decision Processes (POMDPs) due to incomplete observations. The primary objective of reinforcement learning is to train a policy capable of generating optimal actions for the current state $\pi_{\theta}(a_{t}|p,s_{\leq t},a_{<t})$. Various methods such as TD learning, policy gradients, etc., can be employed to achieve this. However, in cases where defining the reward function proves challenging, imitation learning is utilized to directly model the action distribution within trajectories devoid of rewards $\tau=(s_{t=1},a_{t=1},\dots,s_{t=T},a_{t=T})$. Furthermore, many multi-task robot models employ language as instructions to determine which task or skill to execute, thereby engaging in language-conditioned reinforcement learning.
III Vision-Language-Action Models
Vision-language-action models (VLAs) are the models that handle multi-modal inputs of vision and language and output robot actions to complete embodied tasks. They serve as the cornerstone for instruction following robotic policies in the field of embodied AI. These models rely on robust vision encoders, language encoders, and action decoders. Consequently, some works focus on improving the constituent components of VLA models (Sec III-A). Others concentrate on refining low-level control policies, adept at handling simple instructions and executing primitive actions (Sec III-B). Additionally, certain VLAs abstract away from low-level control and primarily aim to solve task decomposition (Sec III-C). Thus, the combination of low-level control policy and high-level task planner can be treated as a hierarchical policy, as illustrated in Figure 3.
III-A Constituents of VLA
A variety of technologies serve as integral components of VLA models. Pioneering works such as Decision Transformer and Trajectory Transformer have paved the way for using Transformers in reinforcement learning (§III-A1), facilitating the development of scalable and high-capacity policies. The vision encoder is also crucial, as it must effectively encode the state and provide sufficient information about the environment to VLAs. Several works aim to pretrain vision encoders to obtain improved pretrained visual representations (PVRs) (§III-A2). Other research emphasizes learning the dynamics of the environment (§III-A3), enabling VLAs to understand the consequences of their actions and make more informed decisions. An extension of dynamics learning involves developing a world model (§III-A4) that can provide additional information, such as reward signals. This approach allows the model to sample imagined rollouts, reducing the need for interaction with the actual environment and improving data efficiency. Additionally, by leveraging internet-scale commonsense knowledge from large language models, LLM-induced world models (§III-A5) have been gaining popularity.
III-A1 Reinforcement Learning Transformer
Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them conducive to Transformer models. Pioneering efforts in this direction include Decision Transformer [48] and Trajectory Transformer [49]. Decision Transformer focuses on learning a policy, i.e., modeling the actions. Trajectory Transformer shares commonalities with Decision Transformer but distinguishes itself by extending its modeling objectives to include both states and returns within the trajectory. These innovative approaches showcase the applicability of Transformer architectures to the intricacies of RL sequence modeling. Gato [50] further extends this paradigm to a multi-modal, multi-task, multi-embodiment setting.
Figure 3: Illustration of a hierarchical robot policy comprising a high-level task planner and a low-level control policy. The high-level task planner generates a plan based on the user instruction, which is then executed step by step by the low-level control policy.
III-A2 Pretrained Visual Representations
The effectiveness of the vision encoder directly influences the performance of the policy, as it supplies crucial information regarding object categories, positions, and environmental affordance. Consequently, numerous methods are devoted to pretraining the vision encoder to improve the quality of PVRs. Their technical details are compared in Table I.
Although Contrastive Language-Image Pre-training (CLIP) [44] was not initially designed for robotics tasks, it has gained widespread adoption as a vision encoder in robotic models, such as CLIPort [13], EmbCLIP [51], and CoW [52]. The training objective of CLIP is to identify the correct text-image pair among all possible combinations in a given batch. By training to enhance alignment between the vision encoder and the language encoder, CLIP proves particularly effective in tasks where text instructions are provided as input. CLIP undergoes training on the WebImageText dataset (WIT), comprising 400 million image-text pairs. This large-scale training allows CLIP to develop a rich understanding of the relationship between visual and textual information. Notably, CLIP conducts a comprehensive comparison of various vision encoders, exploring different configurations of ResNet [29] and ViT [30]. This analysis provides valuable insights into the trade-off between accuracy and efficiency.
Reusable Representations for Robotic Manipulation (R3M) [53] proposes two main pretraning objectives: time contrastive learning and video-language alignment. The objective of time contrastive learning is to minimize the distance between video frames that are temporally close while simultaneously increasing the separation between frames that are temporally distant. This objective aims to create PVRs that capture the temporal relationships within the video sequence. On the other hand, video-language alignment is to learn whether a video corresponds to a language instruction. This objective enriches the semantic relevance embedded in the PVRs.
Masked Visual Pre-training (MVP) [14] adopts the methodology of Masked Autoencoder (MAE) from computer vision. The MAE involves masking out a portion of input patches to a ViT model and training it to reconstruct the corrupted patches. This approach closely resembles the masked language modeling technique used in BERT [33] and falls within the purview of self-supervised training. MVP extends this MAE objective across diverse robotic datasets, demonstrating that the pre-trained vision encoder significantly enhances performance in downstream manipulation tasks.
Value-Implicit Pre-training (VIP) [54] capitalizes on the temporal sequences present in videos, distinguishing itself from R3M [53] by attracting both the initial and target frames while simultaneously repelling successive frames. This objective aims to capture long-range temporal relationships and uphold local smoothness. While their own experiments demonstrate superior performance compared to R3M in specific tasks, subsequent research presents conflicting findings through more comprehensive evaluations [15].
Visual Cortex (VC-1) [15] conducts an in-depth examination of prior PVRs and introduces an enhanced PVR model by systematically exploring optimal ViT configurations across diverse datasets. Additionally, they perform a comprehensive comparative analysis of their model against previous methods on various manipulation and navigation datasets, shedding light on the critical factors that contribute to the improvement of PVR. Another study [55] also compares previous PVRs obtained under supervised learning or self-supervised learning.
Voltron [56] introduces a novel pretraining objective by incorporating language conditioning and language generation into the masked autoencoding (MAE) objective. Employing an encoder-decoder Transformer structure [7], the pretraining alternates between language-conditioned masked image reconstruction and language generation from masked images. This enhances the alignment between language and vision modalities, consequently resulting in a significantly improved success rate on language-conditioned imitation tasks.
RPT [57] undergoes pretraining with a focus not only on reconstructing visual inputs and robotic actions but also on proprioceptive states. In evaluating three distinct masking schemes, it was noted that token masking, in particular, yielded the most significant improvement in the model’s performance.
GR-1 [58] introduces a novel pretraining task of video prediction tailored for a GPT-style model. This video prediction objective is also subsequently employed during the finetuning phase, specifically utilizing robot data. The rationale behind this approach is that the ability to anticipate future frames contributes to more accurate action prediction. The experimental results, in the domain of robot manipulation, provide empirical support for their assertion.
SpawnNet [59] utilizes a two-stream architecture incorporating adapter layers to fuse features from both a pretrained vision encoder and features learned from scratch. This innovative approach eliminates the necessity of training the pretrained vision encoders while surpassing the performance of parameter-efficient finetuning (PEFT) methods, as evidenced by experimental results in robot manipulation tasks.
Table I: Pretrained visual representations. V: Vision. L: Language. Net: Backbone network. CL: contrastive learning. MAE: masked auto-encoding. TFM: Transformer. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation. [SC]: self-collect data. For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc. $\mathcal{S}(\cdot)$ is the similarity measurement. (Ego-Data): Ego4D [60], Epic Kitchens [61], Something-Something-v2 [62], 100DOH [63].
Model V-Encoder Net Objective Formula Notation Dataset Robotic Tasks
CLIP [44] ViT-B [30] - VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair WIT [SC] (EmbCLIP [51])
R3M [53] ResNet-50 [29] - Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Ego4D Sim-Mani: Meta-World, Franka Kitchen, Adroit
MVP [14] ViT-B
ViT-L
- MAE [64] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch (Ego-Data) Real-Mani (xArm7): pick, reach block, push cube, close fridge
VIP [54] ResNet-50 - Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{0},x_{j})\right)}{\exp\left(\mathcal{S}(x_{0},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{j})\right)}$ $x_{i}$: $i$-th video frame, $0<i<j$ Ego4D Real-Mani (Franka): pick and place, close drawer, push bottle, fold towel
VC-1 [15] ViT-L - MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP, including most previous datasets: (Ego-Data), Real Estate 10K [65], OpenHouse24 [SC]) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0;
Sim-Navi: Habitat
Voltron [56] TFM [7] MAP [66] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
$\sum_{i=1}^{N^{\prime}}-\log P\left(y_{i}\mid x,y_{<i}\right)$
$x_{i}$: image patch;
$y$: language instruction
Something-Something-v2 Sim-Mani: Franka Kitchen;
Real-Mani (Franka): custom study desk
RPT [57] ViT [30, 14] TFM MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x_{i}$: token from any modality; $y,z$: other modalities [SC] Real-Mani (Franka): stack, pick, pick from bin
GR-1 [58] ViT-MAE TFM MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: image patch; $y$: language instruction Ego4D Sim-Mani: CALVIN
[…]
Table II: Dynamics learning methods for VLAs. $f(\cdot)$ is the dynamic model. Fwd: forward. Inv: Inverse.
Model V-Encoder Objective Formula Notation Robotic Tasks
Vi-PRoM [71] ResNet [29] Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$ predicts the frame index in original seq. $x$, given a shuffled seq. $x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open and close drawer/door
MaskDP [72] ViT-MAE [64] MAE (implicit fwd & inv dynamics) $\sum_{i=1}^{T}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: state or action token; $y$: the other modality Sim: DeepMind Control Suite
MIDAS [73] ViT, Mask R-CNN [74] Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: VIMA-Bench
SMART [75] CNN Forward & inverse dynamics $\sum_{i=1}^{T}\big{(}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$
$+\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)\big{)}$
$s_{t}$, $a_{t}$: state, action Sim: DeepMind Control Suite
PACT [76] ResNet-18, PointNet [77] Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Navi: Habitat; Real-Navi (MuSHR vehicle)
VPT [78] ResNet Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim: Minecraft
III-A3 Dynamics Learning
Dynamics learning encompasses objectives aimed at imbuing the model with an understanding of forward or inverse dynamics. Forward dynamics involves predicting the subsequent state resulting from a given action, whereas inverse dynamics entails determining the action required to transition from a previous state to a known subsequent state. Some research approaches also frame these objectives as reordering problems for shuffled state sequences. While forward dynamics models closely relate to world models, this subsection focuses specifically on works that leverage dynamics learning as an auxiliary task to enhance the performance of the primary robot task. We compare them in Table II.
Vi-PRoM [71] presents three distinct pretraining objectives. The first involves a contrastive self-supervised learning objective designed to distinguish between different videos. The remaining two objectives are centered around supervised learning tasks: temporal dynamics learning, aimed at recovering shuffled video frames, and image classification employing pseudo labels. Through a comprehensive comparison with preceding pretraining methods, Vi-PRoM demonstrates its effectiveness for behavior cloning and PPO.
MIDAS [73] introduces an inverse dynamics prediction task as part of its pretraining. The objective is to train the model to predict actions from observations, formulated as a motion-following task. This approach enhances the model’s understanding of the transition dynamics of the environment.
SMART [75] presents a pretraining scheme encompassing three distinct objectives: forward dynamics prediction, inverse dynamics prediction, and randomly masked hindsight control. The forward dynamics prediction task involves predicting the next latent state, while the inverse dynamics prediction task entails predicting the last action. In the case of hindsight control, the entire control sequence is provided as input, with some actions masked, and the model is trained to recover these masked actions. The incorporation of the first two dynamics prediction tasks facilitates capturing local and short-term dynamics, while the third task is designed to capture global and long-term temporal dependencies.
MaskDP [72] features the masked decision prediction task, wherein both state and action tokens are masked for reconstruction. This masked modeling task is specifically crafted to equip the model with an understanding of both forward and inverse dynamics. Notably, in contrast to preceding masked modeling approaches like BERT or MAE, MaskDP is applied directly to downstream tasks
Perception-Action Causal Transformer (PACT) [76] introduces a pretraining objective aimed at modeling state-action transitions. PACT receives sequences of states and actions as input, and predicts each state and action token in an autoregressive manner. This pretrained model serves as a dynamics model, which can then be finetuned for various downstream tasks such as localization, mapping, and navigation.
Video Pretraining (VPT) [78] proposes a method that harnesses unlabeled internet data to pretrain a foundational model for the game of Minecraft. The approach begins by training an inverse dynamics model using a limited amount of labeled data, which is then utilized to label internet videos. Subsequently, this newly auto-labeled data is employed to train the VPT foundation model through behavior cloning. This methodology follows semi-supervised imitation learning. As a result of this process, the model demonstrates human-level performance across a multitude of tasks.
III-A4 Classical World Model
Dreamer [79] employs three primary modules to construct a latent dynamics model: the representation model, responsible for encoding images into latent states; the transition model, which captures transitions between latent states; and the reward model, predicting the reward associated with a given state. Under the actor-critic framework, Dreamer utilizes an action model and a value model to learn behavior through imagination by propagating analytic gradients through the learned dynamics. Building upon this foundation, DreamerV2 [80] introduces a discrete latent state space along with an improved objective. DreamerV3 [81] extends its focus to a broader range of domains with fixed hyperparameters. DayDreamer [82] applies this method to physical robots performing real-world tasks.
Masked World Model (MWM) [83] innovates by modifying the vision encoder of DreamerV2 to a hybrid composition of Convolutional Neural Network and Vision Transformer. In training this novel Convolutional-ViT vision encoder, MWM draws inspiration from the approach proposed in MAE [64]. Through the incorporation of an auxiliary reward prediction loss, the resulting learned latent dynamics model exhibits a notable performance improvement across various visual robotic tasks.
Iso-Dream [84] introduces two key enhancements to the Dreamer framework. The first enhancement focuses on optimizing inverse dynamics by decoupling controllable and noncontrollable dynamics. The second enhancement involves optimizing agent behavior based on the decoupled latent imaginations. This approach is particularly advantageous, as the noncontrollable state transition branch can be rolled out independently of actions, yielding benefits for long-horizon decision-making processes.
Transformer-based World Model (TWM) [85] shares the same motivation as Dreamer but adopts a different approach by constructing a world model based on the Transformer-XL architecture [86]. This Transformer-based world model trains a model-free agent in latent imagination by generating new trajectories. Additionally, TWM suggests a modification to the balanced KL divergence loss from DreamerV2 and introduces a novel thresholded entropy loss tailored for the advantage actor-critic framework.
IRIS [87] employs a GPT-like autoregressive Transformer as the foundation of its world model, with a VQ-VAE serving as the vision encoder. Subsequently, a policy is trained using imagined trajectories, which are unrolled from a real observation by the world model, akin to TWM’s approach.
SWIM [88] advocates for the use of human videos in training a world model due to the availability of large-scale human-centric data. However, acknowledging the substantial gap between human and robot data, SWIM addresses this disparity by grounding actions in visual affordance maps. This approach involves inferring target poses based on the affordance maps, facilitating an effective transfer of knowledge from human data to enhance robot control.
Genie [89] introduces a new class of generative models, termed Generative Interactive Environments. It consists of three main components: a spatiotemporal video tokenizer, an autoregressive dynamics model, and a latent action model. After being trained on unlabeled videos in an unsupervised manner, Genie allows users to interact with the generative environment on a frame-by-frame basis. Consequently, Genie establishes a foundation world model.
III-A5 LLM-induced World Model
DECKARD [90] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs, specifically tailored for the task of item crafting in Minecraft. DECKARD iterates between two phases: in the Dream phase, it samples a subgoal guided by the AWM; in the Wake phase, DECKARD executes the subgoal and updates the AWM through interactions with the game. This guided approach, centered around AWMs, enables DECKARD to achieve markedly faster item crafting compared to baselines lacking such guidance.
LLM-DM [91] uses an LLM to construct world models in planning domain definition language (PDDL)—a capability that LLM+P [92] did not achieve, as its PDDL world model was hand-crafted. The LLM also acts as an interface, mediating between the generated PDDL model and corrective feedback from syntactic validators and human experts. Finally, the PDDL world model serves as a symbolic simulator, assisting the LLM planner in generating plans.
RAP [93] repurposes an LLM to act as both a policy that predicts actions and a world model that provides the state transition distribution. Unlike previous chain-of-thought prompting methods, RAP incorporates Monte Carlo Tree Search (MCTS) to enable principled planning, allowing the LLM to build a reasoning tree incrementally. This reasoning strategy helps RAP find a high-reward path that balances exploration and exploitation. Tree-Planner [94] improves efficiency by prompting the LLM only once to generate diverse paths within an action tree.
LLM-MCTS [95] builds upon RAP but extends the problem setting to Partially Observable Markov Decision Processes (POMDPs). It proposes using the LLM as both a world model and a heuristic policy within an MCTS planning algorithm. As a world model, the LLM generates the initial belief of the current state; as a policy, it serves as a heuristic to guide action selection. By leveraging its commonsense knowledge, the LLM reduces the search space of MCTS, thereby improving search efficiency.
E2WM [96] is an LLM training paradigm that focuses on collecting embodied experiences from world models using MCTS. Unlike prior works, E2WM treats simulators as world models, as they emulate real-world physical interactions, and MCTS is used solely for data collection. By fine-tuning LLMs with this embodied knowledge, the models gain enhanced embodied abilities, such as plan generation, activity recognition, and tracking.
3D-VLA [97] proposes a 3D world model capable of goal generation. It processes visual inputs, such as images, depth map, and point clouds, and then generates a goal state—either as an image or a point cloud—using diffusion models in response to the user’s query. The generated goal state can subsequently be used to guide robot control.
III-A6 Reasoning
Reasoning has become an important capability in LLMs, as pioneered by chain-of-thought (CoT) methods [98, 99]. Recent OpenAI o1 models [100] further demonstrate its potential in building more powerful AI systems. A similar trend has also emerged in the field of embodied AI. EmbodiedGPT [101] uses CoT to enhance embodied planning, and ThinkBot [102] applies it to recover missing action descriptions. RAT [103] combines retrieval-augmented generation and CoT to elicit context-aware reasoning, improving long-horizon generation. ReAct [104] interleaves verbal reasoning traces and actions to address diverse decision-making tasks. Tree-Planner [94] employs a Tree-of-Thoughts approach. Text2Motion [105] uses symbolic and geometric reasoning to verify the feasibility of an LLM-generated task plan. Reflexion [106] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
Pros and Cons. Pretrained visual representations underscore the importance of the vision encoder, as visual observation plays a crucial role in perceiving the current state of the environment. Consequently, it sets an upper bound on the performance of the overall model. In VLAs, general vision models are pretrained using robot or human data to enhance their capabilities in tasks such as object detection, affordance map extraction, and even vision-language alignment, which are essential for robotic tasks. In contrast, dynamics learning focuses on understanding the transition between states. This involves not only mapping visual observations to good state representations but also comprehending how different actions lead to different states, and vice versa. Existing dynamics learning methods typically aim to capture the relationship between states and actions using simple masked modeling or reordering objectives. On the other hand, a world model aims to completely model the dynamics of the world, enabling the robot models to roll out states multiple steps into the future based on the current state, facilitating better prediction of optimal actions. Therefore, while a world model is more desired, it is also more challenging to achieve.
Table III: Low-level control policies. We also include some non-VLA models as they are closely related, marked in gray text. BC: behavior cloning (action type, cont/disc: continuous/discrete). TFM: Transformer. Xattn: cross-attention. Concat: concatenation. $p$/$s$: prompt/state vision encoder. (SC): self-collect data. LMP: Latent Motor Plans [107]. MPC: model predictive control.
Model V-Encoder L-Encoder VL-Fusion Action-Decoder Objective Training Dataset Robotic Tasks
CLIPort [13] CLIP-ResNet50, Transporter-ResNet [108] CLIP-GPT Hadamard LingUNet [109] BC ($\textbf{SE}(2)$) (SC) Sim: Ravens [108]
MCIL [110] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP + (SC) Sim: 3D Playroom environment
HULC [111],
HULC++ [112]
MCIL CNN Sentence-BERT [113] LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
Language costs [114] CLIP-ViT,
UNet [115]
CLIP-GPT Concat STORM [116] Max Likelihood Est. (cost map) (SC) Sim & Real (Franka): pick, place
Interactive Language [117] ResNet [29] CLIP-GPT Xattn TFM BC (cont) Language-Table (SC) Sim & Real (xArm6): rearrangement
Hiveformer [118] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data [119] Sim: RLBench
PerAct [120] 3D CNN CLIP-GPT Concat PerceiverIO [121] 3D affordance RLBench data [119] Sim: RLBench; Real (Franka): put marker, place food, etc.
Act3D CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
RVT, RVT-2 [122, 123] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data [119], (SC) Sim: RLBench; Real (Franka): stack, press, place
RoboPoint [124] ViT-L/14 Vicuna-V1.5 13B [125] Concat - 2D affordance (project to 3D) (SC) Real (Franka): pick, place
Gato [50] ViT Sent.Piece [126] Concat TFM BC (cont & disc) (SC) Sim & Real (Sawyer): RGB-stacking
RoboCat [127] VQ-GAN ($p,s$) [128] - - TFM BC, Observation prediction Self-improve Sim & Real (Sawyer, Franka, KUKA): stacking, building, lifting, insertion, removal
VIMA [129] ViT, Mask R-CNN [74] T5 [10] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data (SC) Sim (Ravens): VIMA-Bench
BC-Z [130] ResNet18 ($p,s$) [29] USE FiLM [131] MLP BC (cont) (SC) Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
RT-1 [3] EfficientNet [132] USE FiLM TFM BC (disc) Kitchen (SC) Real (Everyday Robots): pick-place, move, knock
MOO [129] OWL-ViT ($p$) [133], EfficientNet ($s$) USE FiLM TFM BC (disc) (SC) Real (Everyday Robots): pick, move near, knock, place upright, place into
Q-Transformer [134] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1 data Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
RT-Trajectory [135] EfficientNet - - TFM BC (disc) (SC) Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
ACT [136] ResNet18 - - CVAE-TFM BC (cont, action chunking) (SC) with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
MT-ACT [137] CNN N/A FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
RT-2 [2] ViT-22B [138],
ViT-4B [139]
PaLI-X [140],
PaLM-E
Concat Symbol-tuning [141] BC (disc), Co-fine-tuning VQA + Kitchen Real: RT-1 evaluation tasks, Language-Table
RT-H [142] ViT-22B [138] PaLI-X 55B [140] Concat Symbol-tuning BC (disc) Diverse + Kitchen Real: Kitchen, Diverse
RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE (SC) Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
OpenVLA [144] DINOv2 [145],
SigLIP [146]
LLaMA 2 [147] (Prismatic-7B VLM [148] Concat Symbol-tuning BC (disc) OXE, DROID [149] Real: BridgeV2, RT-1 evaluation tasks
RoboFlamingo [46] CLIP-ViT-L/14 LLaMA [150], GPT-NeoX [151], MPT [152] Xattn LSTM, TFM, MLP BC (cont) CALVIN data Sim: CALVIN
RoboUniView [153] UVFormer (Model design follows RoboFlamingo) BC (cont) CALVIN data Sim: CALVIN
VoxPoser [154] ViLD [155],
MDETR [156],
OWL-ViT [133],
SAM [31]
GPT-4 [157] Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
UniPi [158] Imagen video [159] T5-XXL [10] Xattn CNN Inverse dynamics PF (SC), Ravens, BridgeV1 Sim: Painting Factory (PF) [160], Ravens
Diffusion Policy [161] ResNet18 [29] - - UNet,
TFM
DDPM [162] (SC) Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
DP3 [163] DP3 Encoder (Model design follows Diffusion Policy) DDIM (SC) Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
SUDD [164] ResNet18 [29] CLIP-GPT Concat UNet,
TFM
DDPM Lang-Guided Data Gen Sim: MuJoCo; Real (UR5e): pick, place
Octo [165] CNN T5-base [10] Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
3D Diffuser Actor [166] (Model design follows Act3D) DDPM RLBench, CALVIN, (SC) Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
MDT [167] CLIP-ViT-B/16 ($p$), Voltron / ResNet18 ($s$) CLIP-GPT Concat DiT [168] DDIM [169] CALVIN, LIBERO, Kitchen (SC) Sim: CALVIN, LIBERO [170]; Real (Franka): pick, push, open, close, etc.
RDT-1B [171] SigLIP T5-XXL Xattn DiT DDPM (Aggregated Datasets) Real (ALOHA dual-arm): wash, pour, fold, etc.
III-B Low-level Control Policy
Through the integration of an action decoder with perception modules, such as vision encoders and language encoders, a control policy network is formed to execute language instructions in both simulated and real-world environments. The diversity among control policy networks arises from the selection of individual modules and the integration strategy. This section explores various approaches to designing a low-level control policy. Based on the backbone of their action decoder, control policies can be broadly categorized into three types: non-Transformer (§III-B1), Transformer-based (§III-B2), and LLM-based (§III-B8). Other approaches focus on specific aspects of control policies, such as multi-modal instructions (§III-B3), 3D Vision (§III-B4), diffusion-based action generation (§III-B5), and motion planning (§III-B7). In Table III, we compare the architectural details of these control policies.
III-B1 Early Non-Transformer Control Policies
CLIPort [13] integrates the vision and language encoders of CLIP [44] with the Transporter network [108], creating a two-stream architecture. In one stream, CLIP’s vision encoder extracts “semantic” information from the RGB image, while in the other stream, the Transporter network extracts “spatial” information from the RGB-D image. The CLIP sentence encoder encodes the language instruction and guides the output action, which is a pair of end-effector poses: the picking and placing poses. CLIPort demonstrates the ability to pick and place objects as specified by the language instruction.
BC-Z [130] processes two types of task instructions: a language instruction or a human demonstration video. A language instruction undergoes encoding using the USE language encoder [172], while the video is encoded with ResNet18 [29]. The environment is presented to the model in the form of an RGB image, also encoded using ResNet18. Then the instruction embedding and the image embedding are combined through the FiLM layer [131], culminating in the generation of actions. This conditional policy is asserted to exhibit zero-shot task generalization to unseen tasks.
MCIL [110] represents a pioneering robot policy that integrates free-form natural language conditioning. This is in contrast to earlier approaches that typically rely on conditions in the form of task id or goal image. MCIL introduces the capability to leverage unlabeled and unstructured demonstration data. This is achieved by training the policy to follow either image or language goals, with a small fraction of the training dataset consisting of paired image and language goals.
HULC [111] introduces several techniques aimed at enhancing robot learning architectures. These include a hierarchical decomposition of robot learning, a multimodal transformer, and discrete latent plans. The Transformer learns high-level behaviors, hierarchically dividing low-level local policies and the global plan. Additionally, HULC incorporates a visuo-lingual semantic alignment loss based on contrastive learning to align VL modalities. HULC++ [112] further integrates a self-supervised affordance model. This model guides HULC to the actionable region specified by a language instruction, enabling it to fulfill tasks within this designated area.
Universal Policy (UniPi) [158] treats the decision-making problem as a text-conditioned video generation problem. To predict actions, UniPi generates a video based on a given text instruction, and actions are extracted from the frames of the video through inverse dynamics. This innovative policy-as-video formulation offers several advantages, including enhanced generalization across diverse robot tasks and the potential for knowledge transfer from internet videos to real robots.
(a) FiLM
(b) Cross-attention
(c) Concatenation
Figure 4: Three common architectures of Transformer-based control policies are characterized by their vision-language fusion methods. FiLM layers (Hadamard product $\odot$) are employed to fuse language and vision in RT-1 models. Some action decoders utilize cross-attention to condition on the instruction. Concatenation ($\oplus$) is the prevailing method of vision-language fusion for action decoders.
III-B2 Transformer-based Control Policies
Interactive Language [117] presents a robotic system wherein the low-level control policy can be guided in real-time by human instructions conveyed through language, enabling the completion of long-horizon rearrangement tasks. The efficacy of such language-based guidance is primarily attributed to the utilization of a meticulously collected dataset containing diverse language instructions, which surpasses previous datasets by an order of magnitude in scale.
Hiveformer [118] places significant emphasis on leveraging multi-view scene observations and maintaining the full observation history for a language-conditioned policy. This approach represents an advancement over previous systems such as CLIPort and BC-Z that only use the current observation. Notably, Hiveformer stands out as one of the early adopters of Transformer architecture as its policy backbone.
BeT [173], an imitation learning variant of the Trajectory Transformer, addresses the noisy and multi-modal nature of non-expert human demonstration data by using k-means-based action discretization coupled with continuous action correction. C-BeT [174] extends BeT by adding a target frame or demonstration as the goal specification. VQ-BeT [175] further improves BeT by incorporating vector quantization instead of k-means, enabling better modeling of long-range actions.
Gato [50] proposes a model that can play Atari games, caption images, and stack blocks, all with a single set of model parameters. This achievement is facilitated by a unified tokenization scheme, harmonizing the input and output across diverse tasks and domains. Consequently, Gato enables the simultaneous training of different tasks. Representing a significant milestone, Gato exemplifies the potential of constructing a “multi-modal, multi-task, multi-embodiment generalist agent”.
RoboCat [127] proposes a self-improvement process designed to enable an agent to rapidly adapt to new tasks with as few as 100 demonstrations. This self-improvement process iteratively finetunes the model and self-generates new data with the finetuned model. Built upon the Gato model [50], RoboCat incorporates the VQ-GAN image encoder [128]. During training, RoboCat predicts not only the next action but also future observations. The effectiveness of the self-improvement process is demonstrated through comprehensive experiments conducted in both simulated and real-world environments under multi-task, multi-embodiment settings.
RT-1 [3], developed by the same team as BC-Z [130], shares similarities with BC-Z but introduces some key distinctions. Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet [132], departing from BC-Z’s use of ResNet18. The language instruction is also encoded using USE [172] and is integrated with image embeddings through FiLM layers [131]. However, RT-1 does not use video as a task instruction, unlike BC-Z. Additionally, RT-1 replaces the MLP action decoder in BC-Z with a Transformer decoder, producing discretized actions. This modification enables RT-1 to attend to past images, enhancing its performance compared to BC-Z.
Q-Transformer [134] extends RT-1 [3] by introducing autoregressive Q-functions. In contrast to RT-1, which learns expert trajectories through imitation learning, Q-Transformer adopts Q-learning methods. Alongside the TD error objective of Q-learning, a conservative regularizer is incorporated to ensure that the maximum value action remains in-distribution. This approach allows Q-Transformer to leverage not only successful demonstrations but also failed trajectories for learning.
RT-Trajectory [135] adopts trajectory sketches as policy conditions instead of relying on language conditions or goal conditions. These trajectory sketches consist of curves that delineate the intended trajectory for the robot end-effector to follow. They can be manually specified through a graphical user interface, extracted from human demonstration videos, or generated by foundation models. RT-Trajectory’s policy is built upon the backbone of RT-1 and trained to control the robot arm to accurately follow the trajectory sketches. This approach facilitates generalization to novel objects, tasks, and skills, as trajectories from various tasks are transferable.
ACT [136] builds a conditional VAE policy with action chunking, requiring the policy to predict a sequence of actions rather than a single one. During inference, the action sequence is averaged using a method called temporal ensembling. RoboAgent [137] extends this approach with its MT-ACT model, demonstrating that action chunking improves temporal consistency. Additionally, RoboAgent introduces a semantic augmentation method that leverages inpainting to augment existing demonstrations.
III-B3 Multi-modal-instruction Control Policies
VIMA [176] places a significant emphasis on multi-modal prompts and the generalization capabilities of models. By incorporating multi-modal prompts, more specific and intricate tasks can be formulated compared to traditional pure text prompts. VIMA introduces four main types of tasks: object manipulation, visual goal reaching, novel concept grounding, one-shot video imitation, visual constraint satisfaction, visual reasoning. These tasks are often challenging or even infeasible to express using only language prompts. VIMA-Bench has been developed to evaluate across four generalizability levels: placement, combinatorial, novel object, novel task.
MOO [129] extends RT-1 to handle multi-modal prompts. Leveraging the backbone of RT-1, MOO incorporates OWL-ViT [129] to encode images within the prompt. By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1. This extension also facilitates new methods of specifying target objects, such as pointing with a finger, and clicking on graphical user interfaces.
III-B4 Control Policies with 3D Vision
We live in a 3D world, and it is intuitive that using 3D representations as visual inputs should provide richer information than 2D images. Point clouds are a popular choice for representing 3D inputs due to their straightforward derivation from RGBD inputs, as demonstrated by both DP3 and 3D Diffuser Actor. However, voxels have also been explored in various works. RoboUniView [153] shows improved performance by injecting 3D information into RoboFlamingo through a novel UVFormer module as its vision encoder, which provides 3D occupancy information from multi-perspective images. VER [177] also proposes voxelizing multi-view images into 3D cells in a coarse-to-fine manner, enhancing performance in vision-language navigation tasks.
Perceiver-Actor (PerAct) [120] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations. This approach offers a robust structural prior for action learning, enabling natural processing of multi-view observations and facilitating data augmentation in 6-DoF. In this framework, the input to the model comprises voxel maps reconstructed from RGBD images, while the output corresponds to the best voxel for guiding the gripper’s movement. By adopting this formulation, PerAct facilitates efficient task learning even with a few demonstrations. On the other hand, Act3D [178] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization. RVT, RVT-2 [122, 123] propose re-rendering images from virtual views and using these images as inputs, rather than directly relying on 3D inputs. RoboPoint [124] leverages VLM to predict affordance points directly on 2D images, which are then projected into 3D space using depth maps.
SceneScript [179] introduces a set of structured language commands that can define an entire scene, specifying the layout and objects. This language-based scene representation distinguishes itself from previous methods—such as meshes, voxel grids, point clouds, or radiance fields—by generating scenes in an autoregressive, token-based fashion. It achieves competitive results against state-of-the-art methods in layout estimation and object detection.
III-B5 Diffusion-based Action Generation
Diffusion-based action generation leverages the success of diffusion models in the field of computer vision [162]. Diffusion policy [161] formulates a robot policy as Denoising Diffusion Probabilistic Models (DDPMs) [162]. This approach incorporates a variety of innovative techniques, including receding horizon control, visual conditioning, and the time-series diffusion transformer. The effectiveness of this diffusion-based visuomotor policy is underscored by its proficiency in multimodal action distributions, high-dimensional action spaces, and training stability.
Scaling Up and Distilling Down (SUDD) [164] presents a framework where LLM guides data generation and subsequently the filtered dataset is distilled into a visuo-linguo-motor policy. This framework achieves language-guided data generation by composing LLM with a suite of primitive robot utilities, such as grasp samplers, motion planners. Then it extends Diffusion Policy [161] through the incorporation of language-based conditioning for multi-task learning. This diffusion-based policy learns from successful trajectories, facilitating the distillation of the filtered dataset.
Octo [165] introduces a transformer-based diffusion policy characterized by a modular open-framework design, allowing for flexible connections from different task definition encoders, observation encoders, and action decoders to the Octo Transformer. Being among the first to utilize the Open X-Embodiment dataset [143], Octo demonstrates positive transfer and generalizability across diverse robots and tasks.
MDT [167] adapts the newly introduced DiT model [168] from computer vision to the action prediction head. DiT, originally proposed as a transformer-based diffusion model, replaces the classical U-Net architecture for video generation. Coupled with two auxiliary objectives—masked generative foresight and contrastive latent alignment—MDT demonstrates better performance than the U-Net-based diffusion model, SUDD.
RDT-1B [171] is a diffusion-based foundation model for bimanual manipulation, also built on DiT. It addresses data scarcity by introducing a unified action format across various robots, enabling pre-training on heterogeneous multi-robot datasets with over 6K trajectories. As a result, RDT scales up to 1.2B parameters and demonstrates zero-shot generalization.
III-B6 Diffusion-based Control Policy with 3D Vision
Several works have proposed combining 3D Vision with diffusion-based policies. DP3 [163] introduces 3D inputs to a diffusion policy, resulting in improved performance. Similarly, 3D Diffuser Actor [166] shares the core idea of DP3 but differs in model architecture by combining Act3D with Diffusion Policy.
III-B7 Motion Planner
Language costs [114] presents a novel approach to robot correction using natural language for human-in-the-loop robot control systems. This method leverages predicted cost maps generated from human instructions, which are then utilized by the motion planner to compute the optimal action. This framework enables users to correct goals, specify preferences, or recover from errors through intuitive language commands.
VoxPoser [154] employs LLM and VLM to create two 3D voxel maps that represent affordance and constraint. It leverages the programming capability of LLM [157], and the perception capability of VLM models: ViLD [155], MDETR [156], OWL-ViT [133], SAM [31]. LLM translates language instructions into executable code, invoking VLM to obtain object coordinates. Based on the composed affordance and constraint maps, VoxPoser employs model predictive control to generate a feasible trajectory for the robot arm’s end-effector. Notably, VoxPoser does not require any training, as it directly connects LLM and VLM for motion planning.
RoboTAP [180] breaks down demonstrations into stages marked by the opening and closing of the gripper. In each stage, RoboTAP uses the TAPIR algorithm to detect active points that track the relevant object from the source pose to the target pose. The path can then be used by classical visual servoing to control the robot. A motion plan is created by stringing together these stages, enabling few-shot visual imitation.
III-B8 LLM-based Control Policies
RT-2 [2] endeavors to harness the capabilities of large multi-modal models in robotics tasks, drawing inspiration from models like PaLI-X [140] and PaLM-E [181]. The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data. This training scheme enhances the model’s generalizability and brings about emergent capabilities. RT-2 represents an effort to integrate low-level control policy [3] with high-level task planners in pursuit of a more comprehensive robotic system.
RT-X [143] builds upon the previous RT-1 and RT-2 models. These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets. OXE comprises 160,266 tasks from 527 skills with 22 embodiments. The resulting models, RT-1-X and RT-2-X, both outperform their original versions, owing to the more diverse and larger OXE dataset. OpenVLA [144] was later developed as an open-source counterpart to RT-X.
RT-H [142] introduces an action hierarchy that includes an intermediate prediction layer of language motions, situated between language instructions and low-level actions (translation and rotation). This additional layer facilitates improved data sharing across different tasks. For example, both the language instructions “pick” and “pour” may involve the language motion “move the arm up”. Moreover, this action hierarchy enables users to specify corrections to recover from failures, which the model can then learn from.
RoboFlamingo [46] adapts the existing VLM, Flamingo [11, 182], to a robot policy by attaching an LSTM-based policy head to the VLM. This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
Pros and Cons.
Architectures
Various VLA architectures explore different methods of fusing vision and language inputs, including cross-attention, FiLM, and concatenation, as illustrated in Figure 4. FiLM is used in RT-1 and thus its follow-up works inherit this mechanism. While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [176].
Action types and their training objectives
Most low-level control policies predict actions for the end-effector pose while abstracting away the motion planning module that controls the motion of individual joints with inverse kinematics. While this abstraction facilitates better generalization to different embodiments, it also imposes limitations on dexterity.
The behavior cloning (BC) objective is used in imitation-learning, with different variants for different action types. The BC objective for continuous action [46] can be written as:
$\mathcal{L}_{\text{Cont}}=\sum_{t}\operatorname{MSE}\left(\mathbf{a}_{t},\hat{\mathbf{a}}_{t}\right),$ (1)
where $\operatorname{MSE}(\cdot)$ stands for mean squared error. $\hat{\mathbf{a}}_{t}=\operatorname{argmax}_{a_{t}}\pi_{\theta}(a_{t}|p,s_{\leq t},a_{<t})$ is predicted action while $\mathbf{a}_{t}$ is the action from expert demonstrations.
Discrete action is achieved by dividing the action value range into a number of bins [3]. Its BC objective is:
$\mathcal{L}_{\text{Disc}}=\sum_{t}\operatorname{CE}\left(\mathbf{a}_{t},\hat{\mathbf{a}}_{t}\right),$ (2)
where $\operatorname{CE}(\cdot)$ stands for cross-entropy loss.
CLIPort [13] and VIMA [176] use SE(2) action and its BC objective can be expressed as follows:
$\begin{split}\mathcal{L}_{\text{{SE}(2)}}&=\operatorname{CE}(\mathbf{a}_{\text{pick}},\hat{\mathbf{a}}_{\text{pick}})+\operatorname{CE}(\mathbf{a}_{\text{place}},\hat{\mathbf{a}}_{\text{place}})\end{split}$ (3)
The DDPM objective in diffusion-based control policies [161] is represented as:
$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}\left(\mathbf{x}^{0}+\varepsilon^{k},k\right)\right)$ (4)
where $\mathbf{x}^{0}$ is the raw example from the dataset and $\varepsilon^{k}$ is a random noise for iteration $k$; $\varepsilon_{\theta}$ is the noise prediction network, i.e., the VLA model.
While discrete action has demonstrated superior performance in RT-1 [3], Octo [165] argues that it leads to early grasping issues. $\textbf{SE}(2)$ action only necessitates the model to predict the two end-effector poses, e.g., pick pose and place pose. While this action type can be predicted in at most two forward passes, it also imposes limitations on action dexterity and generalizability.
LLM vs non-LLM
While LLM-based control policies can greatly enhance instruction-following abilities because LLM can better parse user intentions, concerns arise regarding their training cost and deployment speed. Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during LLM inference.
RT series
RT-1 [3] inspired a series of “Robotic Transformer” models. Preceding RT-1 was BC-Z [130], which solely utilized MLP layers for action prediction. Subsequent to RT-1, several works emerged, each introducing new capabilities. MOO [129] adapted RT-1 to accommodate multi-modal prompts. RT-Trajectory [135] enabled RT-1 to process trajectory sketches as prompts. Q-Transformer [134] utilized Q-learning to train RT-1. RT-2 [2], based on ViT and LLM, introduced a completely different architecture from RT-1. RT-X [143] retrained RT-1 and RT-2 with a significantly larger dataset, resulting in improved performance. The Transformer backbone surpasses previous RNN backbones [110, 111, 112] by harnessing the higher capacity of Transformers to absorb larger robot datasets. Based on RT-2, RT-H [142] introduces action hierarchies for better data sharing.
(a) Language-based
(b) Code-based
Figure 5: Different approaches to connect LLM to multi-modal modules in high-level task planner.
III-C High-level Task Planner
A high-level task planner aims to divide a complex task into a series of executable sub-tasks. This approach is often referred to as task decomposition, and is also related to the classical Task and Motion Planning (TAMP) setting. Many high-level task planners are constructed on top of LLMs. While integrating multi-modal modules into LLMs in an end-to-end fashion (§III-C1) is a common approach, training with multi-modal data can be expensive. Consequently, some task planners explore to utilize language (§III-C3) or code (§III-C4) as the medium for exchanging multi-modal information instead, as they can be natively processed by LLMs.
III-C1 End-to-end
While control policies are effective in understanding and executing simple language instructions, they tend to struggle in long-horizon tasks that involve multiple sub-tasks. Large language models have been identified as powerful tools to interpret such complex long-horizon tasks. Consequently, several approaches are dedicated to integrating LLMs as high-level task planners. Their goal is to decompose long-horizon tasks into simpler sub-tasks, enabling low-level control policies to execute them sequentially, contributing to the development of a hierarchical robotic system. Table IV outlines important details about these high-level task planners.
SayCan [183] is a framework designed to integrate high-level LLM planners with low-level control policies. In this framework, the LLM planner accepts users’ high-level instruction and “says” what the most probable next low-level skill is, a concept referred to as task-grounding. The low-level policy provides the value function as the affordance function, determining the possibility that the policy “can” complete the skill, known as world-grounding. By considering both LLM’s plan and affordance, the framework selects the optimal skill for the current state.
LID [184] introduces a novel data collection procedure termed Active Data Gathering (ADG). A key aspect of ADG is hindsight relabeling, which reassigns labels to unsuccessful trajectories, effectively maximizing the utilization of data irrespective of their success. By converting all environmental inputs into textual descriptions, their language-model-based policy demonstrates enhanced combinatorial generalization.
Table IV: High-level task planners. VL: vision-language fusion. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation.
Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
SayCan [183] CNN PaLM [185]
FLAN [186]
Concat BC-Z [130]
MT-Opt [12]
Yes Real-Navi&Mani (Everyday Robots): kitchen
LID [184] CNN GPT-2 [187] Concat - Yes Sim: VirtualHome
[…]
Codex [190]
- - No Sim: VirtualHome
(SL)3 [191] ResNet18 [29] T5-small [10] Xattn HLSM [192] Yes Sim-Navi: ALFRED
EmbodiedGPT [101] EVA ViT-G/14 [193] LLaMA-7B [150] Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
PaLM-E [181] ViT [30],
OSRT [194]
PaLM [185] Concat Interactive Language policy [117], RT-1 [3] Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
LEO [195] ConvNeXt [196], PointNet++ [197] Vicuna-7B [125] Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
3D-LLM [198] CLIP, 3D-CLR [199], ConceptFusion [200] Flamingo [11], BLIP-2 [16] Concat DD-PPO [201] Yes Sim-Navi: Habitat
StructFormer [202], StructDiffusion [203] PCT [204] - Concat [205], RRT-connect [206] Yes Sim-Mani: PyBullet; Real-Mani (Franka): rearrange
ShapeLLM [207] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
Inner Monologue [208] MDETR [156], CLIP InstructGPT [209]
PaLM [185]
Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
LLM-Planner [210] HLSM object detector [192] GPT-3 [189] Language HLSM [192] Yes Sim-Navi: ALFRED
SMs [211] CLIP, ViLD [155] GPT-3 [189], RoBERTa [212] Language & Code CLIPort [13] No Sim-Mani: pick, place
ProgPrompt [213] ViLD [155] GPT-3 [189] Code API [214, 215] No Sim: VirtualHome
ChatGPT for Robotics [216] YOLOv8 [217] ChatGPT [1] Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
CaP [218] ViLD [155], MDETR [156] GPT-3 [189]
Codex [190]
Code API No Sim-Mani & Sim-Navi: pick, place, etc.
DEPS [219] CLIP ChatGPT [220] Code MC-Controller, Steve-1 No Sim: Minecraft
ConceptGraphs [221] SAM [31], CLIP, LLaVA [45] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
[…]
Semi-Supervised Skill Learning with Latent Language ((SL)3) [191] is a learning algorithm that alternates between three steps: segmentation, labeling, and parameter update. In the segmentation step, high-level subtasks are aligned with low-level actions, subtask descriptions are then inferred in the labeling step, and finally, the network parameters are updated. This approach enables a hierarchical policy to discover reusable skills with sparse natural language annotations.
EmbodiedGPT [101] introduces the embodied-former, which outputs task-relevant instance-level features. This is achieved by incorporating information from vision encoder embeddings and embodied planning information provided by LLM. The instance feature serves to inform the low-level policy about the immediate next action to take.
PaLM-E [181] integrates ViT [30] and PaLM [185] to create a large embodied multi-modal language model capable of performing high-level embodied reasoning tasks. Based on perceived images and high-level language instructions, PaLM-E generates a text plan that serves as instructions for low-level robotic policies. In a mobile manipulation environment, they map the generated plan to executable low-level instructions with SayCan [183]. As the low-level policy executes actions, PaLM-E can also replan based on changes in the environment. With PaLM as its backbone, PaLM-E can handle both normal visual question answering (VQA) tasks, along with the additional embodied VQA tasks.
III-C2 3D Vision
3D information is vital in embodied tasks, and thus it is also explored for task planners. Because the majority of current VLMs deal with images as visual inputs, it requires architectural change to VLMs to incorporate 3D visual inputs and therefore they are usually end-to-end models.
LEO [195] identifies the conventional use of image inputs to be a limiting factor for multi-modal generalist agents in interacting with the 3D world. This novel approach involves training an LLM-based architecture using a new dataset with two stages. The first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning. LEO demonstrates proficiency not only in 3D captioning and question-answering tasks but also in embodied tasks, including embodied reasoning, embodied navigation, and robotic manipulation.
3D-LLM [198] injects 3D information into LLMs and empowers them to perform 3D tasks, such as 3D-assisted dialog, and navigation. The 3D features can come in diverse forms, including point cloud, gradSLAM, and neural voxel field. MultiPLY [223] is an object-centric embodied LLM that incorporates even more modalities, including audio, tactile, and thermal.
StructFormer [202] addresses the task of arranging objects into complex structures. Previous methods rely on pairwise semantic relations and sequential manipulation, while StructFormer explicitly reasons about the relationships among all objects and predicts their target positions. StructDiffusion [203] extends this by incorporating a diffusion-based generation approach.
ShapeLLM [207] is built on the novel 3D vision encoder ReCon++, which distills knowledge from multi-view image and text teachers, along with a point cloud MAE. By integrating ReCon++ with LLaMA, ShapeLLM improves embodied interaction performance on their newly proposed 3D benchmark, 3D MM-Vet.
Table V: Recent datasets for robot learning. Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects. $\ast$ Dataset collected in simulators rather than in the real world. Some datasets are continuously updated; this table is based on the information from their original papers. This table is partially based on [224, 149].
Name Collection Visual Instruction Embodiment Skills Tasks Scenes Objects Episodes
MIME [225] Human Teleop RGB, D Demo Baxter 12 20 1 - 8.3K
RoboTurk [226] $\ast$ Human Teleop RGB - Sawyer 2 - 1 - 2.1K
RoboNet [227] Scripted RGB Goal image (7 robots) - - 10 - 162K
MT-Opt [12] Scripted RGB Lang (7 robots) 2 12 1 - 800K
BC-Z [130] Human Teleop RGB Lang, Demo Everyday 3 100 1 - 25.9K
RT-1 [3] Human Teleop RGB Lang Everyday 12 700+ 2 16 130K
MOO [129] Human Teleop RGB Multi-modal Everyday 5 - - 106 59.1K
VIMA [176] $\ast$ Scripted RGB Multi-modal UR5 17 - 1 29 650K
RoboSet [228] Human, Scripted RGB, D Lang Franka Panda 12 38 11 - 98.5K
BridgeData V2 [229] Human, Scripted RGB, D Lang WidowX 250 13 - 24 100+ 60.1K
RH20T [224] Human Teleop RGB, D Lang (4 robots) 42 147 7 - 110K+
DROID [149] Human Teleop RGB, D Lang Franka Panda 86 - 564 - 76K
OXE [143] Aggregated Datasets RGB, D Lang (22 robots) 527 160,266 311 - 1M+
III-C3 Language-based
Inner Monologue [208] sits between high-level command and low-level policy to enable closed-loop control planning. It employs LLM to generate language instructions for low-level control policies and dynamically updates these instructions based on feedback received from the control policies. The feedback encompasses various sources: success feedback, object and scene feedback, and human feedback. As the feedback is communicated to the LLM in textual format, no additional training is required for the LLM. A similar approach is also proposed in ReAct [104], which supports more flexible reasoning traces, allowing for more diverse types of reasoning for various tasks.
LLM-Planner [210] introduces a novel approach to constructing a hierarchical policy comprising a high-level planner and a low-level planner. The high-level planner harnesses the capabilities of LLM to generate natural language plans, while the low-level planner translates each subgoal within the plan into primitive actions. While sharing similarities with previous methods in its overall architecture, LLM-Planner distinguishes itself by incorporating a re-planning mechanism, aiding the robot to “get unstuck”.
Socratic Models (SMs) [211] presents a unique framework wherein diverse pretrained models are effectively composed without the need for finetuning. The framework is based on the key component, named multimodal-informed prompting, facilitating information exchange among models with varied multimodal capabilities. The idea is to utilize multimodal models to convert non-language inputs into language descriptions, effectively unifying different modalities within the language space. Beyond excelling in conventional multimodal tasks, SMs showcases its versatility in robot perception and planning. In addition to natural language plans, task plans can also be represented in the form of pseudocode.
III-C4 Code-based
ProgPrompt [213] introduces a novel task-planning approach by prompting LLMs with program-like specifications detailing available actions and objects. This enables LLMs to generate high-level plans for household tasks in a few-shot manner. Environmental feedback can be incorporated through assertions within the program. This prompting scheme leverages the world knowledge and programming skills of LLMs.
ChatGPT for Robotics [216] take advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology. The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code. In this procedure, ChatGPT serves as a high-level task planner, similar to PaLM-E [181], and actions are generated through function calls to corresponding low-level APIs.
Code as policies (CaP) [218] also leverages the code-writing capability of LLMs. It employs GPT-3 [189] or Codex [190] to generate policy code, which, in turn, invokes perception modules and control APIs. CaP exhibits proficiency on spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives. By leveraging the multimodal capabilities of GPT-4V, COME-robot [230] eliminates the need for perception APIs in CaP. This also opens up possibilities for open-ended reasoning and adaptive planning within a closed-loop framework, enabling capabilities such as failure recovery and free-form instruction following.
DEPS [219] stands for “Describe, Explain, Plan and Select”. This approach employs an LLM to generate plans and explain failures based on feedback descriptions collected from the environment—a process referred to as “self-explanation”, aiding in re-planning. Additionally, DEPS introduces a trainable goal selector to choose among parallel candidate sub-goals based on how easily they can be achieved, a crucial aspect often overlooked by other high-level task planners.
ConceptGraphs [221] introduces a method to convert observation sequences into open-vocabulary 3D scene graphs. Objects are extracted from RGB images using 2D segmentation models, and VLMs are employed to caption objects and establish inter-object relations, resulting in the formation of the 3D scene graph. This graph can then be translated into a text description (JSON), offering rich semantic and spatial relationships between entities to LLMs for task planning.
Pros and Cons. While end-to-end task planners like SayCan [183] share a similar architecture to low-level control policies and can be optimized for specific tasks, their training costs can be prohibitive due to the large model size of the combined LLM and visual transformer. Language-based task planners offer the advantage of seamless integration with existing language-conditioned control policies. However, they often require finetuning or alignment methods to map the generated plan to executable language instructions for low-level control policies. On the other hand, code-based task planners leverage the programming capability of LLMs to connect perception and action modules. This approach does not require additional training, but its performance may be constrained by the capabilities of existing models.
IV Datasets and Benchmarks
Table VI: Simulators for benchmarking VLA models. D: Depth. Seg: segmentation. A: Audio. N: Normals. Force: simulated contact force between the end-effector and the item. PD: Pre-defined. Vers: versions. This table is partially based on [231, 232].
Name Scenes
/Rooms
Objects
/Cat
UI Engine Task Observation Action Agent Description Related
Gibson [233] 572/- - - Pybullet Navi RGB, D, N, Seg - - Navi only -
iGibson [234, 235, 232] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, Seg, N, Flow, LiDAR Force Husky, TurtleBot v2, LoCoBot, etc. VR, Continuous Extended States. Vers: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [231], BEHAVIOR-1K [236]
SAPIEN [237] - 2346/- Code PhysX Navi, Mani RGB, D, Seg Force Franka Articulation, Ray Tracing VoxPoser
AI2-THOR [238] -/120 118/118 Mouse Unity Navi, Mani RGB, D, Seg, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Vers: [239, 240] Benchmarks: ALFRED, RoPOR [241]
VirtualHome [242] 7/- -/509 Lang Unity Navi, Mani RGB, D, Seg, F Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
TDW [243] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, Seg, A Force Fetch, Sawyer, Baxter Audio, Fluids -
RLBench [119] 1/- 28/28 Code Bullet Mani RGB, D, Seg Force Franka Tiered Task Difficulty Hiveformer, PerAct
Meta-World [244] 1/- 80/7 Code MuJoCo Mani Pose Force Sawyer Meta-RL R3M, VC-1, Vi-PRoM, EmbodiedGPT
CALVIN [245] 4/- 7/5 - Pybullet Mani RGB, D Force Franka Long-horizon Lang-cond tasks GR-1, HULC, RoboFlamingo
Franka Kitchen [246] 1/- 10/6 VR MuJoCo Mani Pose Force Franka Extended by R3M with RGB R3M, Voltron, Vi-PRoM, Diffusion Policy, EmbodiedGPT
Habitat [247, 248] Matterport + Gibson Mouse Bullet Navi RGB, D, Seg, A Force Fetch, Franka, AlienGO Fast, Navi only. Vers: Rearrangement [249], Habitat 2.0 VC-1, PACT; OVMM [250]
ALFRED [251] -/120 84/84 - Unity Navi, Mani RGB, D, Seg PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
DMC [252] 1/- 4/4 Code MuJoCo Control RGB, D Force - Continuous RL VC-1, SMART
OpenAI Gym [253] 1/- 4/4 Code MuJoCo Control RGB Force - Single agent RL environments -
The collection of real-world robot data presents significant challenges due to various factors. Firstly, the collection process is hindered by the substantial costs associated with procuring robot equipment, setting up environments, and dedicating extensive human resources. Secondly, the gathering of expert demonstration data requires considerable time investments. Thirdly, the diverse types and configurations of robots introduce inconsistency in sensory data, control modes, gripper types, etc. Lastly, achieving precision in capturing object 6D poses and exactly replicating or resetting setups remains elusive. As a result of these difficulties, public real-world robot datasets are relatively scarce. Furthermore, evaluating a robot system’s performance in real-world conditions introduces another layer of complexity, as it is challenging to precisely reproduce a setup and often requires human supervision. We summarize the robot datasets from recent VLAs in Table V.
Consequently, many researchers resort to simulated environments as a means to mitigate these obstacles and accelerate the data collection process. We compare the most relevant simulators in Table VI. Nevertheless, this strategy presents its own challenges, chief among them being the sim-to-real gap. This discrepancy arises when models trained on simulated data exhibit poor performance in real-world deployment. The causes of this gap are multifaceted, encompassing differences in rendering quality, inaccuracies in physics simulations, and domain shifts characterized by unrealistic object properties and robot motion planners. For instance, simulating non-rigid objects like liquids or towels presents significant difficulties. Moreover, incorporating new objects into simulators requires considerable effort, often involving techniques such as 3D scanning and mesh editing. Despite these hurdles, simulated environments provide automated evaluation metrics that aid researchers in consistently evaluating robotic models. Many benchmarks are based on simulators because only simulated environments can reproduce the experimental setup exactly, while real-world evaluation varies between different models, making it infeasible to compare them reliably. The HomeRobot OVMM Benchmark [250] proposes a sim-to-real benchmark, yet its consistency remains to be observed.
Several approaches advocate for automated data collection. RoboGen [254] employs a generative simulation paradigm that proposes interesting skills, simulates corresponding environments, and selects optimal learning approaches to train policies for acquiring those skills. AutoRT [255] functions as a robot orchestrator driven by LLMs, generating tasks, filtering them by affordance, and utilizing either autonomous policies or human teleoperators to collect and evaluate data. DIAL [256] focuses on augmenting language instructions in existing datasets using VLMs. RoboPoint [124] generates scenes procedurally with randomized 3D layouts, objects, and camera viewpoints.
An alternative strategy to address the data scarcity issue in real-world settings is to leverage human data. Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity. However, this strategy also comes with inherent drawbacks. Capturing and transferring human hand/body motions to robot embodiments is inherently difficult. Moreover, the inconsistency in human data poses a hurdle, as some data may be egocentric while others are captured from third-person perspectives. Additionally, filtering human data to extract useful information can be labor-intensive. These obstacles underscore the complexities involved in incorporating human data into robot learning processes. Due to limited space, we refer interested readers to a comprehensive comparison of existing datasets in [257].
Table VII: Embodied question answering benchmarks. Acc: accuracy. PPL: perplexity.
# QA # Videos Video source Answer type Active Data collection Metrics
EQA [258] 5K 750 envs House3D simulator Answer set 172 Yes Template Acc
IQUAD [259] 75K 30 rooms AI2-THOR Multiple-choice Yes Template Acc
MT-EQA [260] 19.3K 588 envs House3D simulator Binary answer Yes Template Acc
MP3D-EQA [261] 1,136 83 envs MatterPort3D Answer set 53 Yes Template Acc
EgoVQA [262] 600 16 IU Multi-view Multiple-choice (1 out of 5) No Human annotators Acc
EgoTaskQA [263] 40K 2K LEMMA Answer set (open-answer and binary verifications) No Human annotators for relationship triples, then template Acc
EgoPlan [264] 3,355 2,432 + 923 Epic-Kitchen + Ego4D Multiple-choice (1 out of 4) No GPT-4 filtering, question template, human verification Acc
OpenEQA [265] 557 + 1079 180 envs HM3D + ScanNet Open-answer (LLM scorer) No Human annotators Score
EgoCOT [101] 5,246 plans 129 Ego4D Caption and planning (not QA) No ChatGPT PPL
EQA-MX [266] 8.2M 750K images CAESAR simulator Answer set No Question templates, answer set (simulator ground truth) Acc
LoTa-Bench [267] 17K instruction-plan pairs AI2-THOR, VirtualHome Goal condition - ALFRED + WAH-NL Success
Furthermore, some datasets and benchmarks may not be directly dedicated to robot manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge. One of the most prominent embodied tasks is embodied question answering (EQA), as summarized in Table VII. EQA is akin to previous visual question answering and video question answering tasks but differs in that the agent can actively explore the environment before providing an answer. EmbodiedQA [258] and IQUAD [259] were among the first works to introduce this task. MT-EQA [260] focuses on questions involving multiple targets, increasing the complexity of understanding and answering the questions. MP3D-EQA [261] converts previous RGB input to point clouds, testing the 3D perception capability. However, active exploration requires access to a simulator, limiting the types of data that can be used, such as real-world videos. Therefore, some EQA benchmarks do not involve active exploration. EgoVQA [262] shifts the focus of VQA to egocentric videos. EgoTaskQA [263] emphasizes spatial, temporal, and causal relationship reasoning. EQA-MX [266] focuses on multimodal expressions (MX), including regular verbal utterances and nonverbal gestures like eye gaze and pointing. OpenEQA [265] evaluates seven main categories, including functional reasoning and world knowledge, which were not covered in previous benchmarks. EgoPlan-Bench [264] and EgoCOT [101] measure the model’s ability to generate task plans, using metrics like accuracy and perlexity. PlanBench [268, 269] comprehensively assesses various aspects of task planning ability, such as cost optimality, plan verification, and replanning, etc. LoTa-Bench [267] directly evaluates task planning by executing the generated plans in simulators and calculating the success rate.
V Challenges and Future Directions
Vision-language-action (VLA) models face several persistent challenges in the domain of Robotics, necessitating focused attention and concerted research efforts:
Scarcity of Robotic Data. Obtaining sufficient real-world robotic data remains a significant obstacle. Collecting such data is time-consuming and resource-intensive, while relying solely on simulation data exacerbates the sim-to-real gap problem. Diverse real-world robotic datasets require close and extensive collaboration among different institutions. Simulated data depends on the development of more realistic and efficient simulators.
Motion Planning. Current motion planning modules often lack the necessary dexterity to address the complexities in various environments. This limitation hampers robots’ abilities to interact effectively with tools, navigate complex environments, and execute high-precision operations, etc. Overcoming these challenges necessitates the development of more robust motion planning algorithms.
Real-Time Responsiveness. Many robotic applications require real-time decision-making and action execution to meet operational requirements. VLA models should be designed to operate responsively, with minimal latency. Furthermore, the entire stack of a robot system necessitates global optimization, from the high-level task planner down to the motion planner.
Integration of Multiple Modalities. VLAs must process and integrate information from multiple modalities, including vision, language, and action. While significant progress has been made in this regard, achieving optimal integration of these modalities remains an ongoing challenge. Addressing this challenge requires advancements in multi-modal representation learning, fusion techniques, and task-specific adaptation. Expanding beyond mere visual and linguistic capacities, robots can greatly benefit from incorporating modalities like audio or speech. Embracing a broader range of perceptual and communicative abilities empowers robots to engage in more effective collaboration with users.
Generalization to Unseen Scenarios. A truly versatile robotics system should be capable of comprehending and executing natural language instructions across diverse and unseen scenarios. Achieving this level of generalization, akin to the capabilities demonstrated by ChatGPT in natural language processing, requires robustness to variations in instructions, environments, objects, and embodiments. This necessitates the development of adaptable and scalable VLA architectures.
Long-Horizon Task Execution. Single instructions can often translate into long-horizon tasks for robots, such as the instruction “clean the room” which involves object rearrangement, floor sweeping, table wiping, and more. Successfully executing such tasks requires the robot to plan and execute a sequence of low-level actions over extended time horizons. While current high-level task planners have achieved initial success, they still fall short in many scenarios because most LLMs are not tuned to embodied tasks. Addressing this challenge entails devising efficient planners that possess powerful perception abilities and a wide range of common sense.
Foundation Model. The exploration of foundation models for VLAs in robot tasks remains uncharted territory, primarily due to the diverse embodiments, environments, and tasks encountered in robotics. Isolated datasets and evaluation setups further compound this challenge. To establish a robust foundation VLA model for robotics, it is imperative to leverage internet-scale embodied datasets and state-of-the-art multi-modal models.
Multi-agent Systems. The large-scale deployment of robots is highly desirable in many scenarios, but it introduces the challenge of multi-agent cooperation. Effective communication, task distribution among multiple robots can be difficult [270]. Fleet heterogeneity may further complicate these issues. While multi-agent systems offer benefits like distributed perception, collaborative fault recovery, and autonomous fleet maintenance, realizing the full potential of these advantages remains elusive.
Benchmarks. Although numerous benchmarks exist for evaluating low-level control policy VLAs, they often vary significantly in terms of the skills they assess. Additionally, the objects and scenes included in these benchmarks are typically limited by what the simulator can provide. To more comprehensively evaluate VLA models, there is a need for benchmarks that cover a diverse set of skills based on realistic simulators. For high-level task planner VLAs, many benchmarks claim to measure planning capability, often in the form of question-answering tasks. However, it is more desirable to evaluate the high-level task planner together with the low-level control policy to execute long-horizon tasks and measure success rate, rather than relying on isolated measures of the planner alone. This approach provides a more holistic assessment of the VLA system’s capabilities.
Safety Considerations. Safety is paramount in robotics, as robots interact directly with the real world. Ensuring the safety of robotic systems requires the integration of real-world commonsense and complex reasoning into their development and deployment processes. This involves the incorporation of robust safety mechanisms, risk assessment frameworks, and human-robot interaction protocols. Interpretability and expandability of VLA decision-making processes are also crucial for enhancing robot safety through error diagnosis and troubleshooting.
Ethical and Societal Implications. The deployment of robots has always raised various ethical, legal, and societal concerns. These include risks related to privacy, safety, job displacement, bias in decision-making, and the impact on social norms and human relationships. Effective regulation plays a crucial role in promoting the ethical use of robots.
VI Conclusion
VLA policy holds immense promise for enabling embodied AI to interact effectively with the world around them. Recent advances have demonstrated the capabilities of these models in completing complex tasks under diverse conditions. However, significant challenges remain in terms of generalization, efficiency, and safety. Further research is needed to address these challenges and pave the way for the widespread adoption of VLA-powered robots in real-world applications.

### v4

@@
 Abstract
-Deep learning has demonstrated remarkable success across many domains, including computer vision, natural language processing, and reinforcement learning.
-Representative artificial neural networks in these fields span convolutional neural networks, Transformers, and deep Q-networks.
-Built upon unimodal neural networks, numerous multi-modal models have been introduced to address a range of tasks such as visual question answering, image captioning, and speech recognition.
-The rise of instruction-following robotic policies in embodied AI has spurred the development of a novel category of multi-modal models known as vision-language-action models (VLAs).
-Their multi-modality capability has become a foundational element in robot learning.
-Various methods have been proposed to enhance traits such as versatility, dexterity, and generalizability.
-Some models focus on refining specific components.
-Others aim to develop control policies adept at predicting low-level actions.
-Certain VLAs serve as high-level task planners capable of decomposing long-horizon tasks into executable subtasks.
-Over the past few years, a myriad of VLAs have emerged, reflecting the rapid advancement of embodied AI.
-Therefore, it is imperative to capture the evolving landscape through a comprehensive survey.
+Embodied AI is widely recognized as a key element of artificial general intelligence because it involves controlling embodied agents to perform tasks in the physical world.
+Building on the success of large language models and vision-language models, a new category of multimodal models—referred to as vision-language-action models (VLAs)—has emerged to address language-conditioned robotic tasks in embodied AI by leveraging their distinct ability to generate actions.
+In recent years, a myriad of VLAs have been developed, making it imperative to capture the rapidly evolving landscape through a comprehensive survey.
+To this end, we present the first survey on VLAs for embodied AI.
+This work provides a detailed taxonomy of VLAs, organized into three major lines of research.
+The first line focuses on individual components of VLAs.
+The second line is dedicated to developing control policies adept at predicting low-level actions.
+The third line comprises high-level task planners capable of decomposing long-horizon tasks into a sequence of subtasks, thereby guiding VLAs to follow more general user instructions.
+Furthermore, we provide an extensive summary of relevant resources, including datasets, simulators, and benchmarks.
+Finally, we discuss the challenges faced by VLAs and outline promising future directions in embodied AI.
+We have created a project associated with this survey, which is available at https://github.com/yueen-ma/Awesome-VLA.
 Index Terms:
-Embodied AI, Robotics, Multi-modality, Computer Vision, Natural Language Processing, Large Language Model
+Vision-Language-Action Model, Embodied AI, Robotics, Multi-modality, Large Language Model, World Model
 I Introduction
-[…]
-rect/.append style=rectangle, rounded corners=2pt, dir tree switch/.style args=at #1for tree= fit=rectangle, , where level=#1 for tree= folder, grow’=0, , delay=child anchor=north, , before typesetting nodes= for tree= content/.wrap value=##1, , if=isodd(n_children(”!r”)) for nodewalk/.wrap pgfmath arg=fake=r,n=##1calign with current edgeint((n_children(”!r”)+1)/2), , , ,
-{forest}
-[…]
-Figure 1: The taxonomy of Vision-Language-Action models.
-Vision-language-action models (VLAs) are a class of multi-modal models within the field of embodied AI, designed to process information from vision, language, and action modalities.
-Unlike conversational AI such as ChatGPT [1], embodied AI requires controlling physical embodiments that interact with the environment.
-Consequently, robotics is the most prominent domain of embodied AI.
-In language-conditioned robotic tasks, the policy must possess the capability to understand language instructions, visually perceive the environment, and generate appropriate actions, necessitating the multi-modal capabilities of VLAs.
+Vision-language-action models (VLAs) are a class of multimodal models within the field of embodied AI, designed to process information from vision, language, and action modalities.
+Unlike conversational AI such as ChatGPT [1], embodied AI requires controlling physical embodiments that interact with the environment and robotics is the most prominent domain of embodied AI.
+In language-conditioned robotic tasks, the policy must possess the capability to understand language instructions, visually perceive the environment, and generate appropriate actions, necessitating the multimodal capabilities of VLAs.
 The term was recently coined by RT-2 [2].
 Compared to earlier deep reinforcement learning approaches, VLAs offer superior versatility, dexterity, and generalizability in complex environments.
-As a result, they are well-suited not only for controlled settings like factories but also for everyday tasks in household environments [3].
+As a result, they are well-suited not only for controlled settings like factories but also for everyday tasks in household environments.
 Early developments in deep learning primarily consist of unimodal models.
-In computer vision (CV), models like AlexNet [4] showcased the potential of artificial neural networks (ANNs) [5].
-Recurrent neural networks (RNNs) [6] laid the groundwork for numerous natural language processing (NLP) models but have seen a transition in recent years, with Transformers [7] taking precedence.
-The Deep Q-network demonstrated that ANNs can successfully tackle reinforcement learning problems.
-Leveraging advancements of unimodal models across diverse machine learning fields, multi-modal models have evolved to be powerful enough to tackle a variety of tasks [8], such as visual question answering, image captioning, speech recognition, etc.
-Conventional robot policies based on reinforcement learning mostly concentrate on a restricted set of tasks, typically within controlled environments like factories and laboratories.
-For instance, [9] trains a policy specifically for grasping items.
-However, there is a growing need for more versatile multi-task policies, akin to recent developments in large language models (LLMs) [10, 1] and vision-language models (VLMs) [11].
-Developing a multi-task policy is more challenging as it requires learning a broader set of skills and adapting to dynamic and uncertain environments.
-Additionally, task specification adds another layer of complexity.
-Some approaches use a one-hot vector to select tasks [12], but they are limited by the number of tasks in the training set.
-Built upon the success of pretrained vision foundation models, LLMs, and VLMs, Vision-Language-Action models have demonstrated their competency in addressing these challenges.
-Pretrained visual representations from state-of-the-art vision encoders assist VLAs in perceiving complex environments [13, 14, 15], providing more precise estimations such as object class, object pose, and object geometry.
-With the increasing power of language models [10, 2], task specification based on language instructions becomes a viable option.
-Numerous ways of integrating vision models and language models have been explored by foundation VLMs, including BLIP-2 [16], Flamingo [11], etc.
-These innovations from different fields empower VLAs to address the challenges of embodied AI.
-Different VLAs exhibit distinct emphases, as depicted by the taxonomy in Figure 1.
-Some works strive to enhance specific constituents of VLAs, such as pretrained visual representations, dynamics learning, and world models.
-Meanwhile, a substantial body of work is dedicated to robotic control policy.
-In this category, language instructions are input into the control policy, which generates primitive actions based on the environment.
-In contrast, another category of VLAs functions as high-level task planners that abstract away low-level controls.
-Instead, these models focus on decomposing long-horizon robotic tasks into subtasks.
-These subtasks can then be completed by control policies one by one, eventually fulfilling the overall task.
+In computer vision (CV), models like AlexNet [3] showcased the potential of artifical neural networks.
+Recurrent neural networks laid the groundwork for numerous natural language processing (NLP) models but have seen a transition in recent years, with Transformers [4] taking precedence.
+The Deep Q-network [5] demonstrated that ANNs can successfully tackle reinforcement learning problems.
+Leveraging advancements of unimodal models across diverse machine learning fields, multimodal models now address a wide range of tasks, such as visual question answering, image captioning, and text-to-video generation.
+Conventional robot policies based on reinforcement learning have largely focused on addressing a limited set of tasks within controlled environments, such as item grasping [6].
+However, there is a growing need for more versatile multi-task policies, akin to recent trends in large language models (LLMs) and vision-language models (VLMs).
+Developing a multi-task policy is challenging, as it requires learning a broader set of skills and adapting to diverse environments.
+Furthermore, specifying tasks via language instructions offers a more intuitive user-robot interface, which necessitates the development of language-conditioned robot policies.
+Figure 1: The general architecture of vision-language-action models. Important related components are shown in the dashed-line boxes.
+Built upon the success of large VLMs, vision-language-action models have demonstrated their potential in addressing these challenges, as illustrated in Figure 1.
+Similar to VLMs, VLAs utilize vision foundation models as vision encoders to obtain pretrained visual representations of the current environmental state, such as object class, pose, and geometry.
+VLAs encode instructions using the token embeddings of their LLMs and employ various strategies to align vision and language embeddings, including approaches like BLIP-2 [7] and LLaVA [8].
+By finetuning on robot data, the LLM can function as a decoder to predict actions and perform language-conditioned robotic tasks.
+These cross-disciplinary innovations drive the advancement of VLAs in embodied AI—a critical building block for artificial general intelligence (AGI).
+(a) Concepts
+(b) Timelines
+Figure 2: (a) A Venn diagram that outlines the main concepts in embodied AI discussed in this paper. (b) Timelines that trace the evolution from unimodal models to vision-language-action models.
+VLAs are closely related to three lines of work, as depicted by the timelines in Figure 2b and the taxonomy in Figure 3.
+Some approaches focus on individual components of VLAs (§III-A), such as pretrained visual representations, dynamics learning, world models, and reasoning.
+Meanwhile, a substantial body of research is dedicated to low-level control policies (§III-B).
+In this category, language instructions and visual perceptions are fed into the control policy, which then generates low-level actions—such as translation and rotation—thereby rendering VLAs an ideal choice for control policies.
+In contrast, another category of models serves as high-level task planners responsible for task decomposition (§IV).
+These models break down long-horizon tasks into a sequence of subtasks that, in turn, guide VLAs toward the overall goal, as illustrated in Figure 4.
+Most current robotic systems adopt such a hierarchical framework [9, 10, 11] because the high-level task planner can leverage models with high capacity while the low-level control policy can focus on speed and precision, similar to hierarchical reinforcement learning.
+To provide a more comprehensive overview of current progress in embodied AI, we propose a generalized definition of “VLA,” as illustrated by the Venn diagram in Figure 2a.
+We define a VLA as any model capable of processing multimodal inputs from vision and language to produce robot actions that accomplish embodied tasks, typically following the architecture in Figure 1.
+The original concept of VLA referred to a model that adapts VLMs to robotic tasks [2].
+Analogous to the distinction between LLMs and more general language models, we designate the original VLAs as “Large VLAs” (LVLAs) because they are based on LLMs or large VLMs.
 Related Work.
-Although surveys for VLAs are currently lacking, existing surveys in related domains provide valuable insights for VLA research.
-In the field of computer vision, surveys cover a wide range of visual models, ranging from convolutional neural networks [17] to Transformers [18].
-Natural language processing models are comprehensively summarized in surveys [19, 20].
-In-depth reviews of reinforcement learning can be found in surveys [21, 22, 23].
-Surveys on graph neural networks are also available [24].
-Furthermore, existing surveys comparing vision-language models offer inspiration for VLAs [8, 25, 26, 27].
-Additionally, a survey on earlier works of embodied AI exists [28].
-Contributions. This survey stands as the first thorough review of the emerging vision-language-action model in the field of embodied AI.
-  • Comprehensive Review. We present a thorough review of emerging VLA models in embodied AI, covering various aspects including architectures, training objectives, and robotic tasks.
+To the best of our knowledge, this survey is the first to review the recent progress of VLA models, a rapidly emerging research area.
+Previous surveys have investigated other facets of embodied AI.
+[12] comprehensively summarized foundation models in robotics up to 2023, while [13] focused on LLMs in robotics.
+[14] examined more recent vision, language, and robotic foundation models for general-purpose robots.
+[15] concentrates on real-world robot applications.
+In contrast, our work emphasizes VLA models, thereby complementing and extending the existing literature on embodied AI.
+Contributions. To the best of our knowledge, this paper is the first comprehensive survey of emerging vision-language-action (VLA) models in the field of embodied AI.
+  • Comprehensive Review. We present a thorough review of emerging VLA models in embodied AI, covering various aspects including components, architectures, training objectives, robotic tasks, etc.
   • Taxonomy.
-We introduce a taxonomy of the hierarchical structure in current robot systems, comprising three main components: pretraining, control policy, and task planner.
-Pretraining techniques aim to enhance specific aspects of VLAs, such as the vision encoder or dynamics model.
-Low-level control policies execute low-level actions based on specified language commands and the perceived environment.
-High-level task planner decomposes long-horizon tasks into subtasks executable by control policies.
-  • Abundant Resources.
-We provide an overview of necessary resources for training and evaluating VLA models.
-We investigate recently introduced datasets and simulators by comparing their key characteristics.
-Additionally, we include widely-adopted benchmarks for tasks such as robot control and embodied reasoning.
-  • Future Directions.
-We outline current challenges and promising future opportunities in the field, such as addressing data scarcity, enhancing robot dexterity, achieving generalization across different tasks, environments, and embodiments, as well as improving robot safety.
-Paper Organization.
-§II-A provides an overview of representative developments and milestones in unimodal models.
-Because vision-language models are closely related to vision-language-action models, recent advancements in vision-language models are compared in §II-B.
-§III explores various types of vision-language-action models.
-§IV summarizes recent datasets and benchmarks for embodied AI.
-Challenges and future directions are included in §V.
+We introduce a taxonomy of VLAs based on the hierarchical framework of current robotic systems, comprising two hierarchies: a low-level control policy and a high-level task planner.
+Control policies execute low-level actions based on specified language commands and the perceived environment.
+Task planners provide guidance to control policies by decomposing long-horizon tasks into subtasks.
+We also discuss various essential components of VLAs.
+  • Resources.
+We summarize the necessary resources for training and evaluating VLA models, including recent datasets and benchmarks in real-world or simulated environments.
+We also discuss various approaches to address current challenges, such as data scarcity and inconsistency.
+  • Future Directions. We outline current challenges and promising future opportunities in the field, such as safety, foundation models, and real-world deployment.
 II Background
-Figure 2: A brief timeline traces the evolution from unimodal models to multimodal models, laying the groundwork for the introduction of VLA models.
+[…]
+[…]
+Figure 3: The taxonomy of VLA models. The organization of this survey follows this taxonomy.
 II-A Unimodal Models
 Vision-language-action models integrate three modalities, often relying on existing unimodal models.
-The transition from convolutional neural networks (e.g., ResNet [29]) to visual Transformers (e.g., ViT [30], SAM [31]) in computer vision has facilitated the development of more general vision models.
+The transition from convolutional neural networks (e.g., ResNet [16]) to visual Transformers (e.g., ViT [17], SAM [18]) in computer vision has facilitated the development of vision foundation models (VFMs).
 In natural language processing, the evolution from recurrent neural networks (e.g., LSTM [19], GRU [20]) to Transformers [4] initially led to the pretrain-finetune paradigm (e.g., BERT [21], ChatGPT [1]), followed by the recent success of prompt tuning driven by large language models.
@@
 II-B Vision-Language Models
-Vision-language tasks, encompassing image captioning [38], visual question answer [39], visual grounding [40], require the fusion of computer vision and natural language processing models.
-Early attempts like Show and Tell [41] capitalized on the success of early CNNs and RNNs.
-The introduction of high-capacity language models such as BERT [33] and GPT [42] ushered in a new era of Transformer-based VLMs, which have evolved in three major directions: self-supervised pretraining, contrastive pretraining, and multi-modal LLMs.
-One of the pioneering self-supervised pretraining methods is ViLBERT [43], while CLIP [44] popularized contrastive pretraining approaches for multi-modal alignment.
-The recent emergence of large language models has driven the development of multi-modal LLMs (MLLMs) or large multi-modal models (LMMs), which achieve state-of-the-art performance on multi-modal instruction-following tasks.
-Representative MLLMs include Flamingo [11], BLIP-2 [16], LLaVA [45].
-VLMs share a close relationship with VLAs, as the multi-modal architectures of VLMs can be readily adopted for VLAs.
-For example, attaching an action decoder to VLMs can transform them into VLAs for low-level control [46].
-VLMs can also serve as high-level task planners if they possess sufficient reasoning abilities.
+Vision-language tasks, encompassing image captioning [25], visual question answering [26], visual grounding [27], require the fusion of computer vision and natural language processing models.
+Early efforts, such as Show and Tell [28], leveraged the success of early CNNs and RNNs.
+The advent of high-capacity language models, including BERT [21] and GPT [29], ushered in a new era of Transformer-based VLMs.
+One of the pioneering self-supervised pretraining methods is ViLBERT [30], while CLIP [31] popularized contrastive pretraining approaches for multimodal alignment.
+The recent emergence of large language models has driven the development of multimodal LLMs (MLLMs) or large multimodal models (LMMs), which achieve state-of-the-art performance on multimodal instruction-following tasks.
+Representative MLLMs include Flamingo [32], BLIP-2 [7], and LLaVA [8].
+VLMs share a close relationship with VLAs, as the multimodal architectures of VLMs can be readily adopted for VLAs.
+Additionally, VLMs can function as high-level task planners if they possess sufficient reasoning capabilities.
 II-C Embodied AI & Robot Learning
@@
 Robot learning is also usually framed as reinforcement learning problems, represented as a Markov Decision Process (MDP) consisting of states ($s$), actions ($a$), and rewards ($r$).
-A MDP trajectory is denoted as $\tau=(s_{t=1},a_{t=1},r_{t=1},\dots,s_{t=T},a_{t=T},r_{t=T})$.
+A MDP trajectory is denoted as $\tau=(s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T})$.
 In certain scenarios, robotic tasks may also be viewed as partially-observable Markov Decision Processes (POMDPs) due to incomplete observations.
-The primary objective of reinforcement learning is to train a policy capable of generating optimal actions for the current state $\pi_{\theta}(a_{t}|p,s_{\leq t},a_{<t})$.
+The primary objective of reinforcement learning is to train a policy capable of generating optimal actions for the current state $\pi(a_{t}|s_{\leq t},a_{<t})$.
 Various methods such as TD learning, policy gradients, etc., can be employed to achieve this.
-However, in cases where defining the reward function proves challenging, imitation learning is utilized to directly model the action distribution within trajectories devoid of rewards $\tau=(s_{t=1},a_{t=1},\dots,s_{t=T},a_{t=T})$.
-Furthermore, many multi-task robot models employ language as instructions to determine which task or skill to execute, thereby engaging in language-conditioned reinforcement learning.
+However, in cases where defining the reward function proves challenging, imitation learning is utilized to directly model the action distribution within trajectories devoid of rewards $\tau=(s_{1},a_{1},\dots,s_{T},a_{T})$.
+Furthermore, many multi-task robot models employ language as instructions $p$ to determine which task or skill to execute, leading to the development of language-conditioned robot policies $\pi(a_{t}|p,s_{\leq t},a_{<t})$.
 III Vision-Language-Action Models
-Vision-language-action models (VLAs) are the models that handle multi-modal inputs of vision and language and output robot actions to complete embodied tasks.
-They serve as the cornerstone for instruction following robotic policies in the field of embodied AI.
-These models rely on robust vision encoders, language encoders, and action decoders.
-Consequently, some works focus on improving the constituent components of VLA models (Sec III-A).
-Others concentrate on refining low-level control policies, adept at handling simple instructions and executing primitive actions (Sec III-B).
-Additionally, certain VLAs abstract away from low-level control and primarily aim to solve task decomposition (Sec III-C).
-Thus, the combination of low-level control policy and high-level task planner can be treated as a hierarchical policy, as illustrated in Figure 3.
-III-A Constituents of VLA
-A variety of technologies serve as integral components of VLA models.
-Pioneering works such as Decision Transformer and Trajectory Transformer have paved the way for using Transformers in reinforcement learning (§III-A1), facilitating the development of scalable and high-capacity policies.
-The vision encoder is also crucial, as it must effectively encode the state and provide sufficient information about the environment to VLAs.
-Several works aim to pretrain vision encoders to obtain improved pretrained visual representations (PVRs) (§III-A2).
-Other research emphasizes learning the dynamics of the environment (§III-A3), enabling VLAs to understand the consequences of their actions and make more informed decisions.
-An extension of dynamics learning involves developing a world model (§III-A4) that can provide additional information, such as reward signals.
-This approach allows the model to sample imagined rollouts, reducing the need for interaction with the actual environment and improving data efficiency.
-Additionally, by leveraging internet-scale commonsense knowledge from large language models, LLM-induced world models (§III-A5) have been gaining popularity.
-III-A1 Reinforcement Learning Transformer
-Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them conducive to Transformer models.
-Pioneering efforts in this direction include Decision Transformer [48] and Trajectory Transformer [49].
-Decision Transformer focuses on learning a policy, i.e., modeling the actions.
-Trajectory Transformer shares commonalities with Decision Transformer but distinguishes itself by extending its modeling objectives to include both states and returns within the trajectory.
-These innovative approaches showcase the applicability of Transformer architectures to the intricacies of RL sequence modeling.
-Gato [50] further extends this paradigm to a multi-modal, multi-task, multi-embodiment setting.
-Figure 3: Illustration of a hierarchical robot policy comprising a high-level task planner and a low-level control policy.
-The high-level task planner generates a plan based on the user instruction, which is then executed step by step by the low-level control policy.
+III-A Components of VLA
+A body of work focuses on individual components of VLA models, drawing on successes in CV, NLP, and RL. We introduce them through the lens of embodied AI.
+III-A1 Reinforcement Learning
+Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them well-suited for Transformer models.
+These RL Transformers have become foundational to many recent VLA models.
+Pioneering efforts in this direction include Decision Transformer (DT) [34] and Trajectory Transformer (TT) [35].
+Gato [36] further extends this paradigm to a multi-modal/task/embodiment setting.
+Reinforcement learning from human feedback (RLHF) has become an important training ingredient for LLMs.
+SEED [37] utilizes RLHF alongside skill-based reinforcement learning to address the sparse reward issue of long-horizon tasks in robot learning.
+Reflexion [38] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
 III-A2 Pretrained Visual Representations
-The effectiveness of the vision encoder directly influences the performance of the policy, as it supplies crucial information regarding object categories, positions, and environmental affordance.
-Consequently, numerous methods are devoted to pretraining the vision encoder to improve the quality of PVRs.
+The effectiveness of the vision encoder directly influences the performance of VLAs, as it provides crucial information regarding the current state $s_{t}$, such as object categories, positions, and affordance.
+Consequently, numerous methods are devoted to improving the quality of pretrained visual representations (PVRs).
 Their technical details are compared in Table I.
-Although Contrastive Language-Image Pre-training (CLIP) [44] was not initially designed for robotics tasks, it has gained widespread adoption as a vision encoder in robotic models, such as CLIPort [13], EmbCLIP [51], and CoW [52].
+CLIP [31] has gained widespread adoption as a vision encoder in robotic models.
 The training objective of CLIP is to identify the correct text-image pair among all possible combinations in a given batch.
-By training to enhance alignment between the vision encoder and the language encoder, CLIP proves particularly effective in tasks where text instructions are provided as input.
-CLIP undergoes training on the WebImageText dataset (WIT), comprising 400 million image-text pairs.
+CLIP undergoes training on the WIT dataset, comprising 400 million image-text pairs.
 This large-scale training allows CLIP to develop a rich understanding of the relationship between visual and textual information.
-Notably, CLIP conducts a comprehensive comparison of various vision encoders, exploring different configurations of ResNet [29] and ViT [30].
-This analysis provides valuable insights into the trade-off between accuracy and efficiency.
-Reusable Representations for Robotic Manipulation (R3M) [53] proposes two main pretraning objectives: time contrastive learning and video-language alignment.
+R3M [39] proposes two main pretraining objectives: time contrastive learning and video-language alignment.
 The objective of time contrastive learning is to minimize the distance between video frames that are temporally close while simultaneously increasing the separation between frames that are temporally distant.
-This objective aims to create PVRs that capture the temporal relationships within the video sequence.
+This objective aims to create PVRs that capture the temporal relationship within the video sequence.
 On the other hand, video-language alignment is to learn whether a video corresponds to a language instruction.
 This objective enriches the semantic relevance embedded in the PVRs.
-Masked Visual Pre-training (MVP) [14] adopts the methodology of Masked Autoencoder (MAE) from computer vision.
+VIP [40] also capitalizes on video temporal relationship.
+MVP [41] adopts masked autoencoder (MAE) from computer vision to robotic datasets.
 The MAE involves masking out a portion of input patches to a ViT model and training it to reconstruct the corrupted patches.
 This approach closely resembles the masked language modeling technique used in BERT [21] and falls within the purview of self-supervised training.
-MVP extends this MAE objective across diverse robotic datasets, demonstrating that the pre-trained vision encoder significantly enhances performance in downstream manipulation tasks.
-Value-Implicit Pre-training (VIP) [54] capitalizes on the temporal sequences present in videos, distinguishing itself from R3M [53] by attracting both the initial and target frames while simultaneously repelling successive frames.
-This objective aims to capture long-range temporal relationships and uphold local smoothness.
-While their own experiments demonstrate superior performance compared to R3M in specific tasks, subsequent research presents conflicting findings through more comprehensive evaluations [15].
-Visual Cortex (VC-1) [15] conducts an in-depth examination of prior PVRs and introduces an enhanced PVR model by systematically exploring optimal ViT configurations across diverse datasets.
-Additionally, they perform a comprehensive comparative analysis of their model against previous methods on various manipulation and navigation datasets, shedding light on the critical factors that contribute to the improvement of PVR.
-Another study [55] also compares previous PVRs obtained under supervised learning or self-supervised learning.
-Voltron [56] introduces a novel pretraining objective by incorporating language conditioning and language generation into the masked autoencoding (MAE) objective.
-Employing an encoder-decoder Transformer structure [7], the pretraining alternates between language-conditioned masked image reconstruction and language generation from masked images.
-This enhances the alignment between language and vision modalities, consequently resulting in a significantly improved success rate on language-conditioned imitation tasks.
-RPT [57] undergoes pretraining with a focus not only on reconstructing visual inputs and robotic actions but also on proprioceptive states.
-In evaluating three distinct masking schemes, it was noted that token masking, in particular, yielded the most significant improvement in the model’s performance.
-GR-1 [58] introduces a novel pretraining task of video prediction tailored for a GPT-style model.
-This video prediction objective is also subsequently employed during the finetuning phase, specifically utilizing robot data.
-The rationale behind this approach is that the ability to anticipate future frames contributes to more accurate action prediction.
-The experimental results, in the domain of robot manipulation, provide empirical support for their assertion.
-SpawnNet [59] utilizes a two-stream architecture incorporating adapter layers to fuse features from both a pretrained vision encoder and features learned from scratch.
-This innovative approach eliminates the necessity of training the pretrained vision encoders while surpassing the performance of parameter-efficient finetuning (PEFT) methods, as evidenced by experimental results in robot manipulation tasks.
+RPT [42] undergoes pretraining with a focus not only on reconstructing visual inputs but also on robotic actions and proprioceptive states.
 Table I: Pretrained visual representations.
@@
 L: Language.
-Net: Backbone network.
 CL: contrastive learning.
-MAE: masked auto-encoding.
 TFM: Transformer.
@@
 Mani/Navi: manipulation/navigation.
-[SC]: self-collect data.
 For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc.
-$\mathcal{S}(\cdot)$ is the similarity measurement.
-(Ego-Data): Ego4D [60], Epic Kitchens [61], Something-Something-v2 [62], 100DOH [63].
-Model V-Encoder Net Objective Formula Notation Dataset Robotic Tasks
-CLIP [44] ViT-B [30] - VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair WIT [SC] (EmbCLIP [51])
-R3M [53] ResNet-50 [29] - Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Ego4D Sim-Mani: Meta-World, Franka Kitchen, Adroit
+$\mathcal{S}(\cdot)$ is some similarity measurement.
+Model V-Encoder Type Objective Notation Robotic Tasks
+CLIP [31] ViT-B [17] VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair (Used by CLIPort [43], EmbCLIP [44], and CoW [45])
+R3M [39] ResNet-50 [16] Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Sim-Mani: Meta-World, Franka Kitchen, Adroit
 MVP [41] ViT-B
 ViT-L
-- MAE [64] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch (Ego-Data) Real-Mani (xArm7): pick, reach block, push cube, close fridge
-VIP [54] ResNet-50 - Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{0},x_{j})\right)}{\exp\left(\mathcal{S}(x_{0},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{j})\right)}$ $x_{i}$: $i$-th video frame, $0<i<j$ Ego4D Real-Mani (Franka): pick and place, close drawer, push bottle, fold towel
-VC-1 [15] ViT-L - MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP, including most previous datasets: (Ego-Data), Real Estate 10K [65], OpenHouse24 [SC]) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0;
-Sim-Navi: Habitat
-Voltron [56] TFM [7] MAP [66] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
+MAE [46] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch Real-Mani (xArm7): pick, reach block, push cube, close fridge
+VIP [40] ResNet-50 Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{0},x_{j})\right)}{\exp\left(\mathcal{S}(x_{0},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{j})\right)}$ $x_{i}$: $i$-th video frame, $0<i<j$ Real-Mani (Franka): pick and place, close drawer, push bottle, fold towel
+VC-1 [47] ViT-L MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP.) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0; Sim-Navi: Habitat
+Voltron [48] TFM [4] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
 $\sum_{i=1}^{N^{\prime}}-\log P\left(y_{i}\mid x,y_{<i}\right)$
@@
 $y$: language instruction
-Something-Something-v2 Sim-Mani: Franka Kitchen;
+Sim-Mani: Franka Kitchen;
 Real-Mani (Franka): custom study desk
-RPT [57] ViT [30, 14] TFM MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x_{i}$: token from any modality; $y,z$: other modalities [SC] Real-Mani (Franka): stack, pick, pick from bin
-GR-1 [58] ViT-MAE TFM MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: image patch; $y$: language instruction Ego4D Sim-Mani: CALVIN
-[…]
-Table II: Dynamics learning methods for VLAs. $f(\cdot)$ is the dynamic model. Fwd: forward. Inv: Inverse.
-Model V-Encoder Objective Formula Notation Robotic Tasks
-Vi-PRoM [71] ResNet [29] Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$ predicts the frame index in original seq.
-$x$, given a shuffled seq.
+RPT [42] ViT [17, 41] MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x,y,z$: three distinct modalities Real-Mani (Franka): stack, pick, pick from bin
+DINOv2 [49] ResNet, ViT Self-distillation $\sum_{x}\sum_{x^{\prime}\neq x}H(P_{t}(x),P_{x}(x^{\prime}))$ $x,x^{\prime}$: image views; $H()$: cross-entropy; $P_{t},P_{s}$: teacher, student (Used by OpenVLA [50], ReKep [51])
+I-JEPA [52] ViT JEPA $\frac{1}{M}\sum_{i=1}^{M}\sum_{j}\left\lVert x_{j}(i)-y_{j}(i)\right\rVert_{2}^{2}$ $x_{j}(i)$: $j$-th masked patch embedding of block $i$; $y_{j}(i)$: unmasked patch embedding -
+Theia [53] ViT-T/S/B Distillation (Distillation of vision foundation models, including ViT, CLIP, SAM, DINOv2, and Depth-Anything.) Sim-Mani&Navi: CortexBench [47]; Real-Mani: pick, place, open door/drawer
+Voltron [48] introduces a pretraining objective by incorporating language conditioning and language generation into the MAE objective.
+Employing an encoder-decoder Transformer structure, the pretraining alternates between language-conditioned masked image reconstruction and language generation from masked images.
+This enhances the alignment between language and vision modalities.
+VC-1 [47] conducts an in-depth examination of prior PVRs and introduces an enhanced PVR model by systematically exploring optimal ViT configurations across diverse datasets.
+Additionally, they perform a comprehensive comparative analysis of their model against previous methods on various manipulation and navigation datasets, shedding light on the critical factors that contribute to the improvement of PVR.
+Another study [54] also compares previous PVRs obtained under supervised learning or self-supervised learning.
+DINOv2 [49] proposes a new self-supervised training paradigm for PVRs that achieves performance beyond that of MAE.
+It employs a self-distillation framework in which the teacher and student networks receive different views of the same image and match their encoded representations.
+The student network is updated using SGD, while the teacher network is maintained as the EMA of the student network.
+I-JEPA [52] is motivated by the joint-embedding predictive architectures proposed by [55].
+It constructs a “primitive” internal world model by comparing the embeddings of patches.
+Unlike DINO, which uses cropped images, I-JEPA employs masked patches.
+Moreover, it differs from MAE because it is a non-generative approach.
+Theia [53] proposes distilling various vision foundation models into a single model.
+By fusing their information from segmentation, depth, semantics, etc., it outperforms previous PVRs while requiring less data and a smaller model size.
+III-A3 Video Representations
+Videos are simple sequences of images and can be represented by concatenating the usual PVRs of each frame.
+However, their multi-view nature enables a variety of unique representation techniques beyond those mentioned above, such as time contrastive learning and MAE.
+NeRF can be extracted from videos and contains rich 3D information for robot learning, as exemplified by F3RM [56] and 3D-LLM [57].
+The recent 3D Gaussian Splatting (3D-GS) method has been gaining popularity in computer vision as well as in embodied AI.
+3D Gaussians can be deformed through physics-based simulation for generative dynamics, as proposed in PhysGaussian [58], or they can serve as 3D representations for VLMs, as enabled by UniGS [59].
+Additionally, many videos contain audio, which can provide important cues for robot policies [60].
+III-A4 Dynamics Learning
+Dynamics learning encompasses objectives aimed at endowing the model $f(\cdot)$ with an understanding of forward or inverse dynamics.
+Forward dynamics involves predicting the subsequent state resulting from a given action, whereas inverse dynamics entails determining the action required to transition from a previous state to a known subsequent state:
+$\begin{split}\text{Forward dynamics: }&\hat{s}_{t+1}\leftarrow f_{\text{fwd}}(s_{t},a_{t}),\\ \text{Inverse dynamics: }&\hat{a}_{t}\leftarrow f_{\text{inv}}(s_{t},s_{t+1}).\end{split}$ (1)
+Some approaches also frame these objectives as reordering problems for shuffled state sequences.
+Some forward dynamics methods closely resemble the image or video prediction pretraining used in PVRs.
+We compare them in Table II.
+Table II: Dynamics learning methods for VLAs. $f(\cdot)$ is the dynamic model. Fwd, inv: forward & inverse dynamics.
+Model V-Encoder Type Objective Notation Robotic Tasks
+Vi-PRoM [61] ResNet Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$: predicts frame index in raw seq.
+$x$ given shuffled seq.
 $x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open and close drawer/door
@@
 VPT [68] ResNet Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim: Minecraft
-III-A3 Dynamics Learning
-Dynamics learning encompasses objectives aimed at imbuing the model with an understanding of forward or inverse dynamics.
-Forward dynamics involves predicting the subsequent state resulting from a given action, whereas inverse dynamics entails determining the action required to transition from a previous state to a known subsequent state.
-Some research approaches also frame these objectives as reordering problems for shuffled state sequences.
-While forward dynamics models closely relate to world models, this subsection focuses specifically on works that leverage dynamics learning as an auxiliary task to enhance the performance of the primary robot task.
-We compare them in Table II.
+GR-1 [69] ViT-MAE Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: CALVIN
 Vi-PRoM [61] presents three distinct pretraining objectives.
@@
 Notably, in contrast to preceding masked modeling approaches like BERT or MAE, MaskDP is applied directly to downstream tasks
-Perception-Action Causal Transformer (PACT) [76] introduces a pretraining objective aimed at modeling state-action transitions.
-PACT receives sequences of states and actions as input, and predicts each state and action token in an autoregressive manner.
+PACT [66] introduces a pretraining objective aimed at modeling state-action transitions.
+It receives sequences of states and actions as input, and predicts each state and action token autoregressively.
 This pretrained model serves as a dynamics model, which can then be finetuned for various downstream tasks such as localization, mapping, and navigation.
-Video Pretraining (VPT) [78] proposes a method that harnesses unlabeled internet data to pretrain a foundational model for the game of Minecraft.
+VPT [68] proposes a video pretraining method that harnesses unlabeled internet data to pretrain a foundational model for the game of Minecraft.
 The approach begins by training an inverse dynamics model using a limited amount of labeled data, which is then utilized to label internet videos.
@@
 As a result of this process, the model demonstrates human-level performance across a multitude of tasks.
-III-A4 Classical World Model
-Dreamer [79] employs three primary modules to construct a latent dynamics model: the representation model, responsible for encoding images into latent states; the transition model, which captures transitions between latent states; and the reward model, predicting the reward associated with a given state.
+GR-1 [69] introduces video prediction pretraining tailored for a GPT-style model.
+The ability to anticipate future frames aligns with forward dynamics learning, contributing to more accurate action prediction.
+III-A5 World Models
+A world model $P(\cdot)$ encodes commonsense knowledge about the world and predicts the future state for a given action [55]:
+$\begin{split}\hat{s}_{t+1}\sim P(\hat{s}_{t+1}|s_{t},a_{t}).\end{split}$ (2)
+It enables model-based control and planning for embodied agents, as they can search for an optimal action sequence in imaginary space before executing any real actions.
+Although forward dynamics learning also attempts to predict the next state, it is usually treated as a pretraining task or an auxiliary loss to enhance the RL-Transformer-based action decoder for the primary robotics tasks, rather than serving as a standalone module.
+Additionally, new embodied experiences can be sampled from visual world models that explicitly generate images or videos of future states.
+Dreamer [70] employs three primary modules to construct a latent dynamics model: a representation model, responsible for encoding images into latent states; a transition model, which captures transitions between latent states; and a reward model, predicting the reward associated with a given state.
 Under the actor-critic framework, Dreamer utilizes an action model and a value model to learn behavior through imagination by propagating analytic gradients through the learned dynamics.
@@
 DayDreamer [73] applies this method to physical robots performing real-world tasks.
-Masked World Model (MWM) [83] innovates by modifying the vision encoder of DreamerV2 to a hybrid composition of Convolutional Neural Network and Vision Transformer.
-In training this novel Convolutional-ViT vision encoder, MWM draws inspiration from the approach proposed in MAE [64].
-Through the incorporation of an auxiliary reward prediction loss, the resulting learned latent dynamics model exhibits a notable performance improvement across various visual robotic tasks.
-Iso-Dream [84] introduces two key enhancements to the Dreamer framework.
-The first enhancement focuses on optimizing inverse dynamics by decoupling controllable and noncontrollable dynamics.
-The second enhancement involves optimizing agent behavior based on the decoupled latent imaginations.
-This approach is particularly advantageous, as the noncontrollable state transition branch can be rolled out independently of actions, yielding benefits for long-horizon decision-making processes.
-Transformer-based World Model (TWM) [85] shares the same motivation as Dreamer but adopts a different approach by constructing a world model based on the Transformer-XL architecture [86].
-This Transformer-based world model trains a model-free agent in latent imagination by generating new trajectories.
-Additionally, TWM suggests a modification to the balanced KL divergence loss from DreamerV2 and introduces a novel thresholded entropy loss tailored for the advantage actor-critic framework.
 IRIS [74] employs a GPT-like autoregressive Transformer as the foundation of its world model, with a VQ-VAE serving as the vision encoder.
-Subsequently, a policy is trained using imagined trajectories, which are unrolled from a real observation by the world model, akin to TWM’s approach.
-SWIM [88] advocates for the use of human videos in training a world model due to the availability of large-scale human-centric data.
-However, acknowledging the substantial gap between human and robot data, SWIM addresses this disparity by grounding actions in visual affordance maps.
-This approach involves inferring target poses based on the affordance maps, facilitating an effective transfer of knowledge from human data to enhance robot control.
+Subsequently, a policy is trained using imagined trajectories, which are unrolled from a real observation by the world model.
+TWM [75] also investigates the application of Transformers in building world models.
+III-A6 LLM-induced World Models
+LLMs encompass a wealth of commonsense knowledge about the world, prompting many approaches to leverage that knowledge for improving VLAs.
+DECKARD [76] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs, specifically tailored for the task of item crafting in Minecraft.
+DECKARD iterates between two phases: in the Dream phase, it samples a subgoal guided by the AWM; in the Wake phase, DECKARD executes the subgoal and updates the AWM through interactions with the game.
+This guided approach enables DECKARD to achieve markedly faster item crafting compared to baselines lacking such guidance.
+LLM-DM [77] uses an LLM to construct world models in planning domain definition language (PDDL)—a capability that LLM+P [78] did not achieve, as its PDDL world model was hand-crafted.
+The LLM also acts as an interface, mediating between the generated PDDL model and corrective feedback from syntactic validators and human experts.
+Finally, the PDDL world model serves as a symbolic simulator, assisting the LLM planner in generating plans.
+RAP [79] repurposes an LLM to act as both a policy that predicts actions and a world model that provides the state transition distribution.
+Unlike previous chain-of-thought prompting methods, RAP incorporates Monte Carlo Tree Search (MCTS) to enable structured planning, allowing the LLM to build a reasoning tree incrementally.
+This reasoning strategy helps RAP find a high-reward path that balances exploration and exploitation.
+Tree-Planner [80] improves efficiency by prompting the LLM only once to generate diverse paths within an action tree.
+LLM-MCTS [81] builds upon RAP but extends the problem setting to POMDPs.
+As a world model, the LLM generates the initial belief of the current state; as a policy, it serves as a heuristic to guide action selection.
+By leveraging its commonsense knowledge, the LLM reduces the search space of MCTS, thereby improving search efficiency.
+III-A7 Visual World Models
+Unlike LLM-induced world models, which are in textual form, visual world models generate images, videos, or 3D scenes of future states—aligning more closely with the physical world.
+They can further be utilized to generate new trajectories.
+This direction has been gaining increasing attention since Sora demonstrated world-simulation capabilities [82], as investigated by a dedicated survey [83].
 Genie [84] introduces a new class of generative models, termed Generative Interactive Environments.
@@
 Consequently, Genie establishes a foundation world model.
-III-A5 LLM-induced World Model
-DECKARD [90] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs, specifically tailored for the task of item crafting in Minecraft.
-DECKARD iterates between two phases: in the Dream phase, it samples a subgoal guided by the AWM; in the Wake phase, DECKARD executes the subgoal and updates the AWM through interactions with the game.
-This guided approach, centered around AWMs, enables DECKARD to achieve markedly faster item crafting compared to baselines lacking such guidance.
-LLM-DM [91] uses an LLM to construct world models in planning domain definition language (PDDL)—a capability that LLM+P [92] did not achieve, as its PDDL world model was hand-crafted.
-The LLM also acts as an interface, mediating between the generated PDDL model and corrective feedback from syntactic validators and human experts.
-Finally, the PDDL world model serves as a symbolic simulator, assisting the LLM planner in generating plans.
-RAP [93] repurposes an LLM to act as both a policy that predicts actions and a world model that provides the state transition distribution.
-Unlike previous chain-of-thought prompting methods, RAP incorporates Monte Carlo Tree Search (MCTS) to enable principled planning, allowing the LLM to build a reasoning tree incrementally.
-This reasoning strategy helps RAP find a high-reward path that balances exploration and exploitation.
-Tree-Planner [94] improves efficiency by prompting the LLM only once to generate diverse paths within an action tree.
-LLM-MCTS [95] builds upon RAP but extends the problem setting to Partially Observable Markov Decision Processes (POMDPs).
-It proposes using the LLM as both a world model and a heuristic policy within an MCTS planning algorithm.
-As a world model, the LLM generates the initial belief of the current state; as a policy, it serves as a heuristic to guide action selection.
-By leveraging its commonsense knowledge, the LLM reduces the search space of MCTS, thereby improving search efficiency.
-E2WM [96] is an LLM training paradigm that focuses on collecting embodied experiences from world models using MCTS.
-Unlike prior works, E2WM treats simulators as world models, as they emulate real-world physical interactions, and MCTS is used solely for data collection.
-By fine-tuning LLMs with this embodied knowledge, the models gain enhanced embodied abilities, such as plan generation, activity recognition, and tracking.
 3D-VLA [85] proposes a 3D world model capable of goal generation.
@@
 The generated goal state can subsequently be used to guide robot control.
-III-A6 Reasoning
-Reasoning has become an important capability in LLMs, as pioneered by chain-of-thought (CoT) methods [98, 99].
-Recent OpenAI o1 models [100] further demonstrate its potential in building more powerful AI systems.
-A similar trend has also emerged in the field of embodied AI.
-EmbodiedGPT [101] uses CoT to enhance embodied planning, and ThinkBot [102] applies it to recover missing action descriptions.
+UniSim [86] builds a generative model based on real-world interaction videos.
+It is capable of simulating visual outcomes for both high-level and low-level actions, which can then be leveraged as new experiences for training embodied agents.
+E2WM [87] even treats existing simulators as world models to collect embodied experiences via MCTS.
+III-A8 Reasoning
+Reasoning has become a key capability of LLMs, as demonstrated by chain-of-thought (CoT) methods [88, 89].
+In embodied AI, researchers are exploring how to leverage CoT reasoning to refine the decision-making process.
+ThinkBot [90] applies it to recover missing action descriptions in sparse human instructions.
 RAT [91] combines retrieval-augmented generation and CoT to elicit context-aware reasoning, improving long-horizon generation.
-ReAct [104] interleaves verbal reasoning traces and actions to address diverse decision-making tasks.
+ReAct [92] interleaves verbal reasoning traces and actions to allow for diverse types of reasoning.
 Tree-Planner [80] employs a tree-of-thoughts approach.
-Text2Motion [105] uses symbolic and geometric reasoning to verify the feasibility of an LLM-generated task plan.
-Reflexion [106] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
+ECoT [93] re-trains OpenVLA [50] using reasoning chain datasets, enabling embodied CoT reasoning that further enhances its performance.
 Pros and Cons.
-Pretrained visual representations underscore the importance of the vision encoder, as visual observation plays a crucial role in perceiving the current state of the environment.
-Consequently, it sets an upper bound on the performance of the overall model.
-In VLAs, general vision models are pretrained using robot or human data to enhance their capabilities in tasks such as object detection, affordance map extraction, and even vision-language alignment, which are essential for robotic tasks.
-In contrast, dynamics learning focuses on understanding the transition between states.
-This involves not only mapping visual observations to good state representations but also comprehending how different actions lead to different states, and vice versa.
-Existing dynamics learning methods typically aim to capture the relationship between states and actions using simple masked modeling or reordering objectives.
-On the other hand, a world model aims to completely model the dynamics of the world, enabling the robot models to roll out states multiple steps into the future based on the current state, facilitating better prediction of optimal actions.
-Therefore, while a world model is more desired, it is also more challenging to achieve.
+PVRs
+Although time contrastive learning with videos and text-guided pretraining methods, such as CLIP, provide image-level information, they lack the pixel-level details offered by MAE-based self-supervised learning methods.
+Pixel-level information contains rich details—including segmentation masks, object positions, and depth estimation—that are generally more useful for robot manipulation tasks requiring high precision, as demonstrated by VC-1’s comparison [47].
+DINOv2 learns both pixel- and image-level features by combining masked image modeling with a momentum encoder and multi-crop augmentation, with benefits on downstream robotic tasks evidenced by OpenVLA [50].
+I-JEPA focuses on patches in the representation space and, as a result, captures low-level image features more effectively than view-invariance methods such as DINO.
+Theia distills various off-the-shelf vision foundation models into a single model that surpasses isolated individual models, as demonstrated by comprehensive evaluations against most existing PVRs in robot learning [53].
+Forward & inverse dynamics
+Forward dynamics learning is generally more challenging than inverse dynamics learning because predicting future states is more complex than predicting past actions.
+Consequently, the difficulty of forward dynamics often leads to greater performance improvements, as demonstrated in SMART.
+However, inverse dynamics models can be used to generate action labels for datasets that contain only states, such as raw robot manipulation videos.
+World Models & Reasoning
+Although both world models and reasoning methods can be applied to low-level control policies and high-level task planners, current approaches remain distinct.
+World models are predominantly used to interact with control policies because they excel at generating the immediate next state given low-level actions.
+In contrast, CoT-based reasoning methods focus on task planning since they express thought chains in text, making them well-suited for refining text-based task plans.
+Figure 4: Illustration of a hierarchical robot policy. The high-level task planner decomposes the user instruction into subtasks, which are then executed step by step by the low-level control policy.
 Table III: Low-level control policies.
@@
 $p$/$s$: prompt/state vision encoder.
-(SC): self-collect data.
-LMP: Latent Motor Plans [107].
+[SC]: self-collect data.
 MPC: model predictive control.
 Model V-Encoder L-Encoder VL-Fusion Action-Decoder Objective Training Dataset Robotic Tasks
-CLIPort [13] CLIP-ResNet50, Transporter-ResNet [108] CLIP-GPT Hadamard LingUNet [109] BC ($\textbf{SE}(2)$) (SC) Sim: Ravens [108]
-MCIL [110] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP + (SC) Sim: 3D Playroom environment
+CLIPort [43] CLIP-ResNet50, Transporter-ResNet CLIP-GPT Hadamard LingUNet BC ($\textbf{SE}(2)$) [SC] Sim: Ravens [94]
+MCIL [95] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP, [SC] Sim: 3D Playroom environment
 HULC [96],
 HULC++ [97]
-MCIL CNN Sentence-BERT [113] LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
+MCIL CNN Sentence-BERT LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
 Language costs [98] CLIP-ViT,
 UNet [99]
-CLIP-GPT Concat STORM [116] Max Likelihood Est. (cost map) (SC) Sim & Real (Franka): pick, place
-Interactive Language [117] ResNet [29] CLIP-GPT Xattn TFM BC (cont) Language-Table (SC) Sim & Real (xArm6): rearrangement
-Hiveformer [118] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data [119] Sim: RLBench
-PerAct [120] 3D CNN CLIP-GPT Concat PerceiverIO [121] 3D affordance RLBench data [119] Sim: RLBench; Real (Franka): put marker, place food, etc.
-Act3D CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
-RVT, RVT-2 [122, 123] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data [119], (SC) Sim: RLBench; Real (Franka): stack, press, place
-RoboPoint [124] ViT-L/14 Vicuna-V1.5 13B [125] Concat - 2D affordance (project to 3D) (SC) Real (Franka): pick, place
-Gato [50] ViT Sent.Piece [126] Concat TFM BC (cont & disc) (SC) Sim & Real (Sawyer): RGB-stacking
+CLIP-GPT Concat STORM Max Likelihood Est. (cost map) [SC] Sim & Real (Franka): pick, place
+Interactive Language [100] ResNet CLIP-GPT Xattn TFM BC (cont) Language-Table [SC] Sim & Real (xArm6): rearrangement
+Hiveformer [101] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data Sim: RLBench [102]
+PerAct [103] 3D CNN CLIP-GPT Concat PerceiverIO 3D affordance RLBench data Sim: RLBench; Real (Franka): put marker, place food, etc.
+Act3D [104] CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
+RVT, RVT-2 [105, 106] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data, [SC] Sim: RLBench; Real (Franka): stack, press, place
+RoboPoint [107] ViT-L/14 Vicuna-V1.5 13B [108] Concat - 2D affordance (project to 3D) [SC] Real (Franka): pick, place
+Gato [36] ViT Sent.Piece Concat TFM BC (cont & disc) [SC] Sim & Real (Sawyer): RGB-stacking
 RoboCat [109] VQ-GAN ($p,s$) [110] - - TFM BC, Observation prediction Self-improve Sim & Real (Sawyer, Franka, KUKA): stacking, building, lifting, insertion, removal
-VIMA [129] ViT, Mask R-CNN [74] T5 [10] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data (SC) Sim (Ravens): VIMA-Bench
-BC-Z [130] ResNet18 ($p,s$) [29] USE FiLM [131] MLP BC (cont) (SC) Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
-RT-1 [3] EfficientNet [132] USE FiLM TFM BC (disc) Kitchen (SC) Real (Everyday Robots): pick-place, move, knock
-MOO [129] OWL-ViT ($p$) [133], EfficientNet ($s$) USE FiLM TFM BC (disc) (SC) Real (Everyday Robots): pick, move near, knock, place upright, place into
-Q-Transformer [134] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1 data Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
-RT-Trajectory [135] EfficientNet - - TFM BC (disc) (SC) Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
-ACT [136] ResNet18 - - CVAE-TFM BC (cont, action chunking) (SC) with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
-MT-ACT [137] CNN N/A FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
-RT-2 [2] ViT-22B [138],
-ViT-4B [139]
-PaLI-X [140],
-PaLM-E
-Concat Symbol-tuning [141] BC (disc), Co-fine-tuning VQA + Kitchen Real: RT-1 evaluation tasks, Language-Table
-RT-H [142] ViT-22B [138] PaLI-X 55B [140] Concat Symbol-tuning BC (disc) Diverse + Kitchen Real: Kitchen, Diverse
-RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE (SC) Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
-OpenVLA [144] DINOv2 [145],
-SigLIP [146]
-LLaMA 2 [147] (Prismatic-7B VLM [148] Concat Symbol-tuning BC (disc) OXE, DROID [149] Real: BridgeV2, RT-1 evaluation tasks
-RoboFlamingo [46] CLIP-ViT-L/14 LLaMA [150], GPT-NeoX [151], MPT [152] Xattn LSTM, TFM, MLP BC (cont) CALVIN data Sim: CALVIN
+VIMA [111] ViT, Mask R-CNN [64] T5 [112] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data [SC] Sim (Ravens): VIMA-Bench
+BC-Z [113] ResNet18 ($p,s$) USE FiLM MLP BC (cont) [SC] Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
+RT-1 [114] EfficientNet [115] USE FiLM TFM BC (disc) RT-1-Kitchen [SC] Real (Everyday Robots): pick-place, move, knock
+MOO [111] OWL-ViT ($p$) [116], EfficientNet ($s$) USE FiLM TFM BC (disc) [SC] Real (Everyday Robots): pick, move near, knock, place upright, place into
+Q-Transformer [117] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1-Kitchen Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
+RT-Trajectory [118] EfficientNet - - TFM BC (disc) [SC] Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
+ACT [119] ResNet18 - - CVAE-TFM BC (cont, action chunking) [SC] with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
+MT-ACT [120] (RoboAgent) CNN - FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
+RoboFlamingo [121] CLIP-ViT-L/14 LLaMA [122], MPT [123] Xattn LSTM, TFM BC (cont) CALVIN data Sim: CALVIN
 RoboUniView [124] UVFormer (Model design follows RoboFlamingo) BC (cont) CALVIN data Sim: CALVIN
@@
 SAM [18]
-GPT-4 [157] Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
-UniPi [158] Imagen video [159] T5-XXL [10] Xattn CNN Inverse dynamics PF (SC), Ravens, BridgeV1 Sim: Painting Factory (PF) [160], Ravens
-Diffusion Policy [161] ResNet18 [29] - - UNet,
+GPT-4 Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
+UniPi [128] Imagen video [129] T5-XXL Xattn CNN Inverse dynamics PF [SC], Ravens, BridgeV1 Sim: Painting Factory (PF), Ravens
+Diffusion Policy [130] ResNet18 - - UNet,
 TFM
-DDPM [162] (SC) Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
-DP3 [163] DP3 Encoder (Model design follows Diffusion Policy) DDIM (SC) Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
-SUDD [164] ResNet18 [29] CLIP-GPT Concat UNet,
+DDPM [131] [SC] Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
+DP3 [132] DP3 Encoder (Model design follows Diffusion Policy) DDIM [SC] Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
+SUDD [133] ResNet18 CLIP-GPT Concat UNet,
 TFM
 DDPM Lang-Guided Data Gen Sim: MuJoCo; Real (UR5e): pick, place
-Octo [165] CNN T5-base [10] Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
-3D Diffuser Actor [166] (Model design follows Act3D) DDPM RLBench, CALVIN, (SC) Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
-MDT [167] CLIP-ViT-B/16 ($p$), Voltron / ResNet18 ($s$) CLIP-GPT Concat DiT [168] DDIM [169] CALVIN, LIBERO, Kitchen (SC) Sim: CALVIN, LIBERO [170]; Real (Franka): pick, push, open, close, etc.
+Octo [134] CNN T5-base Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
+3D Diffuser Actor [135] (Model design follows Act3D) DDPM RLBench, CALVIN, [SC] Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
+MDT [136] CLIP-ViT-B/16 ($p$), Voltron/ResNet18 ($s$) CLIP-GPT Concat DiT [137] DDIM [138] CALVIN, LIBERO, [SC] Sim: CALVIN, LIBERO; Real (Franka): pick, push, open, close, etc.
 RDT-1B [139] SigLIP T5-XXL Xattn DiT DDPM (Aggregated Datasets) Real (ALOHA dual-arm): wash, pour, fold, etc.
-III-B Low-level Control Policy
-Through the integration of an action decoder with perception modules, such as vision encoders and language encoders, a control policy network is formed to execute language instructions in both simulated and real-world environments.
-The diversity among control policy networks arises from the selection of individual modules and the integration strategy.
-This section explores various approaches to designing a low-level control policy.
-Based on the backbone of their action decoder, control policies can be broadly categorized into three types: non-Transformer (§III-B1), Transformer-based (§III-B2), and LLM-based (§III-B8).
-Other approaches focus on specific aspects of control policies, such as multi-modal instructions (§III-B3), 3D Vision (§III-B4), diffusion-based action generation (§III-B5), and motion planning (§III-B7).
-In Table III, we compare the architectural details of these control policies.
-III-B1 Early Non-Transformer Control Policies
+RT-2 [2] ViT-22B [140],
+ViT-4B [141]
+PaLI-X, PaLM-E Concat Symbol-tuning BC (disc), Co-fine-tuning VQA, RT-1-Kitchen Real: RT-1 evaluation tasks, Language-Table
+RT-H [142] ViT-22B [140] PaLI-X 55B Concat Symbol-tuning BC (disc) Diverse+Kitchen Real: Diverse+Kitchen eval tasks
+RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE [SC] Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
+OpenVLA [50] DINOv2, SigLIP [144] Prismatic-7B [145] Concat Symbol-tuning BC (disc) OXE, DROID [146] Real: BridgeV2, RT-1 evaluation tasks
+$\pi_{0}$ [147] SigLIP PaliGemma Concat Action expert Flow matching OXE, $\pi$-cross-embod. [SC] Real (general robot control): sweep, open, pack, etc.
+III-B Low-level Control Policies
+Through the integration of an action decoder with perception modules, such as vision encoders and language encoders, a VLA model $\pi_{\theta}$ with parameters $\theta$ is formed as a control policy to execute language instructions $p$:
+$\begin{split}\hat{a}_{t}\sim\pi_{\theta}(\hat{a}_{t}|p,s_{\leq t},a_{<t}).\end{split}$ (3)
+The diversity among control policy networks arises from individual modules and overall architectures.
+The most typical architecture is shown in Figure 1.
+This section explores various approaches to designing low-level control policies.
+Table III summarizes their technical details.
+III-B1 Non-Transformer Control Policies
+Prior to the adoption of Transformer models, early control policies for language-conditioned robotic tasks varied significantly in architecture.
 CLIPort [43] integrates the vision and language encoders of CLIP [31] with the Transporter network [94], creating a two-stream architecture.
@@
 BC-Z [113] processes two types of task instructions: a language instruction or a human demonstration video.
-A language instruction undergoes encoding using the USE language encoder [172], while the video is encoded with ResNet18 [29].
-The environment is presented to the model in the form of an RGB image, also encoded using ResNet18.
+The environment is presented to the model in the form of an RGB image.
 Then the instruction embedding and the image embedding are combined through the FiLM layer [148], culminating in the generation of actions.
@@
 This model guides HULC to the actionable region specified by a language instruction, enabling it to fulfill tasks within this designated area.
-Universal Policy (UniPi) [158] treats the decision-making problem as a text-conditioned video generation problem.
+UniPi [128] treats the decision-making problem as a text-conditioned video generation problem.
 To predict actions, UniPi generates a video based on a given text instruction, and actions are extracted from the frames of the video through inverse dynamics.
 This innovative policy-as-video formulation offers several advantages, including enhanced generalization across diverse robot tasks and the potential for knowledge transfer from internet videos to real robots.
-(a) FiLM
-(b) Cross-attention
-(c) Concatenation
-Figure 4: Three common architectures of Transformer-based control policies are characterized by their vision-language fusion methods.
-FiLM layers (Hadamard product $\odot$) are employed to fuse language and vision in RT-1 models.
-Some action decoders utilize cross-attention to condition on the instruction.
-Concatenation ($\oplus$) is the prevailing method of vision-language fusion for action decoders.
 III-B2 Transformer-based Control Policies
+Since the introduction of Transformers, control policies have converged to similar Transformer-based architectures.
 Interactive Language [100] presents a robotic system wherein the low-level control policy can be guided in real-time by human instructions conveyed through language, enabling the completion of long-horizon rearrangement tasks.
@@
 Notably, Hiveformer stands out as one of the early adopters of Transformer architecture as its policy backbone.
-BeT [173], an imitation learning variant of the Trajectory Transformer, addresses the noisy and multi-modal nature of non-expert human demonstration data by using k-means-based action discretization coupled with continuous action correction.
-C-BeT [174] extends BeT by adding a target frame or demonstration as the goal specification.
-VQ-BeT [175] further improves BeT by incorporating vector quantization instead of k-means, enabling better modeling of long-range actions.
 Gato [36] proposes a model that can play Atari games, caption images, and stack blocks, all with a single set of model parameters.
@@
 This self-improvement process iteratively finetunes the model and self-generates new data with the finetuned model.
-Built upon the Gato model [50], RoboCat incorporates the VQ-GAN image encoder [128].
+Built upon the Gato model, RoboCat incorporates the VQ-GAN image encoder [110].
 During training, RoboCat predicts not only the next action but also future observations.
 The effectiveness of the self-improvement process is demonstrated through comprehensive experiments conducted in both simulated and real-world environments under multi-task, multi-embodiment settings.
-RT-1 [3], developed by the same team as BC-Z [130], shares similarities with BC-Z but introduces some key distinctions.
-Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet [132], departing from BC-Z’s use of ResNet18.
-The language instruction is also encoded using USE [172] and is integrated with image embeddings through FiLM layers [131].
-However, RT-1 does not use video as a task instruction, unlike BC-Z.
+RT-1 [114], developed by the same team as BC-Z, shares similarities with BC-Z but introduces some key distinctions.
+Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet [115], departing from BC-Z’s use of ResNet.
+However, RT-1 does not use video as a task instruction.
 Additionally, RT-1 replaces the MLP action decoder in BC-Z with a Transformer decoder, producing discretized actions.
 This modification enables RT-1 to attend to past images, enhancing its performance compared to BC-Z.
-Q-Transformer [134] extends RT-1 [3] by introducing autoregressive Q-functions.
+Q-Transformer [117] extends RT-1 by introducing autoregressive Q-functions.
 In contrast to RT-1, which learns expert trajectories through imitation learning, Q-Transformer adopts Q-learning methods.
@@
 Additionally, RoboAgent introduces a semantic augmentation method that leverages inpainting to augment existing demonstrations.
-III-B3 Multi-modal-instruction Control Policies
-VIMA [176] places a significant emphasis on multi-modal prompts and the generalization capabilities of models.
-By incorporating multi-modal prompts, more specific and intricate tasks can be formulated compared to traditional pure text prompts.
+RoboFlamingo [121] adapts the existing VLM, Flamingo [32, 149], to a robot policy by attaching an LSTM-based policy head to the VLM.
+This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
+III-B3 Control Policies for Multimodal Instructions
+Multimodal instruction enables new ways to specify tasks, such as through demonstrations, by naming novel objects, or by pointing with a finger.
+VIMA [150] places a significant emphasis on multimodal prompts and the generalization capabilities of models.
+By incorporating multimodal prompts, more specific and intricate tasks can be formulated compared to traditional pure text prompts.
 VIMA introduces four main types of tasks: object manipulation, visual goal reaching, novel concept grounding, one-shot video imitation, visual constraint satisfaction, visual reasoning.
@@
 VIMA-Bench has been developed to evaluate across four generalizability levels: placement, combinatorial, novel object, novel task.
-MOO [129] extends RT-1 to handle multi-modal prompts.
-Leveraging the backbone of RT-1, MOO incorporates OWL-ViT [129] to encode images within the prompt.
+MOO [111] extends RT-1 to handle multimodal prompts.
+Leveraging the backbone of RT-1, MOO incorporates OWL-ViT to encode images within the prompt.
 By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1.
@@
 VER [151] also proposes voxelizing multi-view images into 3D cells in a coarse-to-fine manner, enhancing performance in vision-language navigation tasks.
-Perceiver-Actor (PerAct) [120] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations.
+PerAct [103] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations.
 This approach offers a robust structural prior for action learning, enabling natural processing of multi-view observations and facilitating data augmentation in 6-DoF.
@@
 On the other hand, Act3D [104] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization.
-RVT, RVT-2 [122, 123] propose re-rendering images from virtual views and using these images as inputs, rather than directly relying on 3D inputs.
-RoboPoint [124] leverages VLM to predict affordance points directly on 2D images, which are then projected into 3D space using depth maps.
-SceneScript [179] introduces a set of structured language commands that can define an entire scene, specifying the layout and objects.
-This language-based scene representation distinguishes itself from previous methods—such as meshes, voxel grids, point clouds, or radiance fields—by generating scenes in an autoregressive, token-based fashion.
-It achieves competitive results against state-of-the-art methods in layout estimation and object detection.
-III-B5 Diffusion-based Action Generation
-Diffusion-based action generation leverages the success of diffusion models in the field of computer vision [162].
-Diffusion policy [161] formulates a robot policy as Denoising Diffusion Probabilistic Models (DDPMs) [162].
-This approach incorporates a variety of innovative techniques, including receding horizon control, visual conditioning, and the time-series diffusion transformer.
+RVT, RVT-2 [105, 106] propose re-rendering images from virtual views of the scene point cloud and using these images as inputs, rather than directly relying on 3D inputs.
+III-B5 Diffusion-based Control Policies
+Diffusion-based action generation leverages the success of diffusion models in the field of computer vision.
+Diffusion Policy [130] formulates a robot policy as a DDPM [131].
+This approach incorporates a variety of techniques, including receding horizon control, visual conditioning, and the time-series diffusion transformer.
 The effectiveness of this diffusion-based visuomotor policy is underscored by its proficiency in multimodal action distributions, high-dimensional action spaces, and training stability.
-Scaling Up and Distilling Down (SUDD) [164] presents a framework where LLM guides data generation and subsequently the filtered dataset is distilled into a visuo-linguo-motor policy.
-This framework achieves language-guided data generation by composing LLM with a suite of primitive robot utilities, such as grasp samplers, motion planners.
-Then it extends Diffusion Policy [161] through the incorporation of language-based conditioning for multi-task learning.
-This diffusion-based policy learns from successful trajectories, facilitating the distillation of the filtered dataset.
+SUDD [133] presents a framework where an LLM guides data generation, and subsequently, the filtered dataset is distilled into a visuo-linguo-motor policy.
+This framework achieves language-guided data generation by composing an LLM with a suite of primitive robot utilities, such as grasp samplers and motion planners.
+It then extends Diffusion Policy by incorporating language-based conditioning for multi-task learning and facilitates the distillation of the filtered dataset.
 Octo [134] introduces a transformer-based diffusion policy characterized by a modular open-framework design, allowing for flexible connections from different task definition encoders, observation encoders, and action decoders to the Octo Transformer.
@@
 RDT-1B [139] is a diffusion-based foundation model for bimanual manipulation, also built on DiT.
-It addresses data scarcity by introducing a unified action format across various robots, enabling pre-training on heterogeneous multi-robot datasets with over 6K trajectories.
+It addresses data scarcity by introducing a unified action format across various robots, enabling pretraining on heterogeneous multi-robot datasets with over 6K trajectories.
 As a result, RDT scales up to 1.2B parameters and demonstrates zero-shot generalization.
-III-B6 Diffusion-based Control Policy with 3D Vision
+III-B6 Diffusion-based Control Policies with 3D Vision
 Several works have proposed combining 3D Vision with diffusion-based policies.
@@
 Similarly, 3D Diffuser Actor [135] shares the core idea of DP3 but differs in model architecture by combining Act3D with Diffusion Policy.
-III-B7 Motion Planner
+III-B7 Control Policies for Motion Planning
+Motion planning involves decomposing movement tasks into discrete waypoints while satisfying constraints such as obstacle avoidance and kinematic limits.
 Language costs [98] presents a novel approach to robot correction using natural language for human-in-the-loop robot control systems.
@@
 VoxPoser [125] employs LLM and VLM to create two 3D voxel maps that represent affordance and constraint.
-It leverages the programming capability of LLM [157], and the perception capability of VLM models: ViLD [155], MDETR [156], OWL-ViT [133], SAM [31].
+It leverages the programming capability of LLMs, and the perception capability of VLMs.
 LLM translates language instructions into executable code, invoking VLM to obtain object coordinates.
@@
 RoboTAP [152] breaks down demonstrations into stages marked by the opening and closing of the gripper.
-In each stage, RoboTAP uses the TAPIR algorithm to detect active points that track the relevant object from the source pose to the target pose.
-The path can then be used by classical visual servoing to control the robot.
+In each stage, RoboTAP uses the TAPIR algorithm to detect active points that track the relevant object from the source to the target pose.
+The path can then be used by visual servoing to control the robot.
 A motion plan is created by stringing together these stages, enabling few-shot visual imitation.
-III-B8 LLM-based Control Policies
-RT-2 [2] endeavors to harness the capabilities of large multi-modal models in robotics tasks, drawing inspiration from models like PaLI-X [140] and PaLM-E [181].
+III-B8 Control Policies with Point-based Action
+Recent research has explored leveraging the capabilities of VLMs to select or predict point-based actions—a cost-effective alternative to building full VLAs.
+PIVOT [153] casts robotics tasks as visual question answering, leveraging VLMs to select the best robot action from a set of visual proposals.
+Visual proposals are annotated in the form of keypoints on images.
+The VLM is iteratively prompted to refine them until the best option is identified.
+RoboPoint [107] finetunes VLM using the task of spatial affordance prediction, which is to point at where to act on the image.
+These affordance points on 2D images are subsequently projected into 3D space using depth maps, forming the predicted robot action.
+ReKep [51] is a constraint function that maps 3D keypoints in the scene to a numerical cost.
+Robotics manipulation tasks can be represented as a sequence of ReKep constraints, which are produced with large vision models and VLMs.
+Consequently, robot actions can be obtained by solving the constrained optimization problem.
+III-B9 Large VLA
+Large VLA is equivalent to the original VLA definition proposed by RT-2 [2], as illustrated in Figure 2a.
+This terminology is analogous to the distinction between LLMs and general language models, or between large VLMs and general VLMs.
+RT-2 [2] endeavors to harness the capabilities of large multimodal models in robotics tasks, drawing inspiration from models like PaLI-X and PaLM-E.
 The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data.
 This training scheme enhances the model’s generalizability and brings about emergent capabilities.
-RT-2 represents an effort to integrate low-level control policy [3] with high-level task planners in pursuit of a more comprehensive robotic system.
-RT-X [143] builds upon the previous RT-1 and RT-2 models.
-These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets.
-OXE comprises 160,266 tasks from 527 skills with 22 embodiments.
-The resulting models, RT-1-X and RT-2-X, both outperform their original versions, owing to the more diverse and larger OXE dataset.
-OpenVLA [144] was later developed as an open-source counterpart to RT-X.
 RT-H [142] introduces an action hierarchy that includes an intermediate prediction layer of language motions, situated between language instructions and low-level actions (translation and rotation).
@@
 Moreover, this action hierarchy enables users to specify corrections to recover from failures, which the model can then learn from.
-RoboFlamingo [46] adapts the existing VLM, Flamingo [11, 182], to a robot policy by attaching an LSTM-based policy head to the VLM.
-This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
+RT-X [143] builds upon the previous RT-1 and RT-2 models.
+These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets.
+The resulting models, RT-1-X and RT-2-X, both outperform their original versions.
+OpenVLA [50] was later developed as an open-source counterpart to RT-2-X. They additionally explored efficient fine-tuning methods, including LoRA and model quantization.
+$\pi_{0}$ [147] proposes a flow-matching architecture for transforming VLMs into VLAs.
+By incorporating an additional action expert based on the mixture-of-experts framework, it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
+(a) FiLM
+(b) Cross-attention
+(c) Concatenation
+Figure 5: Three common vision-language fusion methods of Transformer-based control policies.
+FiLM layers (Hadamard product $\odot$) are used in RT-1 models.
+Some utilize cross-attention to condition on instructions.
+Concatenation ($\oplus$) is the prevailing method in LVLAs.
 Pros and Cons.
@@
 Action types and their training objectives
-Most low-level control policies predict actions for the end-effector pose while abstracting away the motion planning module that controls the motion of individual joints with inverse kinematics.
+Most low-level control policies predict actions for the end-effector pose while abstracting away the motion planning module that produces more fine-grained motions.
 While this abstraction facilitates better generalization to different embodiments, it also imposes limitations on dexterity.
 The behavior cloning (BC) objective is used in imitation-learning, with different variants for different action types. The BC objective for continuous action [121] can be written as:
-$\mathcal{L}_{\text{Cont}}=\sum_{t}\operatorname{MSE}\left(\mathbf{a}_{t},\hat{\mathbf{a}}_{t}\right),$ (1)
-where $\operatorname{MSE}(\cdot)$ stands for mean squared error.
-$\hat{\mathbf{a}}_{t}=\operatorname{argmax}_{a_{t}}\pi_{\theta}(a_{t}|p,s_{\leq t},a_{<t})$ is predicted action while $\mathbf{a}_{t}$ is the action from expert demonstrations.
+$\mathcal{L}_{\text{Cont}}=\sum_{t}\operatorname{MSE}\left(a_{t},\hat{a}_{t}\right),$ (4)
+where $\operatorname{MSE}(\cdot)$ stands for mean squared error. $a_{t}$ is the action annotation from expert demonstrations.
 Discrete action is achieved by dividing the action value range into a number of bins [114]. Its BC objective is:
-$\mathcal{L}_{\text{Disc}}=\sum_{t}\operatorname{CE}\left(\mathbf{a}_{t},\hat{\mathbf{a}}_{t}\right),$ (2)
+$\mathcal{L}_{\text{Disc}}=\sum_{t}\operatorname{CE}\left(a_{t},\hat{a}_{t}\right),$ (5)
 where $\operatorname{CE}(\cdot)$ stands for cross-entropy loss.
 CLIPort [43] and VIMA [150] use $SE(2)$ action and its BC objective can be expressed as follows:
-$\begin{split}\mathcal{L}_{\text{{SE}(2)}}&=\operatorname{CE}(\mathbf{a}_{\text{pick}},\hat{\mathbf{a}}_{\text{pick}})+\operatorname{CE}(\mathbf{a}_{\text{place}},\hat{\mathbf{a}}_{\text{place}})\end{split}$ (3)
+$\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}})\end{split}$ (6)
 The DDPM objective in diffusion-based control policies [130] is represented as:
-$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}\left(\mathbf{x}^{0}+\varepsilon^{k},k\right)\right)$ (4)
-where $\mathbf{x}^{0}$ is the raw example from the dataset and $\varepsilon^{k}$ is a random noise for iteration $k$; $\varepsilon_{\theta}$ is the noise prediction network, i.e., the VLA model.
+$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right)$ (7)
+where $\varepsilon^{k}$ is a random noise for iteration $k$; $\varepsilon_{\theta}$ is the noise prediction network, i.e., the VLA model.
 While discrete action has demonstrated superior performance in RT-1 [114], Octo [134] argues that it leads to early grasping issues.
-$\textbf{SE}(2)$ action only necessitates the model to predict the two end-effector poses, e.g., pick pose and place pose.
-While this action type can be predicted in at most two forward passes, it also imposes limitations on action dexterity and generalizability.
-LLM vs non-LLM
-While LLM-based control policies can greatly enhance instruction-following abilities because LLM can better parse user intentions, concerns arise regarding their training cost and deployment speed.
-Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during LLM inference.
+$SE(2)$ actions require the model to predict only the pick and place end-effector poses, which is sufficient for many tabletop manipulation tasks.
+However, more complex tasks—such as “pouring water into a cup”—may require additional DoFs, thereby necessitating $SE(3)$ actions [51].
+Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner [51].
+Table IV: High-level task planners. VL: vision-language fusion method.
+Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
+SayCan [10] - PaLM, FLAN - BC-Z, MT-Opt No Real-Navi&Mani (Everyday): office kitchen
+[…]
+(SL)3 [155] - T5-small - seq2seq Yes Sim-Navi: ALFRED
+EmbodiedGPT [156] EVA ViT-G/14 [157] LLaMA-7B Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
+PaLM-E [11] ViT, OSRT [158] PaLM Concat Interactive Language, RT-1 Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
+LEO [159] ConvNeXt [160], PointNet++ [161] Vicuna-7B Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
+3D-LLM [57] CLIP, 3D-CLR [162], ConceptFusion [163] Flamingo, BLIP-2 Concat DD-PPO [164] Yes Sim-Navi: Habitat
+ShapeLLM [165] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
+Inner Monologue [9] MDETR, CLIP InstructGPT, PaLM Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
+LLM-Planner [166] HLSM GPT-3 Language HLSM Yes Sim-Navi: ALFRED
+LID [167] CNN GPT-2 Language & Embed - Yes Sim: VirtualHome
+SMs [168] CLIP, ViLD GPT-3, RoBERTa [169] Language & Code CLIPort No Sim-Mani: pick, place
+ConceptGraphs [170] SAM, CLIP, LLaVA [8] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
+ProgPrompt [171] ViLD GPT-3 Code API No Sim: VirtualHome
+ChatGPT for Robotics [172] YOLOv8 [173] ChatGPT Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
+CaP [174] ViLD, MDETR GPT-3, Codex Code API No Sim-Mani & Sim-Navi: pick, place, etc.
+DEPS [175] CLIP ChatGPT Code MC-Controller, Steve-1 No Sim: Minecraft
 RT series
 RT-1 [114] inspired a series of “Robotic Transformer” models.
-Preceding RT-1 was BC-Z [130], which solely utilized MLP layers for action prediction.
+The Transformer backbone surpasses previous RNN backbones by harnessing the higher capacity of Transformers to absorb larger robot datasets.
+Preceding RT-1 was BC-Z, which solely utilized MLP layers for action prediction.
 Subsequent to RT-1, several works emerged, each introducing new capabilities.
-MOO [129] adapted RT-1 to accommodate multi-modal prompts.
-RT-Trajectory [135] enabled RT-1 to process trajectory sketches as prompts.
-Q-Transformer [134] utilized Q-learning to train RT-1.
-RT-2 [2], based on ViT and LLM, introduced a completely different architecture from RT-1.
-RT-X [143] retrained RT-1 and RT-2 with a significantly larger dataset, resulting in improved performance.
-The Transformer backbone surpasses previous RNN backbones [110, 111, 112] by harnessing the higher capacity of Transformers to absorb larger robot datasets.
+MOO adapted RT-1 to accommodate multimodal prompts.
+RT-Trajectory enabled RT-1 to process trajectory sketches as prompts.
+Q-Transformer utilized Q-learning to train RT-1.
+RT-2, based on ViT and LLM, introduced a completely different architecture from RT-1.
+RT-X retrained RT-1 and RT-2 with a significantly larger dataset, resulting in improved performance.
 Based on RT-2, RT-H [142] introduces action hierarchies for better data sharing.
-(a) Language-based
-(b) Code-based
-Figure 5: Different approaches to connect LLM to multi-modal modules in high-level task planner.
-III-C High-level Task Planner
-A high-level task planner aims to divide a complex task into a series of executable sub-tasks.
-This approach is often referred to as task decomposition, and is also related to the classical Task and Motion Planning (TAMP) setting.
-Many high-level task planners are constructed on top of LLMs.
-While integrating multi-modal modules into LLMs in an end-to-end fashion (§III-C1) is a common approach, training with multi-modal data can be expensive.
-Consequently, some task planners explore to utilize language (§III-C3) or code (§III-C4) as the medium for exchanging multi-modal information instead, as they can be natively processed by LLMs.
-III-C1 End-to-end
-While control policies are effective in understanding and executing simple language instructions, they tend to struggle in long-horizon tasks that involve multiple sub-tasks.
-Large language models have been identified as powerful tools to interpret such complex long-horizon tasks.
-Consequently, several approaches are dedicated to integrating LLMs as high-level task planners.
-Their goal is to decompose long-horizon tasks into simpler sub-tasks, enabling low-level control policies to execute them sequentially, contributing to the development of a hierarchical robotic system.
-Table IV outlines important details about these high-level task planners.
+LVLA vs VLA
+While LVLAs can greatly enhance instruction-following abilities because they can better parse user intentions, concerns arise regarding their training cost and deployment speed.
+Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during inference.
+Therefore, several methods have been proposed to improve efficiency.
+TinyVLA [176] focused on inference speed and data efficiency through smaller VLM and diffusion head for robot action.
+DeeR-VLA [177] proposes only partially activating the model with dynamic inference with early-exit.
+IV Task Planners
+A high-level task planner $\pi_{\phi}$ aims to divide a complex task $\ell$ into a sequence of subtasks $p_{i}$ (i.e., task plan), each serving as a language instruction to low-level control policies $\pi_{\theta}$:
+$\begin{split}\left[p_{1},p_{2},\dots,p_{N}\right]\sim\pi_{\phi}(\ell,s_{t}).\end{split}$ (8)
+This process is often referred to as task decomposition and is also related to the task and motion planning (TAMP) setting.
+Ideally, task plans should also incorporate optimal scheduling for these subtasks.
+When equipped with task planners, VLAs can complete more complex, long-horizon tasks, as illustrated in Figure 4.
+Table IV outlines important details about task planners.
+IV-A Monolithic Task Planners
+A single LLM or multimodal LLM (MLLM) can typically generate task plans by employing a tailored framework or through finetuning on embodied datasets. We refer to these as monolithic models.
+IV-A1 Grounded Task Planners
+Grounded task planning involves generating high-level actions while considering whether they can be executed by low-level control policies.
 SayCan [10] is a framework designed to integrate high-level LLM planners with low-level control policies.
@@
 By considering both LLM’s plan and affordance, the framework selects the optimal skill for the current state.
-LID [184] introduces a novel data collection procedure termed Active Data Gathering (ADG).
-A key aspect of ADG is hindsight relabeling, which reassigns labels to unsuccessful trajectories, effectively maximizing the utilization of data irrespective of their success.
-By converting all environmental inputs into textual descriptions, their language-model-based policy demonstrates enhanced combinatorial generalization.
-Table IV: High-level task planners. VL: vision-language fusion. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation.
-Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
-SayCan [183] CNN PaLM [185]
-FLAN [186]
-Concat BC-Z [130]
-MT-Opt [12]
-Yes Real-Navi&Mani (Everyday Robots): kitchen
-LID [184] CNN GPT-2 [187] Concat - Yes Sim: VirtualHome
 […]
-Codex [190]
-- - No Sim: VirtualHome
-(SL)3 [191] ResNet18 [29] T5-small [10] Xattn HLSM [192] Yes Sim-Navi: ALFRED
-EmbodiedGPT [101] EVA ViT-G/14 [193] LLaMA-7B [150] Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
-PaLM-E [181] ViT [30],
-OSRT [194]
-PaLM [185] Concat Interactive Language policy [117], RT-1 [3] Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
-LEO [195] ConvNeXt [196], PointNet++ [197] Vicuna-7B [125] Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
-3D-LLM [198] CLIP, 3D-CLR [199], ConceptFusion [200] Flamingo [11], BLIP-2 [16] Concat DD-PPO [201] Yes Sim-Navi: Habitat
-StructFormer [202], StructDiffusion [203] PCT [204] - Concat [205], RRT-connect [206] Yes Sim-Mani: PyBullet; Real-Mani (Franka): rearrange
-ShapeLLM [207] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
-Inner Monologue [208] MDETR [156], CLIP InstructGPT [209]
-PaLM [185]
-Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
-LLM-Planner [210] HLSM object detector [192] GPT-3 [189] Language HLSM [192] Yes Sim-Navi: ALFRED
-SMs [211] CLIP, ViLD [155] GPT-3 [189], RoBERTa [212] Language & Code CLIPort [13] No Sim-Mani: pick, place
-ProgPrompt [213] ViLD [155] GPT-3 [189] Code API [214, 215] No Sim: VirtualHome
-ChatGPT for Robotics [216] YOLOv8 [217] ChatGPT [1] Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
-CaP [218] ViLD [155], MDETR [156] GPT-3 [189]
-Codex [190]
-Code API No Sim-Mani & Sim-Navi: pick, place, etc.
-DEPS [219] CLIP ChatGPT [220] Code MC-Controller, Steve-1 No Sim: Minecraft
-ConceptGraphs [221] SAM [31], CLIP, LLaVA [45] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
-[…]
-Semi-Supervised Skill Learning with Latent Language ((SL)3) [191] is a learning algorithm that alternates between three steps: segmentation, labeling, and parameter update.
+(SL)3 [155] is a learning algorithm that alternates between three steps: segmentation, labeling, and parameter update.
 In the segmentation step, high-level subtasks are aligned with low-level actions, subtask descriptions are then inferred in the labeling step, and finally, the network parameters are updated.
 This approach enables a hierarchical policy to discover reusable skills with sparse natural language annotations.
-EmbodiedGPT [101] introduces the embodied-former, which outputs task-relevant instance-level features.
-This is achieved by incorporating information from vision encoder embeddings and embodied planning information provided by LLM.
-The instance feature serves to inform the low-level policy about the immediate next action to take.
-PaLM-E [181] integrates ViT [30] and PaLM [185] to create a large embodied multi-modal language model capable of performing high-level embodied reasoning tasks.
+IV-A2 End-to-end Task Planners
+Similar to LVLAs, task planners can be implemented as end-to-end multimodal LLMs, leveraging their Internet-scale knowledge for task planning. Clearly, they are also monolithic models.
+PaLM-E [11] integrates ViT and PaLM to create a large embodied multimodal language model capable of performing high-level embodied reasoning tasks.
 Based on perceived images and high-level language instructions, PaLM-E generates a text plan that serves as instructions for low-level robotic policies.
@@
 With PaLM as its backbone, PaLM-E can handle both normal visual question answering (VQA) tasks, along with the additional embodied VQA tasks.
-III-C2 3D Vision
-3D information is vital in embodied tasks, and thus it is also explored for task planners.
-Because the majority of current VLMs deal with images as visual inputs, it requires architectural change to VLMs to incorporate 3D visual inputs and therefore they are usually end-to-end models.
-LEO [195] identifies the conventional use of image inputs to be a limiting factor for multi-modal generalist agents in interacting with the 3D world.
-This novel approach involves training an LLM-based architecture using a new dataset with two stages.
-The first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning.
-LEO demonstrates proficiency not only in 3D captioning and question-answering tasks but also in embodied tasks, including embodied reasoning, embodied navigation, and robotic manipulation.
+EmbodiedGPT [156] introduces the embodied-former, which outputs task-relevant instance-level features.
+This is achieved by incorporating information from vision encoder embeddings and embodied planning information provided by LLM.
+The instance feature serves to inform the low-level policy about the immediate next action to take.
+IV-A3 End-to-end Task Planners with 3D Vision
+Some task planners also explore the use of 3D vision.
+Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore they are usually end-to-end models.
+LEO [159] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning.
+LEO demonstrates proficiency not only in 3D question-answering tasks but also in manipulation, navigation, and task planning.
 3D-LLM [57] injects 3D information into LLMs and empowers them to perform 3D tasks, such as 3D-assisted dialog, and navigation.
@@
 MultiPLY [179] is an object-centric embodied LLM that incorporates even more modalities, including audio, tactile, and thermal.
-StructFormer [202] addresses the task of arranging objects into complex structures.
-Previous methods rely on pairwise semantic relations and sequential manipulation, while StructFormer explicitly reasons about the relationships among all objects and predicts their target positions.
-StructDiffusion [203] extends this by incorporating a diffusion-based generation approach.
 ShapeLLM [165] is built on the novel 3D vision encoder ReCon++, which distills knowledge from multi-view image and text teachers, along with a point cloud MAE.
 By integrating ReCon++ with LLaMA, ShapeLLM improves embodied interaction performance on their newly proposed 3D benchmark, 3D MM-Vet.
-Table V: Recent datasets for robot learning.
-Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects.
-$\ast$ Dataset collected in simulators rather than in the real world.
-Some datasets are continuously updated; this table is based on the information from their original papers.
-This table is partially based on [224, 149].
-Name Collection Visual Instruction Embodiment Skills Tasks Scenes Objects Episodes
-MIME [225] Human Teleop RGB, D Demo Baxter 12 20 1 - 8.3K
-RoboTurk [226] $\ast$ Human Teleop RGB - Sawyer 2 - 1 - 2.1K
-RoboNet [227] Scripted RGB Goal image (7 robots) - - 10 - 162K
-MT-Opt [12] Scripted RGB Lang (7 robots) 2 12 1 - 800K
-BC-Z [130] Human Teleop RGB Lang, Demo Everyday 3 100 1 - 25.9K
-RT-1 [3] Human Teleop RGB Lang Everyday 12 700+ 2 16 130K
-MOO [129] Human Teleop RGB Multi-modal Everyday 5 - - 106 59.1K
-VIMA [176] $\ast$ Scripted RGB Multi-modal UR5 17 - 1 29 650K
-RoboSet [228] Human, Scripted RGB, D Lang Franka Panda 12 38 11 - 98.5K
-BridgeData V2 [229] Human, Scripted RGB, D Lang WidowX 250 13 - 24 100+ 60.1K
-RH20T [224] Human Teleop RGB, D Lang (4 robots) 42 147 7 - 110K+
-DROID [149] Human Teleop RGB, D Lang Franka Panda 86 - 564 - 76K
-OXE [143] Aggregated Datasets RGB, D Lang (22 robots) 527 160,266 311 - 1M+
-III-C3 Language-based
+(a) Language-based
+(b) Code-based
+Figure 6: Different approaches to connect LLM to multi-modal modules in modular task planners.
+IV-B Modular Task Planners
+Finetuning end-to-end models on embodied data can be expensive, and there are approaches that adopt a modular design by assembling off-the-shelf LLMs and VLMs into task planners.
+IV-B1 Language-based Task Planners
+Language-based approaches use natural language descriptions as the medium for exchanging multimodal information, as illustrated in Figure 6a.
 Inner Monologue [9] sits between high-level command and low-level policy to enable closed-loop control planning.
@@
 As the feedback is communicated to the LLM in textual format, no additional training is required for the LLM.
-A similar approach is also proposed in ReAct [104], which supports more flexible reasoning traces, allowing for more diverse types of reasoning for various tasks.
+A similar approach is used in ReAct [92].
 LLM-Planner [166] introduces a novel approach to constructing a hierarchical policy comprising a high-level planner and a low-level planner.
@@
 While sharing similarities with previous methods in its overall architecture, LLM-Planner distinguishes itself by incorporating a re-planning mechanism, aiding the robot to “get unstuck”.
+LID [167] introduces a novel data collection procedure termed Active Data Gathering (ADG).
+A key aspect of ADG is hindsight relabeling, which reassigns labels to unsuccessful trajectories, effectively maximizing the utilization of data irrespective of their success.
+By converting all environmental inputs into textual descriptions, their language-model-based policy demonstrates enhanced combinatorial generalization.
 Socratic Models (SMs) [168] presents a unique framework wherein diverse pretrained models are effectively composed without the need for finetuning.
@@
 In addition to natural language plans, task plans can also be represented in the form of pseudocode.
-III-C4 Code-based
+Table V: Recent datasets for robot learning.
+Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects.
+$\ast$ Dataset collected in simulators rather than in the real world.
+Some datasets are continuously updated; this table is based on the information from their original papers.
+Table adapted from [180, 146].
+Name Collection Visual Instruction Embodiment Skills Tasks Scenes Objects Episodes
+MIME [181] Human Teleoperation RGB, D Demo Baxter 12 20 1 - 8.3K
+RoboTurk [182] $\ast$ Human Teleoperation RGB - Sawyer 2 - 1 - 2.1K
+RoboNet [183] Scripted RGB Goal image (7 robots) - - 10 - 162K
+MT-Opt [184] Scripted RGB Lang (7 robots) 2 12 1 - 800K
+BC-Z [113] Human Teleoperation RGB Lang, Demo Everyday 3 100 1 - 25.9K
+RT-1-Kitchen [114] Human Teleoperation RGB Lang Everyday 12 700+ 2 16 130K
+MOO [111] Human Teleoperation RGB Multi-modal Everyday 5 - - 106 59.1K
+VIMA [150] $\ast$ Scripted RGB Multi-modal UR5 17 - 1 29 650K
+RoboSet [185] Human, Scripted RGB, D Lang Franka Panda 12 38 11 - 98.5K
+BridgeData V2 [186] Human, Scripted RGB, D Lang WidowX 250 13 - 24 100+ 60.1K
+RH20T [180] Human Teleoperation RGB, D Lang (4 robots) 42 147 7 - 110K+
+DROID [146] Human Teleoperation RGB, D Lang Franka Panda 86 - 564 - 76K
+OXE [143] Aggregated Datasets RGB, D Lang (22 robots) 527 160,266 311 - 1M+
+IV-B2 Code-based Task Planners
+Code-based task planners leverage the coding ability of LLMs to generate task plans in the form of a program.
+Object detectors, VLMs, and control policies can be invoked via APIs, as shown in Figure 6b.
 ProgPrompt [171] introduces a novel task-planning approach by prompting LLMs with program-like specifications detailing available actions and objects.
@@
 Environmental feedback can be incorporated through assertions within the program.
-This prompting scheme leverages the world knowledge and programming skills of LLMs.
 ChatGPT for Robotics [172] take advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology.
 The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code.
-In this procedure, ChatGPT serves as a high-level task planner, similar to PaLM-E [181], and actions are generated through function calls to corresponding low-level APIs.
+In this procedure, ChatGPT serves as a high-level task planner and actions are generated through function calls to corresponding low-level APIs.
 Code as policies (CaP) [174] also leverages the code-writing capability of LLMs.
-It employs GPT-3 [189] or Codex [190] to generate policy code, which, in turn, invokes perception modules and control APIs.
+It employs GPT-3 or Codex to generate policy code, which, in turn, invokes perception modules and control APIs.
 CaP exhibits proficiency on spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives.
@@
 Pros and Cons.
-While end-to-end task planners like SayCan [183] share a similar architecture to low-level control policies and can be optimized for specific tasks, their training costs can be prohibitive due to the large model size of the combined LLM and visual transformer.
-Language-based task planners offer the advantage of seamless integration with existing language-conditioned control policies.
-However, they often require finetuning or alignment methods to map the generated plan to executable language instructions for low-level control policies.
-On the other hand, code-based task planners leverage the programming capability of LLMs to connect perception and action modules.
-This approach does not require additional training, but its performance may be constrained by the capabilities of existing models.
-IV Datasets and Benchmarks
-Table VI: Simulators for benchmarking VLA models.
-D: Depth.
-Seg: segmentation.
-A: Audio.
-N: Normals.
-Force: simulated contact force between the end-effector and the item.
+Monolithic task planners that utilize grounded task planning focus on generating executable plans.
+End-to-end models like PaLM-E share an architecture similar to most LVLAs and can be finetuned on specialized embodied data to achieve better performance.
+However, the training costs of such large models can be a significant concern.
+In contrast, modular task planners are more readily deployable because they leverage off-the-shelf LLMs and VLMs.
+Language-based task planners offer the advantage of seamless integration of LLMs and VLMs, as they are designed to operate in the natural language space.
+However, they often require extra steps to align the generated task plans with language instructions that are admissible to low-level control policies.
+On the other hand, while code-based task planners may require manually wrapping VLMs and control policies in APIs and preparing clear documentation in advance, they enable code debugging and provide greater controllability.
+Nevertheless, their performance can be constrained by the programming capabilities of existing models.
+V Datasets and Benchmarks
+Table VI: Simulators and simulated benchmarks.
+Control: Continuous control tasks.
+D, S, A, N: Depth, segmentation, audio, normals.
+Force: simulated contact force between end-effector and item.
 PD: Pre-defined.
-Vers: versions.
-This table is partially based on [231, 232].
+Table adapted from [188, 189].
 Name Scenes
@@
 /Cat
-UI Engine Task Observation Action Agent Description Related
-Gibson [233] 572/- - - Pybullet Navi RGB, D, N, Seg - - Navi only -
-iGibson [234, 235, 232] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, Seg, N, Flow, LiDAR Force Husky, TurtleBot v2, LoCoBot, etc.
+UI Physics Engine Task Observation Action Agent Description Related
+Gibson [190] 572/- - - Pybullet Navi RGB, D, N, S - - Navi only -
+iGibson [191, 192, 189] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, S, N, Flow, LiDAR Force TurtleBot v2, LoCoBot, etc.
 VR, Continuous Extended States.
-Vers: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [231], BEHAVIOR-1K [236]
-SAPIEN [237] - 2346/- Code PhysX Navi, Mani RGB, D, Seg Force Franka Articulation, Ray Tracing VoxPoser
-AI2-THOR [238] -/120 118/118 Mouse Unity Navi, Mani RGB, D, Seg, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Vers: [239, 240] Benchmarks: ALFRED, RoPOR [241]
-VirtualHome [242] 7/- -/509 Lang Unity Navi, Mani RGB, D, Seg, F Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
-TDW [243] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, Seg, A Force Fetch, Sawyer, Baxter Audio, Fluids -
-RLBench [119] 1/- 28/28 Code Bullet Mani RGB, D, Seg Force Franka Tiered Task Difficulty Hiveformer, PerAct
+Versions: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [188], BEHAVIOR-1K [193]
+SAPIEN [194] - 2346/- Code PhysX Navi, Mani RGB, D, S Force Franka Articulation, Ray Tracing VoxPoser. Benchmarks: SIMPLER [195]
+AI2-THOR [196] -/120 118/118 Mouse Unity Navi, Mani RGB, D, S, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Versions: [197, 198] Benchmarks: ALFRED, RoPOR [199]
+VirtualHome [200] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
+TDW [201] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids -
+RLBench [102] 1/- 28/28 Code Bullet Mani RGB, D, S Force Franka Tiered Task Difficulty Hiveformer, PerAct
 Meta-World [202] 1/- 80/7 Code MuJoCo Mani Pose Force Sawyer Meta-RL R3M, VC-1, Vi-PRoM, EmbodiedGPT
@@
 Franka Kitchen [204] 1/- 10/6 VR MuJoCo Mani Pose Force Franka Extended by R3M with RGB R3M, Voltron, Vi-PRoM, Diffusion Policy, EmbodiedGPT
-Habitat [247, 248] Matterport + Gibson Mouse Bullet Navi RGB, D, Seg, A Force Fetch, Franka, AlienGO Fast, Navi only. Vers: Rearrangement [249], Habitat 2.0 VC-1, PACT; OVMM [250]
-ALFRED [251] -/120 84/84 - Unity Navi, Mani RGB, D, Seg PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
+Habitat [205, 206] Matterport + Gibson Mouse Bullet Navi RGB, D, S, A Force Fetch, Franka, AlienGO Fast, Navi only. Versions: Rearrangement [207], Habitat 2.0 VC-1, PACT; OVMM [208]
+ALFRED [209] -/120 84/84 - Unity Navi, Mani RGB, D, S PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
 DMC [210] 1/- 4/4 Code MuJoCo Control RGB, D Force - Continuous RL VC-1, SMART
 OpenAI Gym [211] 1/- 4/4 Code MuJoCo Control RGB Force - Single agent RL environments -
-The collection of real-world robot data presents significant challenges due to various factors.
-Firstly, the collection process is hindered by the substantial costs associated with procuring robot equipment, setting up environments, and dedicating extensive human resources.
-Secondly, the gathering of expert demonstration data requires considerable time investments.
-Thirdly, the diverse types and configurations of robots introduce inconsistency in sensory data, control modes, gripper types, etc.
-Lastly, achieving precision in capturing object 6D poses and exactly replicating or resetting setups remains elusive.
-As a result of these difficulties, public real-world robot datasets are relatively scarce.
-Furthermore, evaluating a robot system’s performance in real-world conditions introduces another layer of complexity, as it is challenging to precisely reproduce a setup and often requires human supervision.
-We summarize the robot datasets from recent VLAs in Table V.
-Consequently, many researchers resort to simulated environments as a means to mitigate these obstacles and accelerate the data collection process.
-We compare the most relevant simulators in Table VI.
-Nevertheless, this strategy presents its own challenges, chief among them being the sim-to-real gap.
-This discrepancy arises when models trained on simulated data exhibit poor performance in real-world deployment.
-The causes of this gap are multifaceted, encompassing differences in rendering quality, inaccuracies in physics simulations, and domain shifts characterized by unrealistic object properties and robot motion planners.
-For instance, simulating non-rigid objects like liquids or towels presents significant difficulties.
-Moreover, incorporating new objects into simulators requires considerable effort, often involving techniques such as 3D scanning and mesh editing.
+Genesis [212] (Rigid, deformable, liquid, etc.) Code (Propri-etary) Navi, Mani RGB, D, S, N Force Franka, Unitree, etc. High-speed comprehensive physics simulation -
+V-A Real-world Robot Datasets & Benchmarks
+Embodied AI faces significant data scarcity issues because real-world robot data is not as readily available as language data.
+Collecting real-world robot datasets poses multiple challenges.
+Firstly, it is impeded by the cost and time required to procure robotic equipment, set up environments, and gather expert data through dedicated policies or human teleoperation.
+Secondly, the diverse types and configurations of robots introduce inconsistencies in sensory data, control modes, gripper types, etc.
+Lastly, accurately capturing object 6D poses and reproducing setups remains elusive.
+We summarize recent robot datasets in Table V.
+In addition, real-world benchmarks are further complicated by the need for human evaluation.
+V-B Simulators, Simulated Robot Datasets & Benchmarks
+Many researchers resort to simulated environments to circumvent real-world obstacles and scale the data collection process.
+We compare simulators and simulated benchmarks in Table VI.
+Nevertheless, this strategy presents its own challenges, chief among which is the sim-to-real gap.
+This discrepancy arises when models trained on simulated data exhibit poor performance during real-world deployment.
+The causes of this gap are multifaceted, encompassing unrealistic rendering quality, inaccuracies in physics simulations, and domain shifts characterized by object properties and robot motion planners.
+For instance, simulating non-rigid objects such as deformable objects or liquids presents significant difficulties.
+Moreover, importing new objects into simulators requires considerable effort, often involving techniques such as 3D scanning and mesh editing.
 Despite these hurdles, simulated environments provide automated evaluation metrics that aid researchers in consistently evaluating robotic models.
-Many benchmarks are based on simulators because only simulated environments can reproduce the experimental setup exactly, while real-world evaluation varies between different models, making it infeasible to compare them reliably.
-The HomeRobot OVMM Benchmark [250] proposes a sim-to-real benchmark, yet its consistency remains to be observed.
-Several approaches advocate for automated data collection.
+Many benchmarks are based on simulators because they can precisely reproduce the experimental setup and yield fair comparisons of different models.
+Another technique, known as real-to-sim, can improve simulation fidelity, recreate failure cases, or facilitate digital twins.
+V-C Automated Dataset Collection
+Several approaches advocate for automated dataset collection.
 RoboGen [213] employs a generative simulation paradigm that proposes interesting skills, simulates corresponding environments, and selects optimal learning approaches to train policies for acquiring those skills.
@@
 RoboPoint [107] generates scenes procedurally with randomized 3D layouts, objects, and camera viewpoints.
-An alternative strategy to address the data scarcity issue in real-world settings is to leverage human data.
+V-D Human Datasets
+An alternative strategy to address data scarcity in real-world settings is to leverage human data.
 Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity.
@@
 These obstacles underscore the complexities involved in incorporating human data into robot learning processes.
-Due to limited space, we refer interested readers to a comprehensive comparison of existing datasets in [257].
+UMI [216] proposes a method to mitigate these issues using hand-held grippers.
+For a more comprehensive comparison of human datasets, we refer interested readers to [217].
+V-E Task Planning Benchmarks
+EgoPlan-Bench [218] focuses on benchmarking real-world task planning with human annotations.
+PlanBench [219, 220] comprehensively assesses various aspects of task planning ability, such as cost optimality, plan verification, and replanning.
+LoTa-Bench [221] directly evaluates task planning by executing the generated plans in simulators and calculating success rates.
+Embodied Agent Interface (EAI) [222] argues that this approach fails to pinpoint issues in LLMs.
+By formalizing the input-output of LLM-based modules for decision-making tasks, EAI enables more fine-grained metrics beyond success rates.
+V-F Embodied Question Answering Benchmarks
+Embodied question answering (EQA) benchmarks, as summarized in Table VII, do not directly evaluate robotic tasks like manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge.
+EQA is akin to previous visual question answering benchmarks, but differs in that the agent can actively explore the environment before providing an answer.
+EmbodiedQA [223] and IQUAD [224] were among the first works to introduce this type of benchmarks.
+MT-EQA [225] focuses on complex questions involving multiple targets.
+MP3D-EQA [226] converts previous RGB inputs to point clouds, testing 3D perception capabilities.
+Active exploration requires access to a simulator, limiting the types of data that can be used, such as real-world videos.
+EgoVQA [227] shifts the focus of VQA to egocentric videos.
+EgoTaskQA [228] emphasizes spatial, temporal, and causal relationship reasoning.
+EQA-MX [229] investigates multimodal expressions (MX), including regular verbal utterances and nonverbal gestures like eye gaze and pointing.
+OpenEQA [230] evaluates seven main categories, including functional reasoning and world knowledge.
 Table VII: Embodied question answering benchmarks. Acc: accuracy. PPL: perplexity.
@@
 EgoTaskQA [228] 40K 2K LEMMA Answer set (open-answer and binary verifications) No Human annotators for relationship triples, then template Acc
-EgoPlan [264] 3,355 2,432 + 923 Epic-Kitchen + Ego4D Multiple-choice (1 out of 4) No GPT-4 filtering, question template, human verification Acc
+EQA-MX [229] 8.2M 750K images CAESAR simulator Answer set No Question templates, answer set (simulator ground truth) Acc
 OpenEQA [230] 557 + 1079 180 envs HM3D + ScanNet Open-answer (LLM scorer) No Human annotators Score
-EgoCOT [101] 5,246 plans 129 Ego4D Caption and planning (not QA) No ChatGPT PPL
-EQA-MX [266] 8.2M 750K images CAESAR simulator Answer set No Question templates, answer set (simulator ground truth) Acc
-LoTa-Bench [267] 17K instruction-plan pairs AI2-THOR, VirtualHome Goal condition - ALFRED + WAH-NL Success
-Furthermore, some datasets and benchmarks may not be directly dedicated to robot manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge.
-One of the most prominent embodied tasks is embodied question answering (EQA), as summarized in Table VII.
-EQA is akin to previous visual question answering and video question answering tasks but differs in that the agent can actively explore the environment before providing an answer.
-EmbodiedQA [258] and IQUAD [259] were among the first works to introduce this task.
-MT-EQA [260] focuses on questions involving multiple targets, increasing the complexity of understanding and answering the questions.
-MP3D-EQA [261] converts previous RGB input to point clouds, testing the 3D perception capability.
-However, active exploration requires access to a simulator, limiting the types of data that can be used, such as real-world videos.
-Therefore, some EQA benchmarks do not involve active exploration.
-EgoVQA [262] shifts the focus of VQA to egocentric videos.
-EgoTaskQA [263] emphasizes spatial, temporal, and causal relationship reasoning.
-EQA-MX [266] focuses on multimodal expressions (MX), including regular verbal utterances and nonverbal gestures like eye gaze and pointing.
-OpenEQA [265] evaluates seven main categories, including functional reasoning and world knowledge, which were not covered in previous benchmarks.
-EgoPlan-Bench [264] and EgoCOT [101] measure the model’s ability to generate task plans, using metrics like accuracy and perlexity.
-PlanBench [268, 269] comprehensively assesses various aspects of task planning ability, such as cost optimality, plan verification, and replanning, etc.
-LoTa-Bench [267] directly evaluates task planning by executing the generated plans in simulators and calculating the success rate.
-V Challenges and Future Directions
-Vision-language-action (VLA) models face several persistent challenges in the domain of Robotics, necessitating focused attention and concerted research efforts:
-Scarcity of Robotic Data.
-Obtaining sufficient real-world robotic data remains a significant obstacle.
-Collecting such data is time-consuming and resource-intensive, while relying solely on simulation data exacerbates the sim-to-real gap problem.
-Diverse real-world robotic datasets require close and extensive collaboration among different institutions.
-Simulated data depends on the development of more realistic and efficient simulators.
-Motion Planning.
-Current motion planning modules often lack the necessary dexterity to address the complexities in various environments.
-This limitation hampers robots’ abilities to interact effectively with tools, navigate complex environments, and execute high-precision operations, etc.
-Overcoming these challenges necessitates the development of more robust motion planning algorithms.
-Real-Time Responsiveness.
-Many robotic applications require real-time decision-making and action execution to meet operational requirements.
-VLA models should be designed to operate responsively, with minimal latency.
-Furthermore, the entire stack of a robot system necessitates global optimization, from the high-level task planner down to the motion planner.
-Integration of Multiple Modalities.
-VLAs must process and integrate information from multiple modalities, including vision, language, and action.
-While significant progress has been made in this regard, achieving optimal integration of these modalities remains an ongoing challenge.
-Addressing this challenge requires advancements in multi-modal representation learning, fusion techniques, and task-specific adaptation.
-Expanding beyond mere visual and linguistic capacities, robots can greatly benefit from incorporating modalities like audio or speech.
-Embracing a broader range of perceptual and communicative abilities empowers robots to engage in more effective collaboration with users.
-Generalization to Unseen Scenarios.
-A truly versatile robotics system should be capable of comprehending and executing natural language instructions across diverse and unseen scenarios.
-Achieving this level of generalization, akin to the capabilities demonstrated by ChatGPT in natural language processing, requires robustness to variations in instructions, environments, objects, and embodiments.
-This necessitates the development of adaptable and scalable VLA architectures.
-Long-Horizon Task Execution.
-Single instructions can often translate into long-horizon tasks for robots, such as the instruction “clean the room” which involves object rearrangement, floor sweeping, table wiping, and more.
-Successfully executing such tasks requires the robot to plan and execute a sequence of low-level actions over extended time horizons.
-While current high-level task planners have achieved initial success, they still fall short in many scenarios because most LLMs are not tuned to embodied tasks.
-Addressing this challenge entails devising efficient planners that possess powerful perception abilities and a wide range of common sense.
-Foundation Model.
-The exploration of foundation models for VLAs in robot tasks remains uncharted territory, primarily due to the diverse embodiments, environments, and tasks encountered in robotics.
-Isolated datasets and evaluation setups further compound this challenge.
-To establish a robust foundation VLA model for robotics, it is imperative to leverage internet-scale embodied datasets and state-of-the-art multi-modal models.
-Multi-agent Systems.
-The large-scale deployment of robots is highly desirable in many scenarios, but it introduces the challenge of multi-agent cooperation.
-Effective communication, task distribution among multiple robots can be difficult [270].
-Fleet heterogeneity may further complicate these issues.
-While multi-agent systems offer benefits like distributed perception, collaborative fault recovery, and autonomous fleet maintenance, realizing the full potential of these advantages remains elusive.
-Benchmarks.
-Although numerous benchmarks exist for evaluating low-level control policy VLAs, they often vary significantly in terms of the skills they assess.
-Additionally, the objects and scenes included in these benchmarks are typically limited by what the simulator can provide.
-To more comprehensively evaluate VLA models, there is a need for benchmarks that cover a diverse set of skills based on realistic simulators.
-For high-level task planner VLAs, many benchmarks claim to measure planning capability, often in the form of question-answering tasks.
-However, it is more desirable to evaluate the high-level task planner together with the low-level control policy to execute long-horizon tasks and measure success rate, rather than relying on isolated measures of the planner alone.
-This approach provides a more holistic assessment of the VLA system’s capabilities.
-Safety Considerations.
-Safety is paramount in robotics, as robots interact directly with the real world.
-Ensuring the safety of robotic systems requires the integration of real-world commonsense and complex reasoning into their development and deployment processes.
-This involves the incorporation of robust safety mechanisms, risk assessment frameworks, and human-robot interaction protocols.
+VI Challenges and Future Directions
+Safety first
+Safety is paramount in robotics, as robots interact directly with the physical world.
+Ensuring the safety of robotic systems requires the integration of real-world commonsense.
+This involves the incorporation of robust safety guardrails, risk assessment frameworks, and human-robot interaction protocols.
+RLHF and “evaluation without execution” can also significantly lower safety risks [37].
 Interpretability and expandability of VLA decision-making processes are also crucial for enhancing robot safety through error diagnosis and troubleshooting.
-Ethical and Societal Implications.
-The deployment of robots has always raised various ethical, legal, and societal concerns.
+Datasets & Benchmarks
+In addition to the issues discussed in §V, comprehensive benchmarks that cover a wide range of skills, objects, embodiments, and environments remain to be developed.
+Moreover, metrics beyond the success rate are needed for a fine-grained diagnosis of issues in VLA models, as highlighted by EAI [222] for LLMs.
+Foundation Models & Generalization
+VLA foundation models or robotic foundation models (RFM) for embodied AI remains an open research topic primarily due to the diversity in embodiments, environments, and tasks.
+Many have made significant progress [139, 147], but still lacks generalization capability on par with LLMs in NLP.
+Attaining such a level of generalization is very challenging, as it requires the development of many core AGI capabilities.
+Multi-Modality
+VLAs inherit many challenges associated with multimodal models, such as obtaining useful embeddings and aligning different modalities.
+Current approaches, like ImageBind [231] and LanguageBind [232], align different modalities to the image or language embedding space, respectively.
+Within a unified embedding space, MLLMs can accommodate diverse modalities and become more general.
+However, whether focusing on embeddings alone is sufficient remains under debate.
+Although beyond the scope of VLAs, other modalities—such as audio [60] and haptics [233]—have proven useful for certain embodied AI applications.
+Incorporating additional modalities is often desirable, but they also introduce extra complexity into the model design.
+Framework for Long-Horizon Tasks
+The hierarchical framework is currently the most practical approach for long-horizon tasks.
+However, it increases system complexity and potential points of failure.
+Frequent task execution failures can trigger re-planning, which can cause significant latency.
+Moreover, monolithic task planners share similar architectures with LVLAs, so employing two large models can be redundant and may hinder scalability.
+Additionally, although modular task planners typically do not require training, they are not plug-and-play: language-based models may generate subtasks that control policies cannot execute, whereas code-based models require modules to be pre-wrapped in APIs manually.
+Therefore, developing a unified framework that directly translates long-horizon tasks into low-level control signals in an end-to-end fashion is worth exploring.
+Real-Time Responsiveness
+Unlike conversational AI, many robotic applications require real-time decision-making to respond to dynamic environments.
+If inference time cannot keep pace with environmental changes, the model may generate obsolete actions repeatedly.
+However, current VLA models—especially LVLAs and task planners—face a trade-off between speed and capacity.
+Novel mechanisms are thus needed to strike an optimal balance.
+Multi-agent Systems
+Cooperative multi-agent systems offer benefits such as distributed perception and collaborative fault recovery.
+However, they also face challenges, including effective communication, coordinated dispatching, and fleet heterogeneity.
+In certain scenarios, individual agents may have conflicting goals, which further increases complexity.
+Ethical and Societal Implications
+Robotics has always raised various ethical, legal, and societal concerns.
 These include risks related to privacy, safety, job displacement, bias in decision-making, and the impact on social norms and human relationships.
-Effective regulation plays a crucial role in promoting the ethical use of robots.
-VI Conclusion
-VLA policy holds immense promise for enabling embodied AI to interact effectively with the world around them.
-Recent advances have demonstrated the capabilities of these models in completing complex tasks under diverse conditions.
-However, significant challenges remain in terms of generalization, efficiency, and safety.
-Further research is needed to address these challenges and pave the way for the widespread adoption of VLA-powered robots in real-world applications.
+Applications
+Most current VLAs focus on household and industrial settings [2], which shape the available datasets and models.
+Given the potential of VLA models, a wide range of new applications is possible.
+One particularly important application is in healthcare, such as surgical robots [234] and care robots [235].
+Healthcare demands higher safety and privacy standards and may necessitate novel techniques such as human-in-the-loop (HITL) control and federated learning.
+Moreover, due to the significant domain gap, specialized vision models might be needed for medical images.
+VII Conclusion
+Vision-language-action models hold immense promise for enabling embodied agents to interact with the physical world and fulfill users’ instructions.
+This paper is the first survey to review large VLAs alongside other generalized VLAs.
+Our taxonomy provides a high-level overview of three main lines of research: key components, control policies, and task planners.
+We meticulously summarize and compare their model designs and technical details.
+Additionally, we highlight essential resources for training and evaluating VLAs, such as datasets, simulators, and benchmarks.
+We hope this survey captures the rapidly evolving landscape of embodied AI and inspires future research.

### v5

@@
 Furthermore, specifying tasks via language instructions offers a more intuitive user-robot interface, which necessitates the development of language-conditioned robot policies.
-Figure 1: The general architecture of vision-language-action models. Important related components are shown in the dashed-line boxes.
+Figure 1: The general architecture of vision-language-action models. Three representative methods for action prediction are shown. Related components are presented in dashed boxes.
 Built upon the success of large VLMs, vision-language-action models have demonstrated their potential in addressing these challenges, as illustrated in Figure 1.
@@
   • Resources.
-We summarize the necessary resources for training and evaluating VLA models, including recent datasets and benchmarks in real-world or simulated environments.
+We summarize the necessary resources for training and evaluating VLAs, including recent datasets and benchmarks in real-world or simulated environments.
 We also discuss various approaches to address current challenges, such as data scarcity and inconsistency.
@@
 Figure 3: The taxonomy of VLA models. The organization of this survey follows this taxonomy.
-II-A Unimodal Models
-Vision-language-action models integrate three modalities, often relying on existing unimodal models.
-The transition from convolutional neural networks (e.g., ResNet [16]) to visual Transformers (e.g., ViT [17], SAM [18]) in computer vision has facilitated the development of vision foundation models (VFMs).
-In natural language processing, the evolution from recurrent neural networks (e.g., LSTM [19], GRU [20]) to Transformers [4] initially led to the pretrain-finetune paradigm (e.g., BERT [21], ChatGPT [1]), followed by the recent success of prompt tuning driven by large language models.
-Reinforcement learning (e.g., DQN [5], AlphaGo [22], PPO [23], Dactyl [24]) has also witnessed a shift towards employing Transformers (§III-A1) to model the Markov Decision Process as autoregressive sequential data.
-II-B Vision-Language Models
-Vision-language tasks, encompassing image captioning [25], visual question answering [26], visual grounding [27], require the fusion of computer vision and natural language processing models.
-Early efforts, such as Show and Tell [28], leveraged the success of early CNNs and RNNs.
-The advent of high-capacity language models, including BERT [21] and GPT [29], ushered in a new era of Transformer-based VLMs.
-One of the pioneering self-supervised pretraining methods is ViLBERT [30], while CLIP [31] popularized contrastive pretraining approaches for multimodal alignment.
-The recent emergence of large language models has driven the development of multimodal LLMs (MLLMs) or large multimodal models (LMMs), which achieve state-of-the-art performance on multimodal instruction-following tasks.
-Representative MLLMs include Flamingo [32], BLIP-2 [7], and LLaVA [8].
-VLMs share a close relationship with VLAs, as the multimodal architectures of VLMs can be readily adopted for VLAs.
-Additionally, VLMs can function as high-level task planners if they possess sufficient reasoning capabilities.
-II-C Embodied AI & Robot Learning
 Embodied AI is a unique form of artificial intelligence that actively interacts with the physical environment.
@@
 III-A1 Reinforcement Learning
-Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them well-suited for Transformer models.
-These RL Transformers have become foundational to many recent VLA models.
+Reinforcement learning (RL) laid the foundation for embodied AI and continues to help advance VLAs.
+Deep Q-Network (DQN) first demonstrated learning policies directly from high-dimensional pixel inputs, underscoring the need for greater model capacity in end-to-end RL.
+RL trajectories—sequences of states, actions, and rewards—naturally align with sequence modeling problems, making them well-suited to transformer architectures.
 Pioneering efforts in this direction include Decision Transformer (DT) [17] and Trajectory Transformer (TT) [18].
-Gato [36] further extends this paradigm to a multi-modal/task/embodiment setting.
-Reinforcement learning from human feedback (RLHF) has become an important training ingredient for LLMs.
-SEED [37] utilizes RLHF alongside skill-based reinforcement learning to address the sparse reward issue of long-horizon tasks in robot learning.
-Reflexion [38] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
+Gato [19] further extended this paradigm to a multimodal, multitask, multi-embodiment setting.
+These methods in RL directly inspired modern VLAs [20, 21], which also process multimodal sequences using transformer-based architectures.
+Furthermore, a synergy has emerged between RL and LLMs, benefiting embodied AI in multiple ways.
+Reinforcement learning from human feedback (RLHF) aligns LLMs with human preferences and has also been applied to robot learning.
+SEED [22] utilizes RLHF alongside skill-based RL to address the sparse reward issue and to improve safety via human evaluation.
+Conversely, LLMs also enable novel RL methods: Reflexion [23] proposes a novel verbal reinforcement learning framework that replaces weight updates in RL models with linguistic feedback, which is naturally applicable to embodied decision making where planning also occurs in language form.
+Eureka [24] shows that LLMs can design reward functions for embodied agents that outperform expert human-engineered ones.
+As LLMs evolve, this synergy accelerates progress in RL and embodied AI.
 III-A2 Pretrained Visual Representations
@@
 For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc.
-$\mathcal{S}(\cdot)$ is some similarity measurement.
-Model V-Encoder Type Objective Notation Robotic Tasks
-CLIP [31] ViT-B [17] VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair (Used by CLIPort [43], EmbCLIP [44], and CoW [45])
-R3M [39] ResNet-50 [16] Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Sim-Mani: Meta-World, Franka Kitchen, Adroit
-MVP [41] ViT-B
-ViT-L
-MAE [46] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch Real-Mani (xArm7): pick, reach block, push cube, close fridge
+$\mathcal{S}(\cdot)$ denotes a similarity measure.
+Model Network Type Objective Notations Robotic Tasks
+CLIP [25] ViT-B VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair (Used by CLIPort [31], EmbCLIP [32], and CoW [33])
+R3M [26] ResNet-50 Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Sim-Mani: Meta-World, Franka Kitchen, Adroit
+MVP [28] ViT-B/L MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch Real-Mani (xArm7): pick, reach block, push cube, close fridge
 VIP [27] ResNet-50 Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{0},x_{j})\right)}{\exp\left(\mathcal{S}(x_{0},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{j})\right)}$ $x_{i}$: $i$-th video frame, $0<i<j$ Real-Mani (Franka): pick and place, close drawer, push bottle, fold towel
-VC-1 [47] ViT-L MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP.) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0; Sim-Navi: Habitat
-Voltron [48] TFM [4] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
+VC-1 [34] ViT-L MAE, CL (A combination of previous works, including CLIP, R3M, MVP, and VIP.)   Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0; Sim-Navi: Habitat
+Voltron [35] TFM MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
 $\sum_{i=1}^{N^{\prime}}-\log P\left(y_{i}\mid x,y_{<i}\right)$
@@
 Real-Mani (Franka): custom study desk
-RPT [42] ViT [17, 41] MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x,y,z$: three distinct modalities Real-Mani (Franka): stack, pick, pick from bin
+RPT [30] ViT MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x,y,z$: three distinct modalities Real-Mani (Franka): stack, pick, pick from bin
 DINOv2 [36] ResNet, ViT Self-distillation $\sum_{x}\sum_{x^{\prime}\neq x}H(P_{t}(x),P_{x}(x^{\prime}))$ $x,x^{\prime}$: image views; $H()$: cross-entropy; $P_{t},P_{s}$: teacher, student (Used by OpenVLA [37], ReKep [38])
-I-JEPA [52] ViT JEPA $\frac{1}{M}\sum_{i=1}^{M}\sum_{j}\left\lVert x_{j}(i)-y_{j}(i)\right\rVert_{2}^{2}$ $x_{j}(i)$: $j$-th masked patch embedding of block $i$; $y_{j}(i)$: unmasked patch embedding -
-Theia [53] ViT-T/S/B Distillation (Distillation of vision foundation models, including ViT, CLIP, SAM, DINOv2, and Depth-Anything.) Sim-Mani&Navi: CortexBench [47]; Real-Mani: pick, place, open door/drawer
+I-JEPA [39] ViT JEPA $\frac{1}{M}\sum_{i=1}^{M}\sum_{j}\left\lVert x_{j}(i)-y_{j}(i)\right\rVert_{2}^{2}$ $x_{j}(i)$: $j$-th masked patch embedding of block $i$; $y_{j}(i)$: unmasked patch embedding
+Theia [40] ViT-T/S/B Distillation (Distillation of vision foundation models: ViT, CLIP, SAM, DINOv2, Depth-Anything.)   Sim-Mani&Navi: CortexBench (VC-1); Real-Mani: pick, place, open door/drawer
 Voltron [35] introduces a pretraining objective by incorporating language conditioning and language generation into the MAE objective.
@@
 Table II: Dynamics learning methods for VLAs. $f(\cdot)$ is the dynamic model. Fwd, inv: forward & inverse dynamics.
-Model V-Encoder Type Objective Notation Robotic Tasks
+Model Vision Encoder Type Objective Notations Robotic Tasks
 Vi-PRoM [48] ResNet Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$: predicts frame index in raw seq.
 $x$ given shuffled seq.
-$x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open and close drawer/door
-MaskDP [62] ViT-MAE [46] MAE (implicit fwd & inv dynamics) $\sum_{i=1}^{T}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: state or action token; $y$: the other modality Sim: DeepMind Control Suite
-MIDAS [63] ViT, Mask R-CNN [64] Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: VIMA-Bench
+$x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open & close drawer/door
+MaskDP [49] ViT (MAE) MAE (implicit fwd & inv dynamics) $\sum_{i=1}^{T}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: state or action token; $y$: the other modality Sim: DeepMind Control Suite
+MIDAS [50] ViT, Mask R-CNN Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: VIMA-Bench
 SMART [51] CNN Forward & inverse dynamics $\sum_{i=1}^{T}\big{(}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$
-$+\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)\big{)}$
+    $\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)\big{)}$
 $s_{t}$, $a_{t}$: state, action Sim: DeepMind Control Suite
-PACT [66] ResNet-18, PointNet [67] Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Navi: Habitat; Real-Navi (MuSHR vehicle)
+PACT [52] ResNet-18, PointNet Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Navi: Habitat; Real-Navi (MuSHR vehicle)
 VPT [53] ResNet Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim: Minecraft
-GR-1 [69] ViT-MAE Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: CALVIN
+GR-1 [54] ViT (MAE) Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: CALVIN
 Vi-PRoM [48] presents three distinct pretraining objectives.
@@
 In embodied AI, researchers are exploring how to leverage CoT reasoning to refine the decision-making process.
-ThinkBot [90] applies it to recover missing action descriptions in sparse human instructions.
-RAT [91] combines retrieval-augmented generation and CoT to elicit context-aware reasoning, improving long-horizon generation.
-ReAct [92] interleaves verbal reasoning traces and actions to allow for diverse types of reasoning.
-Tree-Planner [80] employs a tree-of-thoughts approach.
-ECoT [93] re-trains OpenVLA [50] using reasoning chain datasets, enabling embodied CoT reasoning that further enhances its performance.
-Pros and Cons.
+Reasoning is naturally compatible with high-level task planning, as both often occur in the language domain.
+ThinkBot [75] applies CoT to recover missing action descriptions in sparse human instructions, thereby enhancing instruction coherence and boosting success rates in difficult tasks.
+ReAct [76] interleaves reasoning traces and actions, where CoT can help create action plans, inject commonsense knowledge, and handle exceptions, ultimately improving embodied decision-making.
+RAT [77] integrates CoT with retrieval-augmented generation (RAG) to mitigate hallucinations and thus improve embodied planning.
+Tree-Planner [65] employs a tree-of-thoughts approach for task planning.
+Another paradigm involves equipping low-level control policies with reasoning capabilities, as pioneered by ECoT [78].
+This method trains OpenVLA [37] to conduct embodied CoT reasoning about plans, sub-tasks, motions, and visual features, before predicting low-level actions.
+By relying on this multi-step reasoning rather than the “muscle memory” of VLAs, it improves success rates on challenging generalization tasks without requiring additional robot data.
+Strengths and Limitations.
 PVRs
@@
 In contrast, CoT-based reasoning methods focus on task planning since they express thought chains in text, making them well-suited for refining text-based task plans.
-Figure 4: Illustration of a hierarchical robot policy. The high-level task planner decomposes the user instruction into subtasks, which are then executed step by step by the low-level control policy.
-Table III: Low-level control policies.
-We also include some non-VLA models as they are closely related, marked in gray text.
-BC: behavior cloning (action type, cont/disc: continuous/discrete).
-TFM: Transformer.
-Xattn: cross-attention.
-Concat: concatenation.
-$p$/$s$: prompt/state vision encoder.
-[SC]: self-collect data.
-MPC: model predictive control.
-Model V-Encoder L-Encoder VL-Fusion Action-Decoder Objective Training Dataset Robotic Tasks
-CLIPort [43] CLIP-ResNet50, Transporter-ResNet CLIP-GPT Hadamard LingUNet BC ($\textbf{SE}(2)$) [SC] Sim: Ravens [94]
-MCIL [95] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP, [SC] Sim: 3D Playroom environment
-HULC [96],
-HULC++ [97]
-MCIL CNN Sentence-BERT LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
-Language costs [98] CLIP-ViT,
-UNet [99]
-CLIP-GPT Concat STORM Max Likelihood Est. (cost map) [SC] Sim & Real (Franka): pick, place
-Interactive Language [100] ResNet CLIP-GPT Xattn TFM BC (cont) Language-Table [SC] Sim & Real (xArm6): rearrangement
-Hiveformer [101] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data Sim: RLBench [102]
-PerAct [103] 3D CNN CLIP-GPT Concat PerceiverIO 3D affordance RLBench data Sim: RLBench; Real (Franka): put marker, place food, etc.
-Act3D [104] CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
-RVT, RVT-2 [105, 106] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data, [SC] Sim: RLBench; Real (Franka): stack, press, place
-RoboPoint [107] ViT-L/14 Vicuna-V1.5 13B [108] Concat - 2D affordance (project to 3D) [SC] Real (Franka): pick, place
-Gato [36] ViT Sent.Piece Concat TFM BC (cont & disc) [SC] Sim & Real (Sawyer): RGB-stacking
-RoboCat [109] VQ-GAN ($p,s$) [110] - - TFM BC, Observation prediction Self-improve Sim & Real (Sawyer, Franka, KUKA): stacking, building, lifting, insertion, removal
-VIMA [111] ViT, Mask R-CNN [64] T5 [112] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data [SC] Sim (Ravens): VIMA-Bench
-BC-Z [113] ResNet18 ($p,s$) USE FiLM MLP BC (cont) [SC] Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
-RT-1 [114] EfficientNet [115] USE FiLM TFM BC (disc) RT-1-Kitchen [SC] Real (Everyday Robots): pick-place, move, knock
-MOO [111] OWL-ViT ($p$) [116], EfficientNet ($s$) USE FiLM TFM BC (disc) [SC] Real (Everyday Robots): pick, move near, knock, place upright, place into
-Q-Transformer [117] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1-Kitchen Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
-RT-Trajectory [118] EfficientNet - - TFM BC (disc) [SC] Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
-ACT [119] ResNet18 - - CVAE-TFM BC (cont, action chunking) [SC] with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
-MT-ACT [120] (RoboAgent) CNN - FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
-RoboFlamingo [121] CLIP-ViT-L/14 LLaMA [122], MPT [123] Xattn LSTM, TFM BC (cont) CALVIN data Sim: CALVIN
-RoboUniView [124] UVFormer (Model design follows RoboFlamingo) BC (cont) CALVIN data Sim: CALVIN
-VoxPoser [125] ViLD [126],
-MDETR [127],
-OWL-ViT [116],
-SAM [18]
-GPT-4 Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
-UniPi [128] Imagen video [129] T5-XXL Xattn CNN Inverse dynamics PF [SC], Ravens, BridgeV1 Sim: Painting Factory (PF), Ravens
-Diffusion Policy [130] ResNet18 - - UNet,
-TFM
-DDPM [131] [SC] Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
-DP3 [132] DP3 Encoder (Model design follows Diffusion Policy) DDIM [SC] Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
-SUDD [133] ResNet18 CLIP-GPT Concat UNet,
-TFM
-DDPM Lang-Guided Data Gen Sim: MuJoCo; Real (UR5e): pick, place
-Octo [134] CNN T5-base Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
-3D Diffuser Actor [135] (Model design follows Act3D) DDPM RLBench, CALVIN, [SC] Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
-MDT [136] CLIP-ViT-B/16 ($p$), Voltron/ResNet18 ($s$) CLIP-GPT Concat DiT [137] DDIM [138] CALVIN, LIBERO, [SC] Sim: CALVIN, LIBERO; Real (Franka): pick, push, open, close, etc.
-RDT-1B [139] SigLIP T5-XXL Xattn DiT DDPM (Aggregated Datasets) Real (ALOHA dual-arm): wash, pour, fold, etc.
-RT-2 [2] ViT-22B [140],
-ViT-4B [141]
-PaLI-X, PaLM-E Concat Symbol-tuning BC (disc), Co-fine-tuning VQA, RT-1-Kitchen Real: RT-1 evaluation tasks, Language-Table
-RT-H [142] ViT-22B [140] PaLI-X 55B Concat Symbol-tuning BC (disc) Diverse+Kitchen Real: Diverse+Kitchen eval tasks
-RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE [SC] Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
-OpenVLA [50] DINOv2, SigLIP [144] Prismatic-7B [145] Concat Symbol-tuning BC (disc) OXE, DROID [146] Real: BridgeV2, RT-1 evaluation tasks
-$\pi_{0}$ [147] SigLIP PaliGemma Concat Action expert Flow matching OXE, $\pi$-cross-embod. [SC] Real (general robot control): sweep, open, pack, etc.
 III-B Low-level Control Policies
@@
 $\begin{split}\hat{a}_{t}\sim\pi_{\theta}(\hat{a}_{t}|p,s_{\leq t},a_{<t}).\end{split}$ (3)
-The diversity among control policy networks arises from individual modules and overall architectures.
-The most typical architecture is shown in Figure 1.
+It can also be referred to as a low-level policy, low-level controller, or action primitive.
+The diversity among VLAs arises from individual modules and overall architectures.
+The general architecture is shown in Figure 1.
 This section explores various approaches to designing low-level control policies.
 Table III summarizes their technical details.
+Figure 4: Illustration of a hierarchical robot policy. The high-level task planner decomposes the user instruction into subtasks, which are then executed step by step by the low-level control policy.
 III-B1 Non-Transformer Control Policies
 Prior to the adoption of Transformer models, early control policies for language-conditioned robotic tasks varied significantly in architecture.
-CLIPort [43] integrates the vision and language encoders of CLIP [31] with the Transporter network [94], creating a two-stream architecture.
-In one stream, CLIP’s vision encoder extracts “semantic” information from the RGB image, while in the other stream, the Transporter network extracts “spatial” information from the RGB-D image.
-The CLIP sentence encoder encodes the language instruction and guides the output action, which is a pair of end-effector poses: the picking and placing poses.
-CLIPort demonstrates the ability to pick and place objects as specified by the language instruction.
+CLIPort [31] integrates CLIP with the Transporter network, creating a two-stream architecture.
+The CLIP vision encoder extracts “semantic” information from the RGB image, while the Transporter network extracts “spatial” information from the RGB-D image.
+The CLIP sentence encoder encodes the language instruction and guides the output $SE(2)$ action, consisting of paired pick and place end-effector poses.
+It represents an early demonstration of language-conditioned pick-and-place capabilities.
 BC-Z [79] processes two types of task instructions: a language instruction or a human demonstration video.
 The environment is presented to the model in the form of an RGB image.
-Then the instruction embedding and the image embedding are combined through the FiLM layer [148], culminating in the generation of actions.
+Then the instruction embedding and the image embedding are combined through the FiLM layer, culminating in the generation of actions.
 This conditional policy is asserted to exhibit zero-shot task generalization to unseen tasks.
@@
 This innovative policy-as-video formulation offers several advantages, including enhanced generalization across diverse robot tasks and the potential for knowledge transfer from internet videos to real robots.
+Table III: VLAs that serve as low-level control policies.
+$\lozenge$ indicates large VLAs.
+Closely related (non-VLA models) are included in brackets.
+The remaining are generalized VLAs.
+BC: behavior cloning (cont/disc: continuous/discrete action).
+Xattn: cross-attention.
+Concat: concatenation.
+Quant.: quantization.
+pp/$s$: prompt/state vision encoder.
+[SC]: self-collect data.
+Model Vision
+Encoder
+Language Encoder / VLM Action Decoder/Head Architecture Training Objectives Training Datasets Environments, Embodiments, Tasks, and Skills
+CLIPort [31] CLIP-ResNet50, Transporter-ResNet CLIP-GPT LingUNet Hadamard BC ($SE(2)$) [SC] Sim: Ravens
+MCIL [80] Custom-CNN USE RNN LMP MCIL Play-LMP, [SC] Sim: 3D Playroom environment
+HULC [81, 82] MCIL CNN Sentence-BERT RNN LMP MCIL CALVIN data Sim: CALVIN
+Language costs [84] CLIP-ViT, UNet CLIP-GPT STORM Concat MLE (cost map) [SC] Sim & Real (Franka): pick, place
+Interactive Language [85] ResNet CLIP-GPT TFM Xattn BC (cont) [SC: Language-Table] Sim & Real (xArm6): rearrangement
+Hiveformer [86] UNet CLIP-GPT TFM, UNet Concat BC (cont) RLBench data Sim: RLBench
+PerAct [87] 3D CNN CLIP-GPT PerceiverIO Concat 3D affordance RLBench data, [SC] Sim: RLBench; Real (Franka): pick, place, stack, open drawer, sweep, insert peg, etc.
+Act3D [88] CLIP-ResNet50 CLIP-GPT TFM Xattn 3D affordance (voxel) RLBench data, [SC] Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
+RVT [89],
+RVT-2 [90]
+CLIP-ResNet50 CLIP-GPT TFM Concat 2D affordance (project to 3D) RLBench data, [SC] Sim: RLBench; Real (Franka): stack, press, place
+RoboPoint [91] ViT-L/14 Vicuna-V1.5 13B   Concat 2D affordance (project to 3D) [SC] Real (Franka): pick, place
+Gato [19] ViT Sent.Piece TFM Concat BC (cont & disc) [SC] Sim & Real (Sawyer): RGB-stacking
+(RoboCat) [92] VQ-GAN ($p,s$)   TFM Quant. BC, observation prediction Self-improvement Sim & Real (Sawyer, Franka, KUKA): stacking, building, lifting, insertion, removal
+VIMA [93] ViT, Mask R-CNN T5 TFM Xattn BC ($SE(2)$) [SC:VIMA-Data] Sim (Ravens): VIMA-Bench
+BC-Z [79] ResNet18 ($p,s$) USE MLP FiLM BC (cont) [SC] Real (EDR): pick-place/wipe/drag, grasp, push
+RT-1 [94] EfficientNet USE TFM FiLM BC (disc) [SC: Fractal] Real (EDR): pick-place, move, knock
+MOO [93] OWL-ViT ($p$), EfficientNet ($s$) USE TFM FiLM BC (disc) [SC] Real (EDR): pick, move near, knock, place upright, place into
+Q-Transformer [95] EfficientNet USE TFM FiLM TD error Fractal, Auto-collect Sim: pick; Real (EDR): pick, place, open/close drawer, move near
+(RT-Trajectory) [96] EfficientNet   TFM   BC (disc) [SC] Real (EDR): pick, place, fold towel, swivel chair, etc.
+(ACT) [97] ResNet18   CVAE-TFM   BC (cont, action chunking) [SC] with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
+RoboAgent / MT-ACT [98] CNN   CVAE-TFM FiLM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
+RoboFlamingo [99] CLIP-ViT-L/14 LLaMA, MPT, GPT-NeoX LSTM, TFM Xattn BC (cont) CALVIN data Sim: CALVIN
+RoboUniView [100] UVFormer (Model design follows RoboFlamingo) BC (cont) CALVIN data Sim: CALVIN
+DeeR-VLA [101] (Model design follows RoboFlamingo++) BC (cont) CALVIN data Sim: CALVIN
+Instruct2Act [102] CLIP, SAM ChatGPT Robot APIs Tool-Use     Sim (Ravens): VIMA-Bench
+VoxPoser [103] ViLD, MDETR, OWL-ViT, SAM GPT-4 MPC Tool-Use     Sim: Sapien; Real (Franka): move&avoid, set table, close drawer, open bottle, sweep trash
+UniPi [83] Imagen video T5-XXL CNN Xattn Inverse dynamics [SC:PF], Ravens, BridgeV1 Sim: Painting Factory (PF), Ravens
+(Diffusion Policy) [104] ResNet18   UNet, TFM   DDPM [SC] Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
+(DP3) [105] DP3 Encoder (Model design follows Diffusion Policy) DDIM [SC] Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
+SUDD [106] ResNet18 CLIP-GPT UNet, TFM Concat DDPM Lang-guided data generation Sim: MuJoCo; Real (UR5e): pick, place
+Octo [107] CNN T5-base TFM Concat DDPM OXE Real: BridgeV2, CMU Baking, Stanford Coffee, Berkeley Peg Insert, etc.
+3D Diffuser Actor [108] (Model design follows Act3D) DDPM RLBench, CALVIN, [SC] Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
+MDT [109] CLIP-ViT-B/16 ($p$), Voltron/ResNet18 ($s$) CLIP-GPT DiT Concat DDIM CALVIN, LIBERO, [SC] Sim: CALVIN, LIBERO; Real (Franka): pick, push, open, close, etc.
+RDT-1B [110] SigLIP T5-XXL DiT Xattn DDPM (Aggregated Datasets) Real (ALOHA dual-arm): wash, pour, fold, etc.
+RT-2 [2] $\lozenge$ ViT-4B, ViT-22B PaLI-X, PaLM-E Symbol-tuning Concat BC (disc), Co-fine-tuning Fractal, VQA Sim: Language-Table; Real: RT-1 evaluation tasks
+RT-H [111] $\lozenge$ (Model design follows RT-2) BC (disc) Diverse+Kitchen Real: Diverse+Kitchen eval tasks
+RT-X [112] $\lozenge$ (Models from RT-1 and RT-2) BC (disc) [SC: OXE] Real: BridgeV2, RT-1 evaluation tasks, etc.
+OpenVLA [37] $\lozenge$ DINOv2, SigLIP Prismatic-7B Symbol-tuning Concat BC (disc) OXE, DROID Real: BridgeV2, RT-1 evaluation tasks, Franka-Tabletop, DROID, etc.
+OpenVLA-OFT [113] $\lozenge$ (Improves OpenVLA with OFT recipe) BC (cont, parallel decode w/ chunk.) LIBERO, [SC] Sim: LIBERO; Real (ALOHA setup): fold, scoop, put
+TraceVLA [114] $\lozenge$ (Model design follows OpenVLA, adding visual trace prompting) BC (disc) BridgeV2, Fractal, [SC] Sim: SimplerEnv; Real (WidowX): pick, push, fold, swipe
+$\pi_{0}$ [115] $\lozenge$ SigLIP PaliGemma Action expert Concat Flow matching OXE, [SC: $\pi$-cross-embod.] Real (general robot control): sweep, open, pack, etc.
+RoboMamba [116] $\lozenge$ CLIP-ViT Mamba MLP Concat BC (cont) [SC] Sim: Sapien; Real (Franka): open door/box, staple
+SpatialVLA [117] $\lozenge$ SigLIP, Ego3D Position Encoding PaliGemma 2 MLP Concat BC (Adaptive Action Grids) OXE, BridgeV2, Fractal, RH20T Sim: SimplerEnv, LIBERO; Real (WidowX): pick, place, close, push
+TinyVLA [118] $\lozenge$ ViT Pythia Diffusion Policy Concat DDPM OXE, [SC] Sim: MetaWorld; Real (Franka, UR5): place, stack, flip mug, close drawer, open box
+CogACT [119] $\lozenge$ DINOv2, SigLIP LLaMA 2 DiT Concat DDIM OXE, [SC] Sim: SimplerEnv; Real (Realman, Franka): pick, place, stack, open/close oven
+DexVLA [20] $\lozenge$ ViT, ResNet-50 Qwen2-VL 2B, DistilBERT, ScaleDP Concat DDPM [SC] Real (Franka, UR5e, AgileX): pick, fold shirt, bus table, pour
+HybridVLA [120] $\lozenge$ DINOv2, SigLIP, CLIP LLaMA 2, Phi-2 MLP Concat DDIM OXE, DROID Sim: RLBench; Real (Franka, AgileX): pick-place, pour, open drawer, fold, etc.
+LAPA [121] $\lozenge$ C-ViViT, VQ-GAN LWM-Chat-1M MLP Concat LAPA OXE, BridgeV2, Something V2 Sim: Language-Table, SimplerEnv; Real (Franka): pick, cover with towel, knock over
+WorldVLA [122] $\lozenge$ VQ-GAN Chameleon TFM Quant. BC (disc, chunk.), world model LIBERO, OpenVLA data Sim: LIBERO
+UniVLA [123] $\lozenge$ MoVQ Emu3 TFM Quant. BC (disc, chunk.), world model (Sim benchmark data) Sim: CALVIN, LIBERO, SimplerEnv
 III-B2 Transformer-based Control Policies
@@
 Additionally, RoboAgent introduces a semantic augmentation method that leverages inpainting to augment existing demonstrations.
-RoboFlamingo [121] adapts the existing VLM, Flamingo [32, 149], to a robot policy by attaching an LSTM-based policy head to the VLM.
+RoboFlamingo [99] adapts the existing Flamingo VLM to a robot policy by attaching an LSTM-based policy head to the VLM.
 This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
+(a) FiLM
+(b) Cross-attention
+(c) Concatenation
+(d) Quantization
+(e) Tool Use
+Figure 5: Representative architectures of VLA models. $\odot$: Hadamard product. $\oplus$: Concatenation.
 III-B3 Control Policies for Multimodal Instructions
-Multimodal instruction enables new ways to specify tasks, such as through demonstrations, by naming novel objects, or by pointing with a finger.
+Multimodal instruction enables new ways to specify tasks, such as through demonstrations, by naming novel objects, or by pointing with a finger or mouse click.
 VIMA [126] places a significant emphasis on multimodal prompts and the generalization capabilities of models.
@@
 MOO [93] extends RT-1 to handle multimodal prompts.
-Leveraging the backbone of RT-1, MOO incorporates OWL-ViT to encode images within the prompt.
+Leveraging the backbone of RT-1, it incorporates OWL-ViT to encode images within the prompt.
 By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1.
@@
 III-B4 Control Policies with 3D Vision
-We live in a 3D world, and it is intuitive that using 3D representations as visual inputs should provide richer information than 2D images.
-Point clouds are a popular choice for representing 3D inputs due to their straightforward derivation from RGBD inputs, as demonstrated by both DP3 and 3D Diffuser Actor.
-However, voxels have also been explored in various works.
-RoboUniView [124] shows improved performance by injecting 3D information into RoboFlamingo through a novel UVFormer module as its vision encoder, which provides 3D occupancy information from multi-perspective images.
-VER [151] also proposes voxelizing multi-view images into 3D cells in a coarse-to-fine manner, enhancing performance in vision-language navigation tasks.
-PerAct [103] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations.
-This approach offers a robust structural prior for action learning, enabling natural processing of multi-view observations and facilitating data augmentation in 6-DoF.
-In this framework, the input to the model comprises voxel maps reconstructed from RGBD images, while the output corresponds to the best voxel for guiding the gripper’s movement.
-By adopting this formulation, PerAct facilitates efficient task learning even with a few demonstrations.
-On the other hand, Act3D [104] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization.
+In our 3D world, it is reasonable to assume that 3D vision provides richer information than 2D images.
+Point clouds are a common representation for 3D visual inputs due to their straightforward derivation from RGB-D inputs, as demonstrated by both DP3 [105] and 3D Diffuser Actor [108].
+Voxels have likewise been extensively studied.
+VER [127] proposes voxelizing multi-view images into 3D cells in a coarse-to-fine manner, enhancing performance in vision-language navigation tasks.
+PerAct [87] facilitates efficient task learning with only a few demonstrations by leveraging voxel representations in both the observation and action spaces.
+The input comprises voxel maps reconstructed from multi-view RGB-D images, while the output corresponds to the best voxel for guiding the gripper’s movement.
+RoboUniView [100] improves performance by injecting 3D information from multi-perspective images through a novel UVFormer vision encoder, which is pretrained on a 3D occupancy task.
+In contrast, Act3D [88] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization.
 RVT, RVT-2 [89, 90] propose re-rendering images from virtual views of the scene point cloud and using these images as inputs, rather than directly relying on 3D inputs.
@@
 Octo [107] introduces a Transformer-based diffusion policy characterized by a modular open-framework design, allowing for flexible connections from different task definition encoders, observation encoders, and action decoders to the Octo Transformer.
-Being among the first to utilize the Open X-Embodiment dataset [143], Octo demonstrates positive transfer and generalizability across diverse robots and tasks.
+Being among the first to utilize the OXE dataset [112], Octo demonstrates positive transfer and generalizability across diverse robots and tasks.
 MDT [109] adapts the newly introduced DiT model [129] from computer vision to the action prediction head.
@@
 Similarly, 3D Diffuser Actor [108] shares the core idea of DP3 but differs in model architecture by combining Act3D with Diffusion Policy.
+3D-MoE [130] explores an efficient mixture-of-experts architecture with DiT-based action diffusion using a rectified flow scheduler.
 III-B7 Control Policies for Motion Planning
@@
 The resulting models, RT-1-X and RT-2-X, both outperform their original versions.
-OpenVLA [50] was later developed as an open-source counterpart to RT-2-X. They additionally explored efficient fine-tuning methods, including LoRA and model quantization.
+OpenVLA [37] was later developed as an open-source counterpart to RT-2-X.
+They additionally explored efficient fine-tuning methods, including LoRA and model quantization.
+OpenVLA-OFT [113] applies an Optimized Fine-Tuning (OFT) recipe for improved efficiency and performance.
+TraceVLA [114] finetunes OpenVLA to enable visual trace prompting, enhancing spatial-temporal awareness.
 $\pi_{0}$ [115] proposes a flow-matching architecture for transforming VLMs into VLAs.
 By incorporating an additional action expert based on the mixture-of-experts framework, it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
-(a) FiLM
-(b) Cross-attention
-(c) Concatenation
-Figure 5: Three common vision-language fusion methods of Transformer-based control policies.
-FiLM layers (Hadamard product $\odot$) are used in RT-1 models.
-Some utilize cross-attention to condition on instructions.
-Concatenation ($\oplus$) is the prevailing method in LVLAs.
-Pros and Cons.
+RoboMamba [116] replaces the costly Transformer with a Mamba state space model that has linear inference complexity, achieving efficient robotic reasoning and action capabilities.
+SpatialVLA [117] introduces Ego3D Position Encoding for injecting 3D information and represents actions with Adaptive Action Grids for enhanced generalizability and robustness.
+LAPA [121] devises the first unsupervised pretraining method for VLAs based on latent actions [69].
+This approach employs a three-stage process to learn from internet-scale unlabeled videos.
+First, a VQ-VAE framework is pretrained to extract quantized latent actions between image frames.
+Next, a VLA model is pretrained to predict these latent actions.
+Finally, the model is finetuned using only a small robotic dataset to map the latent actions to actual robot actions.
+The integration of large VLAs with diffusion has also been gaining popularity.
+TinyVLA [118] leverages the Diffusion Policy, while CogACT [119] utilizes a DiT action diffusion module.
+DexVLA [20] proposes embodied curriculum learning to progressively train a diffusion action expert, incorporating sub-step reasoning for decomposing long-horizon tasks.
+HybridVLA [120] integrates diffusion with the autoregression paradigm to fully leverage VLMs’ reasoning capabilities.
+Visual autoregressive modeling (VAR) [133] with quantized visual tokens demonstrates improved performance over diffusion models in image generation.
+This suggests that the three modalities of VLAs can be unified under the autoregressive paradigm [134].
+WorldVLA [122] and UniVLA [123] advance this direction by integrating VLAs with world models.
+It quantizes multimodal data into discrete tokens, forming a shared vocabulary of quantized multimodal tokens.
+Consequently, all modalities can be modeled autoregressively, enabling not only action and text generation but also image generation, thereby constituting a world model.
+A recent trend in LLMs is equipping them with tool-use capabilities by generating code that calls tools via APIs [135].
+Instruct2Act [102] follows this paradigm by integrating vision and action tools, enabling LLMs to perform robotic tasks.
+Strengths and Limitations.
 Architectures
-Various VLA architectures explore different methods of fusing vision and language inputs, including cross-attention, FiLM, and concatenation, as illustrated in Figure 5.
-FiLM is used in RT-1 and thus its follow-up works inherit this mechanism.
+Representative VLA architectures are illustrated in Figure 5.
+FiLM is used in RT-1 and thus its follow-up models inherit this mechanism.
 While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [126].
+Quantization unifies multimodal tokens into a shared vocabulary, thus enabling integration with world models.
+The tool-use paradigm of LLMs can also be applied to robotic tasks.
 Action types and their training objectives
@@
 CLIPort [31] and VIMA [126] use $SE(2)$ action and its BC objective can be expressed as follows:
-$\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}})\end{split}$ (6)
+$\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}}).\end{split}$ (6)
 The DDPM objective in diffusion-based control policies [104] is represented as:
-$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right)$ (7)
+$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right),$ (7)
 where $\varepsilon^{k}$ is a random noise for iteration $k$; $\varepsilon_{\theta}$ is the noise prediction network, i.e., the VLA model.
@@
 Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner [38].
-Table IV: High-level task planners. VL: vision-language fusion method.
-Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
-SayCan [10] - PaLM, FLAN - BC-Z, MT-Opt No Real-Navi&Mani (Everyday): office kitchen
+Table IV: High-level task planners.
+Model Vision
+Model
+Language Model / VLM Low-level Control Policy Architecture Require Training Environments, Embodiments, Tasks, and Skills
+PaLM-E [11] ViT, OSRT PaLM Interactive Language, RT-1 Concat Sim: TAMP; Sim&Real-Mani: Language-Table; Real: SayCan setup
+EmbodiedGPT [136] EVA ViT-G/14 LLaMA-7B MLP Xattn Sim-Mani: Franka Kitchen, Meta-World
+LEO [137] ConvNeXt, PointNet++ Vicuna-7B MLP Concat Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
+3D-LLM [44] CLIP, 3D-CLR, ConceptFusion Flamingo, BLIP-2 $D$-PPO Concat Sim-Navi: Habitat
+ShapeLLM [138] ReCon++ LLaMA   Concat 3D MM-Vet Benchmark
+SayCan [10]   PaLM, FLAN BC-Z, MT-Opt   Real-Navi&Mani (EDR): office kitchen
 […]
-(SL)3 [155] - T5-small - seq2seq Yes Sim-Navi: ALFRED
-EmbodiedGPT [156] EVA ViT-G/14 [157] LLaMA-7B Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
-PaLM-E [11] ViT, OSRT [158] PaLM Concat Interactive Language, RT-1 Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
-LEO [159] ConvNeXt [160], PointNet++ [161] Vicuna-7B Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
-3D-LLM [57] CLIP, 3D-CLR [162], ConceptFusion [163] Flamingo, BLIP-2 Concat DD-PPO [164] Yes Sim-Navi: Habitat
-ShapeLLM [165] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
-Inner Monologue [9] MDETR, CLIP InstructGPT, PaLM Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
-LLM-Planner [166] HLSM GPT-3 Language HLSM Yes Sim-Navi: ALFRED
-LID [167] CNN GPT-2 Language & Embed - Yes Sim: VirtualHome
-SMs [168] CLIP, ViLD GPT-3, RoBERTa [169] Language & Code CLIPort No Sim-Mani: pick, place
-ConceptGraphs [170] SAM, CLIP, LLaVA [8] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
-ProgPrompt [171] ViLD GPT-3 Code API No Sim: VirtualHome
-ChatGPT for Robotics [172] YOLOv8 [173] ChatGPT Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
-CaP [174] ViLD, MDETR GPT-3, Codex Code API No Sim-Mani & Sim-Navi: pick, place, etc.
-DEPS [175] CLIP ChatGPT Code MC-Controller, Steve-1 No Sim: Minecraft
+(SL)3 [140]   T5-small seq2seq   Sim-Navi: ALFRED
+Inner Monologue [9] MDETR, CLIP InstructGPT, PaLM CLIPort, heuristics, SayCan policies Language Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
+LLM-Planner [141] HLSM object detector GPT-3 HLSM Language Sim-Navi: ALFRED
+LID [142] CNN GPT-2   Lang., Embed Sim: VirtualHome
+SMs [143] CLIP, ViLD GPT-3, RoBERTa CLIPort Lang., Code Sim-Mani: pick, place
+ProgPrompt [144] ViLD GPT-3 API Code Sim: VirtualHome
+ChatGPT for Robotics [145] YOLOv8 ChatGPT API Code Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
+CaP [146] ViLD, MDETR GPT-3, Codex API Code Sim-Mani & Sim-Navi: pick, place, etc.
+DEPS [147] CLIP ChatGPT MC-Controller, Steve-1 Code Sim: Minecraft
+ConceptGraphs [148] SAM, CLIP, LLaVA GPT-4 API Code (JSON) Sim: AI2-THOR; Real (Spot Arm): pick-place
 RT series
@@
 Based on RT-2, RT-H [111] introduces action hierarchies for better data sharing.
-LVLA vs VLA
+LVLA vs generalized VLA
 While LVLAs can greatly enhance instruction-following abilities because they can better parse user intentions, concerns arise regarding their training cost and deployment speed.
@@
 DeeR-VLA [101] proposes only partially activating the model with dynamic inference with early-exit.
+Scaling Law
+Similar to LLMs, scaling laws have also been observed in robotics [149], revealing the importance of model size, data quality, and diversity in environments and objects.
+Further research in this area can guide the development of VLAs with improved in-the-wild generalization capabilities.
 IV Task Planners
-A high-level task planner $\pi_{\phi}$ aims to divide a complex task $\ell$ into a sequence of subtasks $p_{i}$ (i.e., task plan), each serving as a language instruction to low-level control policies $\pi_{\theta}$:
+A high-level task planner $\pi_{\phi}$ aims to divide a complex task $\ell$ into a sequence of subtasks $p_{i}$ (i.e., task plan), each serving as an instruction to low-level control policies $\pi_{\theta}$:
 $\begin{split}\left[p_{1},p_{2},\dots,p_{N}\right]\sim\pi_{\phi}(\ell,s_{t}).\end{split}$ (8)
-This process is often referred to as task decomposition and is also related to the task and motion planning (TAMP) setting.
+This process is sometimes referred to as task or subgoal decomposition and is closely related to task and motion planning (TAMP) and embodied decision making.
+When equipped with task planners, VLAs can complete more complex, long-horizon tasks, as illustrated in Figure 4.
+Details are summarized in Table IV.
 Ideally, task plans should also incorporate optimal scheduling for these subtasks.
-When equipped with task planners, VLAs can complete more complex, long-horizon tasks, as illustrated in Figure 4.
-Table IV outlines important details about task planners.
 IV-A Monolithic Task Planners
 A single LLM or multimodal LLM (MLLM) can typically generate task plans by employing a tailored framework or through finetuning on embodied datasets. We refer to these as monolithic models.
-IV-A1 Grounded Task Planners
+IV-A1 End-to-end Task Planners
+Similar to LVLAs, task planners can be implemented as end-to-end multimodal LLMs, leveraging their Internet-scale knowledge for task planning.
+PaLM-E [11] integrates ViT and PaLM to create a large embodied multimodal language model capable of performing high-level embodied reasoning tasks.
+Based on perceived images and high-level language instructions, PaLM-E generates a text plan that serves as instructions for low-level robotic policies.
+In a mobile manipulation environment, they map the generated plan to executable low-level instructions with SayCan [10].
+As the low-level policy executes actions, PaLM-E can also replan based on changes in the environment.
+With PaLM as its backbone, PaLM-E can handle both normal visual question answering (VQA) tasks, along with the additional embodied VQA tasks.
+EmbodiedGPT [136] introduces the embodied-former, which outputs task-relevant instance-level features.
+This is achieved by incorporating information from vision encoder embeddings and embodied planning information provided by LLM.
+The instance feature serves to inform the low-level policy about the immediate next action to take.
+IV-A2 End-to-end Task Planners with 3D Vision
+Some task planners also explore the use of 3D vision.
+Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore they are usually end-to-end models.
+LEO [137] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning.
+LEO performs well not only in 3D question-answering tasks but also in manipulation, navigation, and task planning.
+3D-LLM [44] injects 3D information into LLMs and empowers them to perform 3D tasks, such as 3D-assisted dialog and navigation, using features from point cloud, gradSLAM, and neural voxel field.
+MultiPLY [150] is an object-centric embodied LLM that incorporates even more modalities, including audio, tactile, and thermal.
+ShapeLLM [138] is built on the novel 3D vision encoder ReCon++, which distills knowledge from multi-view image and text teachers, along with a point cloud MAE.
+By integrating ReCon++ with LLaMA, ShapeLLM improves embodied interaction performance on their newly proposed 3D benchmark, 3D MM-Vet.
+IV-A3 Grounded Task Planners
 Grounded task planning involves generating high-level actions while considering whether they can be executed by low-level control policies.
@@
 This approach enables a hierarchical policy to discover reusable skills with sparse natural language annotations.
-IV-A2 End-to-end Task Planners
-Similar to LVLAs, task planners can be implemented as end-to-end multimodal LLMs, leveraging their Internet-scale knowledge for task planning. Clearly, they are also monolithic models.
-PaLM-E [11] integrates ViT and PaLM to create a large embodied multimodal language model capable of performing high-level embodied reasoning tasks.
-Based on perceived images and high-level language instructions, PaLM-E generates a text plan that serves as instructions for low-level robotic policies.
-In a mobile manipulation environment, they map the generated plan to executable low-level instructions with SayCan [10].
-As the low-level policy executes actions, PaLM-E can also replan based on changes in the environment.
-With PaLM as its backbone, PaLM-E can handle both normal visual question answering (VQA) tasks, along with the additional embodied VQA tasks.
-EmbodiedGPT [156] introduces the embodied-former, which outputs task-relevant instance-level features.
-This is achieved by incorporating information from vision encoder embeddings and embodied planning information provided by LLM.
-The instance feature serves to inform the low-level policy about the immediate next action to take.
-IV-A3 End-to-end Task Planners with 3D Vision
-Some task planners also explore the use of 3D vision.
-Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore they are usually end-to-end models.
-LEO [159] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning.
-LEO demonstrates proficiency not only in 3D question-answering tasks but also in manipulation, navigation, and task planning.
-3D-LLM [57] injects 3D information into LLMs and empowers them to perform 3D tasks, such as 3D-assisted dialog, and navigation.
-The 3D features can come in diverse forms, including point cloud, gradSLAM, and neural voxel field.
-MultiPLY [179] is an object-centric embodied LLM that incorporates even more modalities, including audio, tactile, and thermal.
-ShapeLLM [165] is built on the novel 3D vision encoder ReCon++, which distills knowledge from multi-view image and text teachers, along with a point cloud MAE.
-By integrating ReCon++ with LLaMA, ShapeLLM improves embodied interaction performance on their newly proposed 3D benchmark, 3D MM-Vet.
+IV-B Modular Task Planners
+Finetuning end-to-end models on embodied data can be expensive, and some approaches adopt a modular design by assembling off-the-shelf LLMs and VLMs into task planners.
+These approaches can also be viewed as following the tool-use architecture [135].
 (a) Language-based
@@
 Figure 6: Different approaches to connect LLM to multi-modal modules in modular task planners.
-IV-B Modular Task Planners
-Finetuning end-to-end models on embodied data can be expensive, and there are approaches that adopt a modular design by assembling off-the-shelf LLMs and VLMs into task planners.
 IV-B1 Language-based Task Planners
@@
 In addition to natural language plans, task plans can also be represented in the form of pseudocode.
-Table V: Recent datasets for robot learning.
-Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects.
-$\ast$ Dataset collected in simulators rather than in the real world.
-Some datasets are continuously updated; this table is based on the information from their original papers.
-Table adapted from [180, 146].
-Name Collection Visual Instruction Embodiment Skills Tasks Scenes Objects Episodes
-MIME [181] Human Teleoperation RGB, D Demo Baxter 12 20 1 - 8.3K
-RoboTurk [182] $\ast$ Human Teleoperation RGB - Sawyer 2 - 1 - 2.1K
-RoboNet [183] Scripted RGB Goal image (7 robots) - - 10 - 162K
-MT-Opt [184] Scripted RGB Lang (7 robots) 2 12 1 - 800K
-BC-Z [113] Human Teleoperation RGB Lang, Demo Everyday 3 100 1 - 25.9K
-RT-1-Kitchen [114] Human Teleoperation RGB Lang Everyday 12 700+ 2 16 130K
-MOO [111] Human Teleoperation RGB Multi-modal Everyday 5 - - 106 59.1K
-VIMA [150] $\ast$ Scripted RGB Multi-modal UR5 17 - 1 29 650K
-RoboSet [185] Human, Scripted RGB, D Lang Franka Panda 12 38 11 - 98.5K
-BridgeData V2 [186] Human, Scripted RGB, D Lang WidowX 250 13 - 24 100+ 60.1K
-RH20T [180] Human Teleoperation RGB, D Lang (4 robots) 42 147 7 - 110K+
-DROID [146] Human Teleoperation RGB, D Lang Franka Panda 86 - 564 - 76K
-OXE [143] Aggregated Datasets RGB, D Lang (22 robots) 527 160,266 311 - 1M+
 IV-B2 Code-based Task Planners
@@
 This graph can then be translated into a text description (JSON), offering rich semantic and spatial relationships between entities to LLMs for task planning.
-Pros and Cons.
+Strengths and Limitations.
 Monolithic task planners that utilize grounded task planning focus on generating executable plans.
-End-to-end models like PaLM-E share an architecture similar to most LVLAs and can be finetuned on specialized embodied data to achieve better performance.
-However, the training costs of such large models can be a significant concern.
+End-to-end models share an architecture similar to most LVLAs and can be finetuned on specialized embodied data to achieve better performance.
+However, the training costs of such large models can be a concern.
 In contrast, modular task planners are more readily deployable because they leverage off-the-shelf LLMs and VLMs.
@@
 However, they often require extra steps to align the generated task plans with language instructions that are admissible to low-level control policies.
-On the other hand, while code-based task planners may require manually wrapping VLMs and control policies in APIs and preparing clear documentation in advance, they enable code debugging and provide greater controllability.
+Conversely, while code-based task planners may require manually wrapping VLMs and control policies in APIs and preparing clear documentation in advance, they enable code debugging and provide greater controllability.
 Nevertheless, their performance can be constrained by the programming capabilities of existing models.
 V Datasets and Benchmarks
-Table VI: Simulators and simulated benchmarks.
-Control: Continuous control tasks.
-D, S, A, N: Depth, segmentation, audio, normals.
-Force: simulated contact force between end-effector and item.
-PD: Pre-defined.
-Table adapted from [188, 189].
-Name Scenes
-/Rooms
-Objects
-/Cat
-UI Physics Engine Task Observation Action Agent Description Related
-Gibson [190] 572/- - - Pybullet Navi RGB, D, N, S - - Navi only -
-iGibson [191, 192, 189] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, S, N, Flow, LiDAR Force TurtleBot v2, LoCoBot, etc.
-VR, Continuous Extended States.
-Versions: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [188], BEHAVIOR-1K [193]
-SAPIEN [194] - 2346/- Code PhysX Navi, Mani RGB, D, S Force Franka Articulation, Ray Tracing VoxPoser. Benchmarks: SIMPLER [195]
-AI2-THOR [196] -/120 118/118 Mouse Unity Navi, Mani RGB, D, S, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Versions: [197, 198] Benchmarks: ALFRED, RoPOR [199]
-VirtualHome [200] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
-TDW [201] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids -
-RLBench [102] 1/- 28/28 Code Bullet Mani RGB, D, S Force Franka Tiered Task Difficulty Hiveformer, PerAct
-Meta-World [202] 1/- 80/7 Code MuJoCo Mani Pose Force Sawyer Meta-RL R3M, VC-1, Vi-PRoM, EmbodiedGPT
-CALVIN [203] 4/- 7/5 - Pybullet Mani RGB, D Force Franka Long-horizon Lang-cond tasks GR-1, HULC, RoboFlamingo
-Franka Kitchen [204] 1/- 10/6 VR MuJoCo Mani Pose Force Franka Extended by R3M with RGB R3M, Voltron, Vi-PRoM, Diffusion Policy, EmbodiedGPT
-Habitat [205, 206] Matterport + Gibson Mouse Bullet Navi RGB, D, S, A Force Fetch, Franka, AlienGO Fast, Navi only. Versions: Rearrangement [207], Habitat 2.0 VC-1, PACT; OVMM [208]
-ALFRED [209] -/120 84/84 - Unity Navi, Mani RGB, D, S PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
-DMC [210] 1/- 4/4 Code MuJoCo Control RGB, D Force - Continuous RL VC-1, SMART
-OpenAI Gym [211] 1/- 4/4 Code MuJoCo Control RGB Force - Single agent RL environments -
-Genesis [212] (Rigid, deformable, liquid, etc.) Code (Propri-etary) Navi, Mani RGB, D, S, N Force Franka, Unitree, etc. High-speed comprehensive physics simulation -
 V-A Real-world Robot Datasets & Benchmarks
@@
 Another technique, known as real-to-sim, can improve simulation fidelity, recreate failure cases, or facilitate digital twins.
+Table V: Embodied Datasets.
+Per RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects.
+$\ast$ Dataset collected in simulation rather than the real world.
+Some datasets are continually updated, and we include only the original version.
+Table adapted from [153, 154].
+Dataset Skills Tasks Scenes Episodes Collection Obs. Instruction Robots
+MIME [155] 12 20 1 8.3K Human teleop. RGBD Demo Baxter
+[…]
+RoboNet [157]     10 162K Script RGB Goal image (7 robots)
+MT-Opt [158] 2 12 1 800K Script RGB Lang (7 robots)
+BC-Z [79] 3 100 1 25.9K Human teleop. RGB Lang, demo EDR
+Fractal [94] 12 700+ 2 130K Human teleop. RGB Lang EDR
+MOO [93] 5     59.1K Human teleop. RGB Multimodal EDR
+[…]
+RoboSet [159] 12 38 11 98.5K Human, script RGBD Lang Franka
+BridgeV2 [160] 13   24 60.1K Human, script RGBD Lang WidowX
+RH20T [153] 42 147 7 110K+ Human teleop. RGBD Lang (4 robots)
+DROID [154] 86   564 76K Human teleop. RGBD Lang Franka
+OXE [112] 527 160K 311 1M+ Aggregate data RGBD Lang (22 robots)
+Table VI: Simulators and simulated benchmarks.
+Control: continuous control tasks.
+D, S, A, N: depth, segmentation, audio, normal.
+Force: simulated contact force between end-effector and item.
+PD: pre-defined.
+Table adapted from [161, 162].
+Name Scenes
+/Rooms
+Objects
+/Cat
+UI Physics Engine Task Observation Action Agent Description Related
+Gibson [163] 572/-     Pybullet Navi RGB, D, N, S     Navi only
+iGibson [164, 165, 162] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, S, N, Flow, LiDAR Force TurtleBot v2, LoCoBot, etc.
+VR, Continuous Extended States.
+Versions: 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [161], BEHAVIOR-1K [166]
+SAPIEN [167]   2346/- Code PhysX Navi, Mani RGB, D, S Force Franka Articulation, Ray Tracing VoxPoser. Benchmark: SIMPLER [168]
+AI2-THOR [169] -/120 118/118 Mouse Unity Navi, Mani RGB, D, S, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Versions: [170, 171] Benchmarks: ALFRED, RoPOR [172]
+VirtualHome [173] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
+TDW [174] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids
+RLBench [175] 1/- 28/28 Code Bullet Mani RGB, D, S Force Franka Tiered Task Difficulty Hiveformer, PerAct
+Meta-World [176] 1/- 80/7 Code MuJoCo Mani Pose Force Sawyer Meta-RL R3M, VC-1, Vi-PRoM, EmbodiedGPT
+CALVIN [177] 4/- 7/5   Pybullet Mani RGB, D Force Franka Long-horizon Lang-cond tasks GR-1, HULC, RoboFlamingo
+Franka Kitchen [178] 1/- 10/6 VR MuJoCo Mani Pose Force Franka Extended by R3M with RGB R3M, Voltron, Vi-PRoM, Diffusion Policy, EmbodiedGPT
+Habitat [179, 180] (Matterport + Gibson) Mouse Bullet Navi RGB, D, S, A Force Fetch, Franka, AlienGO Fast, Navi only. Versions: 1.0, 2.0, Rearrangement [181] VC-1, PACT; Benchmark: OVMM [182]
+ALFRED [183] -/120 84/84   Unity Navi, Mani RGB, D, S PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
+DMC [184] 1/- 4/4 Code MuJoCo Control RGB, D Force   Continuous RL VC-1, SMART
+OpenAI Gym [185] 1/- 4/4 Code MuJoCo Control RGB Force   Single agent RL environments
+Genesis [186] (Rigid, deformable, liquid, etc.) Code (Propri-etary) Navi, Mani RGB, D, S, N Force Franka, Unitree, etc. High-speed comprehensive physics simulation
 V-C Automated Dataset Collection
@@
 OpenEQA [204] evaluates seven main categories, including functional reasoning and world knowledge.
-Table VII: Embodied question answering benchmarks. Acc: accuracy. PPL: perplexity.
-# QA # Videos Video source Answer type Active Data collection Metrics
-EQA [223] 5K 750 envs House3D simulator Answer set 172 Yes Template Acc
-IQUAD [224] 75K 30 rooms AI2-THOR Multiple-choice Yes Template Acc
-MT-EQA [225] 19.3K 588 envs House3D simulator Binary answer Yes Template Acc
-MP3D-EQA [226] 1,136 83 envs MatterPort3D Answer set 53 Yes Template Acc
-EgoVQA [227] 600 16 IU Multi-view Multiple-choice (1 out of 5) No Human annotators Acc
-EgoTaskQA [228] 40K 2K LEMMA Answer set (open-answer and binary verifications) No Human annotators for relationship triples, then template Acc
-EQA-MX [229] 8.2M 750K images CAESAR simulator Answer set No Question templates, answer set (simulator ground truth) Acc
-OpenEQA [230] 557 + 1079 180 envs HM3D + ScanNet Open-answer (LLM scorer) No Human annotators Score
+Table VII: Embodied question answering benchmarks. Explore: active exploration.
+Benchmark QA Pairs Scenes Source Answer Type Collection Explore Metrics
+EQA [197] 5K 750 envs House3D simulator Answer set (172 answers) Template Accuracy
+IQUAD [198] 75K 30 rooms AI2-THOR Multiple choice Template Accuracy
+MT-EQA [199] 19.3K 588 envs House3D simulator Binary answer Template Accuracy
+MP3D-EQA [200] 1,136 83 envs MatterPort3D Answer set (53 answers) Template Accuracy
+EgoVQA [201] 600 16 videos IU Multi-view Multiple choice (1 out of 5) Human annotators Accuracy
+EgoTaskQA [202] 40K 2K videos LEMMA dataset Open answer, binary verification Human annotators & template Accuracy
+EQA-MX [203] 8.2M 750K images CAESAR simulator Answer set Question templates, answer set Accuracy
+OpenEQA [204] 557 + 1079 180 envs HM3D + ScanNet Open answer Human annotators LLM Score
 VI Challenges and Future Directions
@@
 Attaining such a level of generalization is very challenging, as it requires the development of many core AGI capabilities.
-Multi-Modality
+Multimodality
 VLAs inherit many challenges associated with multimodal models, such as obtaining useful embeddings and aligning different modalities.
@@
 However, whether focusing on embeddings alone is sufficient remains under debate.
-Although beyond the scope of VLAs, other modalities—such as audio [60] and haptics [233]—have proven useful for certain embodied AI applications.
-Incorporating additional modalities is often desirable, but they also introduce extra complexity into the model design.
+Although beyond the scope of VLAs, other modalities—such as audio [47], haptics [207], and gaze [208]—have proven useful for certain embodied AI applications.
+For instance, modeling human gaze data using an additional gaze network [209] or incorporating it via an auxiliary loss [210] has been shown to enhance the performance of the primary policy network.
+These approaches demonstrate that RL policies can benefit from human-like visual attention, as human gaze often reveals the most salient locations in the environment [211].
+While incorporating additional modalities is often advantageous, it inevitably introduces extra complexity into the model design.
 Framework for Long-Horizon Tasks
@@
 Ethical and Societal Implications
-Robotics has always raised various ethical, legal, and societal concerns.
-These include risks related to privacy, safety, job displacement, bias in decision-making, and the impact on social norms and human relationships.
+Robotics has always raised various ethical, societal, and legal concerns.
+These include risks related to privacy, job displacement, decision-making bias, and the impact on social norms and human relationships.
 Applications
-Most current VLAs focus on household and industrial settings [2], which shape the available datasets and models.
-Given the potential of VLA models, a wide range of new applications is possible.
-One particularly important application is in healthcare, such as surgical robots [234] and care robots [235].
+Most current VLAs focus on household or industrial settings, but a wider range of applications is possible, such as virtual assistants, autonomous vehicles, and agricultural robots.
+Various embodiments may also call for specialized VLAs, including dexterous hands, drones, quadruped robots, and humanoid robots.
+One particularly important field is healthcare, encompassing surgical robots [212], care robots [213].
 Healthcare demands higher safety and privacy standards and may necessitate novel techniques such as human-in-the-loop (HITL) control and federated learning.
@@
 Vision-language-action models hold immense promise for enabling embodied agents to interact with the physical world and fulfill users’ instructions.
-This paper is the first survey to review large VLAs alongside other generalized VLAs.
+This paper is the first survey to review large VLAs alongside generalized VLAs.
 Our taxonomy provides a high-level overview of three main lines of research: key components, control policies, and task planners.
-We meticulously summarize and compare their model designs and technical details.
+We meticulously analyze and compare their technical details, including model architectures, training strategies, and individual modules.
 Additionally, we highlight essential resources for training and evaluating VLAs, such as datasets, simulators, and benchmarks.

### v7

@@
 Abstract
-Embodied AI is widely recognized as a key element of artificial general intelligence because it involves controlling embodied agents to perform tasks in the physical world.
+Embodied AI is widely recognized as a cornerstone of artificial general intelligence because it involves controlling embodied agents to perform tasks in the physical world.
 Building on the success of large language models and vision-language models, a new category of multimodal models—referred to as vision-language-action models (VLAs)—has emerged to address language-conditioned robotic tasks in embodied AI by leveraging their distinct ability to generate actions.
-In recent years, a myriad of VLAs have been developed, making it imperative to capture the rapidly evolving landscape through a comprehensive survey.
+The recent proliferation of VLAs necessitates a comprehensive survey to capture the rapidly evolving landscape.
 To this end, we present the first survey on VLAs for embodied AI.
@@
 The first line focuses on individual components of VLAs.
-The second line is dedicated to developing control policies adept at predicting low-level actions.
+The second line is dedicated to developing VLA-based control policies adept at predicting low-level actions.
 The third line comprises high-level task planners capable of decomposing long-horizon tasks into a sequence of subtasks, thereby guiding VLAs to follow more general user instructions.
 Furthermore, we provide an extensive summary of relevant resources, including datasets, simulators, and benchmarks.
-Finally, we discuss the challenges faced by VLAs and outline promising future directions in embodied AI.
-We have created a project associated with this survey, which is available at https://github.com/yueen-ma/Awesome-VLA.
-Index Terms:
-Vision-Language-Action Model, Embodied AI, Robotics, Multi-modality, Large Language Model, World Model
+Finally, we discuss the challenges facing VLAs and outline promising future directions in embodied AI.
+A curated repository associated with this survey is available at: https://github.com/yueen-ma/Awesome-VLA.
 I Introduction
-Vision-language-action models (VLAs) are a class of multimodal models within the field of embodied AI, designed to process information from vision, language, and action modalities.
-Unlike conversational AI such as ChatGPT [1], embodied AI requires controlling physical embodiments that interact with the environment and robotics is the most prominent domain of embodied AI.
+Vision-language-action models (VLAs) are a class of multimodal models within the field of embodied AI, designed to process information from three modalities: vision, language, and action.
+Unlike conversational AI, embodied AI requires controlling physical embodiments that interact with the environment, and robotics is the most prominent domain of embodied AI.
 In language-conditioned robotic tasks, the policy must possess the capability to understand language instructions, visually perceive the environment, and generate appropriate actions, necessitating the multimodal capabilities of VLAs.
@@
 Early developments in deep learning primarily consist of unimodal models.
-In computer vision (CV), models like AlexNet [3] showcased the potential of artifical neural networks.
-Recurrent neural networks laid the groundwork for numerous natural language processing (NLP) models but have seen a transition in recent years, with Transformers [4] taking precedence.
-The Deep Q-network [5] demonstrated that ANNs can successfully tackle reinforcement learning problems.
+In computer vision (CV), models like AlexNet [2] showcased the potential of artificial neural networks (ANNs).
+Recurrent neural networks laid the groundwork for numerous natural language processing (NLP) models, but have seen a transition in recent years, with Transformers [3] taking precedence.
+The Deep Q-network [4] demonstrated that ANNs can successfully tackle reinforcement learning (RL) problems.
 Leveraging advancements of unimodal models across diverse machine learning fields, multimodal models now address a wide range of tasks, such as visual question answering, image captioning, and text-to-video generation.
@@
 II Background
-[…]
-[…]
+{forest}
 Figure 3: The taxonomy of VLA models. The organization of this survey follows this taxonomy.
 Embodied AI is a unique form of artificial intelligence that actively interacts with the physical environment.
-This sets it apart from other AI models, such as conversational AI, which primarily handles textual conversations (ChatGPT [1]), or generative AI models that focus on tasks like text-to-video generation (Sora [16]).
+This sets it apart from other AI models, such as conversational AI, which primarily handles textual conversations, or generative AI models that focus on tasks like text-to-video generation.
 Embodied AI encompasses a broad spectrum of embodiments, including smart appliances, smart glasses, autonomous vehicles, and more.
 Among them all, robots stand out as one of the most prominent embodiments.
-Robot learning is also usually framed as reinforcement learning problems, represented as a Markov Decision Process (MDP) consisting of states ($s$), actions ($a$), and rewards ($r$).
+Robot learning is also usually framed as a reinforcement learning problem, represented as a Markov Decision Process (MDP) consisting of states ($s$), actions ($a$), and rewards ($r$).
 A MDP trajectory is denoted as $\tau=(s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T})$.
-In certain scenarios, robotic tasks may also be viewed as partially-observable Markov Decision Processes (POMDPs) due to incomplete observations.
+In certain scenarios, robotic tasks may also be viewed as Partially Observable Markov Decision Processes (POMDPs) due to incomplete observations.
 The primary objective of reinforcement learning is to train a policy capable of generating optimal actions for the current state $\pi(a_{t}|s_{\leq t},a_{<t})$.
-Various methods such as TD learning, policy gradients, etc., can be employed to achieve this.
+Various methods, such as TD learning, policy gradients, etc., can be employed to achieve this.
 However, in cases where defining the reward function proves challenging, imitation learning is utilized to directly model the action distribution within trajectories devoid of rewards $\tau=(s_{1},a_{1},\dots,s_{T},a_{T})$.
@@
 Gato [17] further extended this paradigm to a multimodal, multitask, multi-embodiment setting.
-These methods in RL directly inspired modern VLAs [20, 21], which also process multimodal sequences using transformer-based architectures.
+$\pi_{0.6}^{*}$ [18] employs RL for VLAs to learn from experience.
 Furthermore, a synergy has emerged between RL and LLMs, benefiting embodied AI in multiple ways.
@@
 SEED [19] utilizes RLHF alongside skill-based RL to address the sparse reward issue and to improve safety via human evaluation.
-Conversely, LLMs also enable novel RL methods: Reflexion [23] proposes a novel verbal reinforcement learning framework that replaces weight updates in RL models with linguistic feedback, which is naturally applicable to embodied decision making where planning also occurs in language form.
+Conversely, LLMs also enable novel RL methods.
+Reflexion [20] proposes a novel verbal reinforcement learning framework that replaces weight updates in RL models with linguistic feedback, which is naturally applicable to embodied decision-making, where planning also occurs in language form.
 Eureka [21] shows that LLMs can design reward functions for embodied agents that outperform expert human-engineered ones.
@@
 R3M [23] proposes two main pretraining objectives: time contrastive learning and video-language alignment.
-The objective of time contrastive learning is to minimize the distance between video frames that are temporally close while simultaneously increasing the separation between frames that are temporally distant.
+The objective of time contrastive learning is to minimize the distance between temporally proximal video frames while simultaneously increasing the separation between temporally distant ones.
 This objective aims to create PVRs that capture the temporal relationship within the video sequence.
@@
 This objective enriches the semantic relevance embedded in the PVRs.
-VIP [27] also capitalizes on video temporal relationship.
+VIP [24] also capitalizes on the video temporal relationship.
 MVP [25] adopts masked autoencoder (MAE) from computer vision to robotic datasets.
@@
 NeRF can be extracted from videos and contains rich 3D information for robot learning, as exemplified by F3RM [40] and 3D-LLM [41].
-The recent 3D Gaussian Splatting (3D-GS) method has been gaining popularity in computer vision as well as in embodied AI.
-3D Gaussians can be deformed through physics-based simulation for generative dynamics, as proposed in PhysGaussian [45], or they can serve as 3D representations for VLMs, as enabled by UniGS [46].
+The recent 3D Gaussian Splatting (3D-GS) [42] surpasses NeRF in visual quality and rendering speed.
 Additionally, many videos contain audio, which can provide important cues for robot policies [43].
@@
 LLMs encompass a wealth of commonsense knowledge about the world, prompting many approaches to leverage that knowledge for improving VLAs.
-DECKARD [61] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs, specifically tailored for the task of item crafting in Minecraft.
+DECKARD [57] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs [58, 59], specifically tailored for the task of item crafting in Minecraft.
 DECKARD iterates between two phases: in the Dream phase, it samples a subgoal guided by the AWM; in the Wake phase, DECKARD executes the subgoal and updates the AWM through interactions with the game.
@@
 3D-VLA [68] proposes a 3D world model capable of goal generation.
-It processes visual inputs, such as images, depth map, and point clouds, and then generates a goal state—either as an image or a point cloud—using diffusion models in response to the user’s query.
+It processes visual inputs, such as images, depth maps, and point clouds, and then generates a goal state—either as an image or a point cloud—using diffusion models in response to the user’s query.
 The generated goal state can subsequently be used to guide robot control.
@@
 By relying on this multi-step reasoning rather than the “muscle memory” of VLAs, it improves success rates on challenging generalization tasks without requiring additional robot data.
+CoT-VLA [77] introduces visual CoT reasoning for VLAs.
+III-A9 Policy Steering
+Policy steering can enhance VLA performance at test-time without the need for expensive retraining.
+V-GPS [78] re-ranks generated actions based on a learned value function, while RoboMonkey [79] employs a VLM-based verifier to select the optimal action from a sampled set.
 Strengths and Limitations.
@@
 MCIL [81] represents a pioneering robot policy that integrates free-form natural language conditioning.
-This is in contrast to earlier approaches that typically rely on conditions in the form of task ID or goal image.
+This is in contrast to earlier approaches that typically rely on conditions in the form of task IDs or goal images.
 MCIL introduces the capability to leverage unlabeled and unstructured demonstration data.
@@
 Hiveformer [87] places significant emphasis on leveraging multi-view scene observations and maintaining the full observation history for a language-conditioned policy.
-This approach represents an advancement over previous systems such as CLIPort and BC-Z that only use the current observation.
+This approach represents an advancement over previous systems, such as CLIPort and BC-Z, that only use the current observation.
 Notably, Hiveformer stands out as one of the early adopters of Transformer architecture as its policy backbone.
@@
 Consequently, Gato enables the simultaneous training of different tasks.
-Representing a significant milestone, Gato exemplifies the potential of constructing a “multi-modal, multi-task, multi-embodiment generalist agent”.
+Astra [126] optimizes this architecture via trajectory attention.
 RoboCat [93] proposes a self-improvement process designed to enable an agent to rapidly adapt to new tasks with as few as 100 demonstrations.
 This self-improvement process iteratively finetunes the model and self-generates new data with the finetuned model.
-Built upon the Gato model, RoboCat incorporates the VQ-GAN image encoder [124].
+Built upon the Gato model, RoboCat incorporates the VQ-GAN image encoder.
 During training, RoboCat predicts not only the next action but also future observations.
@@
 RT-1 [95], developed by the same team as BC-Z, shares similarities with BC-Z but introduces some key distinctions.
-Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet [125], departing from BC-Z’s use of ResNet.
+Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet, departing from BC-Z’s use of ResNet.
 However, RT-1 does not use video as a task instruction.
@@
 During inference, the action sequence is averaged using a method called temporal ensembling.
-RoboAgent [98] extends this approach with its MT-ACT model, demonstrating that action chunking improves temporal consistency.
-Additionally, RoboAgent introduces a semantic augmentation method that leverages inpainting to augment existing demonstrations.
-RoboFlamingo [99] adapts the existing Flamingo VLM to a robot policy by attaching an LSTM-based policy head to the VLM.
+RoboAgent [99] extends ACT via MT-ACT, demonstrating that action chunking improves temporal consistency.
+It also introduces an inpainting-based semantic augmentation method.
+RoboFlamingo [100] adapts the existing Flamingo VLM to a robot policy by attaching an LSTM-based policy head.
 This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
+A recent trend in LLMs is equipping them with tool-use capabilities by generating code that calls tools via APIs [127].
+Instruct2Act [103] follows this paradigm by integrating vision and action tools, enabling LLMs to perform robotic tasks.
 (a) FiLM
@@
 (c) Concatenation
-(d) Quantization
-(e) Tool Use
-Figure 5: Representative architectures of VLA models. $\odot$: Hadamard product. $\oplus$: Concatenation.
+(d) Tool Use
+(e) Standard LVLA
+(f) Action Expert
+(g) Dual-System
+(h) VLA + World Model
+(i) Quantization
+Figure 5: Representative architectures of VLA models. Some correspond to the “Architecture” column in Table III. $\odot$: Hadamard product. $\oplus$: Concatenation.
 III-B3 Control Policies for Multimodal Instructions
@@
 VIMA [128] places a significant emphasis on multimodal prompts and the generalization capabilities of models.
-By incorporating multimodal prompts, more specific and intricate tasks can be formulated compared to traditional pure text prompts.
-VIMA introduces four main types of tasks: object manipulation, visual goal reaching, novel concept grounding, one-shot video imitation, visual constraint satisfaction, visual reasoning.
+By incorporating multimodal prompts, more specific and intricate tasks can be formulated compared to pure text prompts.
+VIMA introduces four main types of tasks: object manipulation, visual goal reaching, novel concept grounding, one-shot video imitation, visual constraint satisfaction, and visual reasoning.
 These tasks are often challenging or even infeasible to express using only language prompts.
-VIMA-Bench has been developed to evaluate across four generalizability levels: placement, combinatorial, novel object, novel task.
+VIMA-Bench has been developed to evaluate across four generalizability levels: placement, novel combinatorial, novel object, and novel task.
 MOO [94] extends RT-1 to handle multimodal prompts.
@@
 By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1.
-This extension also facilitates new methods of specifying target objects, such as pointing with a finger, and clicking on graphical user interfaces.
+This extension also facilitates new methods of specifying target objects, such as pointing with a finger and clicking on graphical user interfaces.
 III-B4 Control Policies with 3D Vision
@@
 RoboUniView [101] improves performance by injecting 3D information from multi-perspective images through a novel UVFormer vision encoder, which is pretrained on a 3D occupancy task.
-In contrast, Act3D [88] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization.
+In contrast, Act3D [89] introduces a continuous resolution 3D feature field, with adaptive resolutions based on the current task, addressing the computational cost of voxelization.
 RVT, RVT-2 [90, 91] propose re-rendering images from virtual views of the scene point cloud and using these images as inputs, rather than directly relying on 3D inputs.
@@
 Diffusion-based action generation leverages the success of diffusion models in the field of computer vision.
-Diffusion Policy [104] formulates a robot policy as a DDPM [128].
+Diffusion Policy [105] formulates a robot policy as a DDPM.
 This approach incorporates a variety of techniques, including receding horizon control, visual conditioning, and the time-series diffusion Transformer.
@@
 Being among the first to utilize the OXE dataset [113], Octo demonstrates positive transfer and generalizability across diverse robots and tasks.
-MDT [109] adapts the newly introduced DiT model [129] from computer vision to the action prediction head.
+MDT [110] adapts the newly introduced DiT model from computer vision to the action prediction head.
 DiT, originally proposed as a Transformer-based diffusion model, replaces the classical U-Net architecture for video generation.
@@
 VoxPoser [104] employs LLM and VLM to create two 3D voxel maps that represent affordance and constraint.
-It leverages the programming capability of LLMs, and the perception capability of VLMs.
+It leverages the programming capability of LLMs and the perception capability of VLMs.
 LLM translates language instructions into executable code, invoking VLM to obtain object coordinates.
@@
 This terminology is analogous to the distinction between LLMs and general language models, or between large VLMs and general VLMs.
-RT-2 [2] endeavors to harness the capabilities of large multimodal models in robotics tasks, drawing inspiration from models like PaLI-X and PaLM-E.
+RT-2 [1] endeavors to harness the capabilities of large multimodal models in robotics tasks, drawing inspiration from models like PaLI-X and PaLM-E (see the architecture in Fig.
+5e).
 The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data.
-This training scheme enhances the model’s generalizability and brings about emergent capabilities.
+This training scheme enhances generalizability and brings about emergent capabilities.
 RT-H [112] introduces an action hierarchy that includes an intermediate prediction layer of language motions, situated between language instructions and low-level actions (translation and rotation).
@@
 RT-X [113] builds upon the previous RT-1 and RT-2 models.
-These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets.
+These models are retrained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets.
 The resulting models, RT-1-X and RT-2-X, both outperform their original versions.
@@
 $\pi_{0}$ [116] proposes a flow-matching architecture for transforming VLMs into VLAs.
-By incorporating an additional action expert based on the mixture-of-experts framework, it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
+By incorporating an additional action expert based on the mixture-of-experts framework (see the architecture in Fig.
+5f), it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
 RoboMamba [117] replaces the costly Transformer with a Mamba state space model that has linear inference complexity, achieving efficient robotic reasoning and action capabilities.
@@
 HybridVLA [122] integrates diffusion with the autoregression paradigm to fully leverage VLMs’ reasoning capabilities.
+GR00T N1 [133] introduces a dual-system architecture to build a robot foundation model for humanoid robots (see the architecture in Fig.
+5g).
+Its VLM-based System 2 processes image observations and language instructions at 10 Hz, while System 1 is a diffusion module that generates closed-loop motor actions in real time (120 Hz).
+NORA-1.5 [134] unifies a VLA with a world model through reward-guided post-training.
+Genie Envisioner [135] is a world foundation platform that integrates a world model and a VLA within a single video-generative framework.
+See the general architecture in Fig.
+5h.
 Visual autoregressive modeling (VAR) [136] with quantized visual tokens demonstrates improved performance over diffusion models in image generation.
 This suggests that the three modalities of VLAs can be unified under the autoregressive paradigm [137].
-WorldVLA [122] and UniVLA [123] advance this direction by integrating VLAs with world models.
+WorldVLA [124] and UniVLA [125] advance this direction by integrating VLAs with world models (see the architecture in Fig.
+5i).
 It quantizes multimodal data into discrete tokens, forming a shared vocabulary of quantized multimodal tokens.
 Consequently, all modalities can be modeled autoregressively, enabling not only action and text generation but also image generation, thereby constituting a world model.
-A recent trend in LLMs is equipping them with tool-use capabilities by generating code that calls tools via APIs [135].
-Instruct2Act [102] follows this paradigm by integrating vision and action tools, enabling LLMs to perform robotic tasks.
 Strengths and Limitations.
@@
 Representative VLA architectures are illustrated in Figure 5.
-FiLM is used in RT-1 and thus its follow-up models inherit this mechanism.
+FiLM is used in RT-1, and thus its follow-up models inherit this mechanism.
 While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [128].
@@
 While this abstraction facilitates better generalization to different embodiments, it also imposes limitations on dexterity.
-The behavior cloning (BC) objective is used in imitation-learning, with different variants for different action types. The BC objective for continuous action [99] can be written as:
+The behavior cloning (BC) objective is used in imitation-learning, with different variants for different action types. The BC objective for continuous action can be written as:
 $\mathcal{L}_{\text{Cont}}=\sum_{t}\operatorname{MSE}\left(a_{t},\hat{a}_{t}\right),$ (4)
 where $\operatorname{MSE}(\cdot)$ stands for mean squared error. $a_{t}$ is the action annotation from expert demonstrations.
-Discrete action is achieved by dividing the action value range into a number of bins [94]. Its BC objective is:
+Discrete action is achieved by dividing the action value range into a fixed number of bins. Its BC objective is:
 $\mathcal{L}_{\text{Disc}}=\sum_{t}\operatorname{CE}\left(a_{t},\hat{a}_{t}\right),$ (5)
 where $\operatorname{CE}(\cdot)$ stands for cross-entropy loss.
-CLIPort [31] and VIMA [126] use $SE(2)$ action and its BC objective can be expressed as follows:
+The BC objective of $SE(2)$ action is:
 $\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}}).\end{split}$ (6)
-The DDPM objective in diffusion-based control policies [104] is represented as:
+The DDPM objective in diffusion-based control policies is:
 $\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right),$ (7)
@@
 However, more complex tasks—such as “pouring water into a cup”—may require additional DoFs, thereby necessitating $SE(3)$ actions [35].
-Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner [38].
+Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner.
 Table IV: High-level task planners.
@@
 Therefore, several methods have been proposed to improve efficiency.
-TinyVLA [118] focused on inference speed and data efficiency through smaller VLM and diffusion head for robot action.
+TinyVLA [119] focused on inference speed and data efficiency through a smaller VLM and diffusion head for robot action.
 DeeR-VLA [102] proposes only partially activating the model with dynamic inference with early-exit.
 Scaling Law
-Similar to LLMs, scaling laws have also been observed in robotics [149], revealing the importance of model size, data quality, and diversity in environments and objects.
-Further research in this area can guide the development of VLAs with improved in-the-wild generalization capabilities.
+Similar to LLMs, scaling laws have also been observed in robotics [151, 152], revealing the importance of model size, dataset size, and the diversity of environments and objects.
+Further research in this area can guide the development of VLAs with robust in-the-wild generalization capabilities.
 IV Task Planners
@@
 Some task planners also explore the use of 3D vision.
-Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore they are usually end-to-end models.
+Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore, they are usually end-to-end models.
 LEO [139] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning.
@@
 IV-B1 Language-based Task Planners
-Language-based approaches use natural language descriptions as the medium for exchanging multimodal information, as illustrated in Figure 6a.
+Language-based approaches use natural language descriptions as the medium for exchanging multimodal information, as shown in Figure 6a.
 Inner Monologue [8] sits between high-level command and low-level policy to enable closed-loop control planning.
-It employs LLM to generate language instructions for low-level control policies and dynamically updates these instructions based on feedback received from the control policies.
+It employs LLM to generate language instructions for low-level policies and dynamically updates these instructions based on feedback received from them.
 The feedback encompasses various sources: success feedback, object and scene feedback, and human feedback.
@@
 By converting all environmental inputs into textual descriptions, their language-model-based policy demonstrates enhanced combinatorial generalization.
-Socratic Models (SMs) [143] presents a unique framework wherein diverse pretrained models are effectively composed without the need for finetuning.
+Socratic Models (SMs) [145] present a unique framework wherein diverse pretrained models are effectively composed without the need for finetuning.
 The framework is based on the key component, named multimodal-informed prompting, facilitating information exchange among models with varied multimodal capabilities.
 The idea is to utilize multimodal models to convert non-language inputs into language descriptions, effectively unifying different modalities within the language space.
-Beyond excelling in conventional multimodal tasks, SMs showcases its versatility in robot perception and planning.
+Beyond excelling in conventional multimodal tasks, SMs showcase their versatility in robot perception and planning.
 In addition to natural language plans, task plans can also be represented in the form of pseudocode.
@@
 Environmental feedback can be incorporated through assertions within the program.
-ChatGPT for Robotics [145] take advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology.
+ChatGPT for Robotics [147] takes advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology.
 The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code.
-In this procedure, ChatGPT serves as a high-level task planner and actions are generated through function calls to corresponding low-level APIs.
+In this procedure, ChatGPT serves as a high-level task planner, and actions are generated through function calls to corresponding low-level APIs.
 Code as policies (CaP) [148] also leverages the code-writing capability of LLMs.
 It employs GPT-3 or Codex to generate policy code, which, in turn, invokes perception modules and control APIs.
-CaP exhibits proficiency on spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives.
+CaP exhibits proficiency in spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives.
 By leveraging the multimodal capabilities of GPT-4V, COME-robot [155] eliminates the need for perception APIs in CaP.
 This also opens up possibilities for open-ended reasoning and adaptive planning within a closed-loop framework, enabling capabilities such as failure recovery and free-form instruction following.
-DEPS [147] stands for “Describe, Explain, Plan and Select”.
+DEPS [149] stands for “Describe, Explain, Plan, and Select”.
 This approach employs an LLM to generate plans and explain failures based on feedback descriptions collected from the environment—a process referred to as “self-explanation”, aiding in re-planning.
@@
 V Datasets and Benchmarks
+Table V: Embodied Datasets.
+Per RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects.
+$\ast$ Dataset collected in simulation rather than the real world.
+Some datasets are continually updated, and we include only the original version.
+Table adapted from [156, 157].
+Dataset Skills Tasks Scenes Episodes Collection Obs. Instruction Robots
+MIME [158] 12 20 1 8.3K Human teleop. RGBD Demo Baxter
+[…]
+RoboNet [160]     10 162K Script RGB Goal image (7 robots)
+MT-Opt [161] 2 12 1 800K Script RGB Lang (7 robots)
+BC-Z [80] 3 100 1 25.9K Human teleop. RGB Lang, demo EDR
+Fractal [95] 12 700+ 2 130K Human teleop. RGB Lang EDR
+MOO [94] 5     59.1K Human teleop. RGB Multimodal EDR
+[…]
+RoboSet [162] 12 38 11 98.5K Human, script RGBD Lang Franka
+BridgeV2 [163] 13   24 60.1K Human, script RGBD Lang WidowX
+RH20T [156] 42 147 7 110K+ Human teleop. RGBD Lang (4 robots)
+DROID [157] 86   564 76K Human teleop. RGBD Lang Franka
+OXE [113] 527 160K 311 1M+ Aggregate data RGBD Lang (22 robots)
 V-A Real-world Robot Datasets & Benchmarks
@@
 Another technique, known as real-to-sim, can improve simulation fidelity, recreate failure cases, or facilitate digital twins.
-Table V: Embodied Datasets.
-Per RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects.
-$\ast$ Dataset collected in simulation rather than the real world.
-Some datasets are continually updated, and we include only the original version.
-Table adapted from [153, 154].
-Dataset Skills Tasks Scenes Episodes Collection Obs. Instruction Robots
-MIME [155] 12 20 1 8.3K Human teleop. RGBD Demo Baxter
-[…]
-RoboNet [157]     10 162K Script RGB Goal image (7 robots)
-MT-Opt [158] 2 12 1 800K Script RGB Lang (7 robots)
-BC-Z [79] 3 100 1 25.9K Human teleop. RGB Lang, demo EDR
-Fractal [94] 12 700+ 2 130K Human teleop. RGB Lang EDR
-MOO [93] 5     59.1K Human teleop. RGB Multimodal EDR
-[…]
-RoboSet [159] 12 38 11 98.5K Human, script RGBD Lang Franka
-BridgeV2 [160] 13   24 60.1K Human, script RGBD Lang WidowX
-RH20T [153] 42 147 7 110K+ Human teleop. RGBD Lang (4 robots)
-DROID [154] 86   564 76K Human teleop. RGBD Lang Franka
-OXE [112] 527 160K 311 1M+ Aggregate data RGBD Lang (22 robots)
 Table VI: Simulators and simulated benchmarks.
@@
 AI2-THOR [172] -/120 118/118 Mouse Unity Navi, Mani RGB, D, S, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Versions: [173, 174] Benchmarks: ALFRED, RoPOR [175]
-VirtualHome [173] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
+VirtualHome [176] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated⟨LM⟩, ProgPrompt
 TDW [177] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids
@@
 An alternative strategy to address data scarcity in real-world settings is to leverage human data.
-Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity.
+Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity [193].
 However, this strategy also comes with inherent drawbacks.
@@
 EQA is akin to previous visual question answering benchmarks, but differs in that the agent can actively explore the environment before providing an answer.
-EmbodiedQA [197] and IQUAD [198] were among the first works to introduce this type of benchmarks.
+EmbodiedQA [201] and IQUAD [202] were among the first works to introduce this type of benchmark.
 MT-EQA [203] focuses on complex questions involving multiple targets.
@@
 Foundation Models & Generalization
-VLA foundation models or robotic foundation models (RFM) for embodied AI remains an open research topic primarily due to the diversity in embodiments, environments, and tasks.
-Many have made significant progress [110, 115], but still lacks generalization capability on par with LLMs in NLP.
+VLA foundation models or robotic foundation models (RFM) for embodied AI remain an open research topic primarily due to the diversity in embodiments, environments, and tasks.
+Many have made significant progress [111, 116], but still lack generalization capability on par with LLMs in NLP.
 Attaining such a level of generalization is very challenging, as it requires the development of many core AGI capabilities.
@@
 We hope this survey captures the rapidly evolving landscape of embodied AI and inspires future research.
+Acknowledgments
+The research presented in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 2410072, RGC R1015-23) and CUHK 7010870.